### v3

# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

Abstract.
We study the multiplicity problem for prime closed orbits of dynamically convex Reeb flows on the boundary of a star-shaped domain in ${\mathbb{R}}^{2n}$. The first of our two main results asserts that such a flow has at least $n$ prime closed Reeb orbits, improving the previously known lower bound by a factor of two and settling a long-standing open question. The second main theorem is that when, in addition, the domain is centrally symmetric and the Reeb flow is non-degenerate, the flow has either exactly $n$ or infinitely many prime closed orbits. This is a higher-dimensional contact variant of Franks’ celebrated $2$-or-infinity theorem and, viewed from the symplectic dynamics perspective, settles a particular case of the contact Hofer–Zehnder conjecture. The proofs are based on several auxiliary results of independent interest on the structure of the filtered symplectic homology and the properties of closed orbits.
Key words and phrases:
Closed orbits, Reeb flows, Floer homology, symplectic homology, Franks’ theorem
2020 Mathematics Subject Classification:
53D40, 37J11, 37J46
The work is partially supported by the NSF grants DMS-2304207 (BG) and DMS-2304206 (VG), the Simons Foundation grants 855299 (BG) and MP-TSM-00002529 (VG), and the ERC Starting Grant 851701 via a postdoctoral fellowship (EÇ)
1. Introduction and main results
1.1. Introduction
We study the multiplicity problem for closed characteristics, aka prime closed orbits, of the Reeb flow on the boundary of a star-shaped domain in ${\mathbb{R}}^{2n}$. The first of our two main results, Theorem A, asserts that such a flow has at least $n$ prime closed Reeb orbits whenever the flow is dynamically convex (and the boundary is smooth), improving the previously known lower bound by a factor of two. The lower bound $n$ is obviously sharp.
The second main theorem, Theorem B, is that when, in addition, the domain is centrally symmetric and the Reeb flow is non-degenerate, the flow has either exactly $n$ or infinitely many prime closed orbits. This is a higher-dimensional contact variant of the celebrated Franks’ theorem on periodic points of area preserving diffeomorphisms of $S^{2}$. From the symplectic dynamics perspective, the theorem settles a contact version of the Hofer–Zehnder (HZ) conjecture discussed below.
The proofs are based on several auxiliary results on the structure of the filtered symplectic homology and the properties of closed orbits, which are of independent interest. For instance, we show that this homology is one-dimensional for any background field and any positive action threshold whenever the number of prime closed orbits is finite and the flow is dynamically convex. Below we provide more context for this work.
Broadly understood, the study of closed orbits of Reeb flows on contact type hypersurfaces $M$ in ${\mathbb{R}}^{2n}$, e.g., on convex hypersurfaces, goes back at least to Lyapunov’s times if not earlier. Here we exclusively focus on the case where the hypersurface is the boundary of a star-shaped domain or, equivalently, on Reeb flows on the standard contact sphere $S^{2n-1}$, thus fixing the underlying contact topology. The following conjecture encompasses much of what we know or can expect to be true about prime closed orbits of such flows.
Conjecture (The $n$-or-$\infty$ conjecture).
Let $M$ be the boundary of a star-shaped domain in ${\mathbb{R}}^{2n}$, which we assume to be smooth. Then
  • (C-M)
    Multiplicity: the Reeb flow on $M$ has at least $n$ prime closed orbits;
  • (C-HZ)
    HZ-conjecture: the Reeb flow on $M$ has exactly $n$ prime closed orbits whenever the flow is a Reeb pseudo-rotation, i.e., the number of prime closed orbits is finite.
Here we deliberately broke down the question into two parts, (C-M) and (C-HZ), for our understanding of them and their status are very different. Overall, very little is known about either part of the conjecture in such a generality. For instance, without extra conditions on $M$, it is not even known that there must be more than one prime orbit or more than two if the flow is non-degenerate. (See, however, [AGKM].) The existence of at least one orbit was proved in [Ra] (see also [We78]) and served as the basis of the Weinstein conjecture, [We79], later established in [Vi87] for all contact type hypersurfaces in ${\mathbb{R}}^{2n}$. We expect the $n$-or-$\infty$ conjecture to be true for a very broad class of Reeb flows on $M$, but possibly not all. Once some additional restrictions on $M$ are imposed, usually along the lines of convexity or dynamical convexity or symmetry, much more is known.
In various settings the conjecture is far from new. For instance, Part (C-M), often referred to as the multiplicity problem, is stated in, say, [ALM] in full generality exactly as above. For convex hypersurfaces in ${\mathbb{R}}^{2n}$, it is included in Ekeland’s monograph, [Ek, p. 198], and apparently it was a well-known open question already then. Thus, since convex hypersurfaces are dynamically convex (see [HWZ]), our first result establishes Ekeland’s variant of the conjecture and, in fact, is more general. (Dynamical convexity is the lower bound $n+1$, the same as for ellipsoids, for the Conley–Zehnder index of closed orbits. This condition is in general less restrictive than convexity; see [CE22, CE25, CGHi, DGRZ] and also [AM22, GiM].) For non-degenerate flows, Ekeland’s conjecture was proved in [LZ] (see also [Lo02]) for convex hypersurfaces and in [GK] in the dynamically convex case. Recently, again in the non-degenerate setting, the dynamical convexity condition has been significantly relaxed to a very minor restriction on the index; see [DL2W18, DL2W24, GGM] and references therein.
Without the non-degeneracy requirement, the previously known lower bound on the number of prime closed orbits is only $\lceil n/2\rceil+1$, i.e., roughly $n/2$, and the full form of dynamical convexity is used in the proofs; see [DL, GG20] and also [LZ, Wa16b] for earlier results. As we have already mentioned, in Theorem A we improve this lower bound to $n$, i.e., by roughly a factor of two. (When $n=3$ and $4$, the lower bounds $3$ and, respectively, $4$ have been established in [Wa16a, WHL].)
Imposing the symmetry condition somewhat simplifies the question. In this case, the existence of $n$ prime closed orbits on convex hypersurfaces was proved in [LLZ] without the non-degeneracy requirement and then in [GiM] under a less restrictive dynamical convexity type condition; see also [ALM] for other relevant results. Our second main result, Theorem B, establishes a particular case of Part (C-HZ) and asserts that a Reeb flow with finitely many prime closed orbits must have exactly $n$ such orbits, provided that the flow is non-degenerate and $M$ is centrally symmetric. The lower bound $n$ is known to hold due to either non-degeneracy or, independently, symmetry, and the new point is the upper bound.
To put this theorem and Part (C-HZ) in perspective, it is illuminating to first discuss similar results and questions in the setting of Hamiltonian diffeomorphisms and Hamiltonian Floer theory.
Recall that by the celebrated theorem of Franks a Hamiltonian diffeomorphism of $S^{2}$ with more than two fixed points must have infinitely many periodic orbits; see [Fr92, Fr96, LeC]. A generalization of Franks’ theorem to higher dimensions was conjectured by Hofer and Zehnder in [HZ, p. 263], asserting that a Hamiltonian diffeomorphism of ${\mathbb{C}}{\mathbb{P}}^{n}$ with more than $n+1$ fixed points must have infinitely many periodic orbits. (Hence, the name “HZ-conjecture”.) Here, by Arnold’s conjecture, $n+1$ is the minimal possible number of fixed points, [Fl89b, Fo, FW]. For instance, $n+1=2$ for ${\mathbb{C}}{\mathbb{P}}^{1}=S^{2}$. The HZ-conjecture was proved in [Sh22] under a minor additional non-degeneracy type condition but for a broader class of manifolds than complex projective spaces; see also, e.g., [Al22a, Al22b, Al23, ABF, AL, BX, Ba15, Ba17, Ba18, ÇGG22, GG14, GG16, Gü13, Su21] for other relevant results, conjectures and more context.
Part (C-HZ) is a contact analogue of the HZ-conjecture. To the best of our knowledge, this conjecture was originally stated in [Gü15]. In that paper the first results along the lines of an upper bound on the number of prime closed orbits were proved for lacunary Reeb flows on $S^{2n-1}$. (See also [AM24] for a generalization of this upper bound to other prequantization bundles.) A more recent result closely related to the contact HZ-conjecture is that a non-degenerate dynamically convex Reeb flow with a hyperbolic closed orbit must have infinitely many prime closed orbits, [ÇG2M23]. (We emphasize that that result is logically independent from what is proved in this paper: neither of the results imply the other. The approaches to the question are also substantially different.) Other than these works, we are not aware of any prior results supporting the conjecture when $2n-1\geq 5$. Theorem B proves the first sufficiently general case of the conjecture, albeit under additional conditions.
From a slightly different perspective, the $n$-or-$\infty$ conjecture is a question about Reeb pseudo-rotations. In the context of symplectic dynamics, the two competing definitions of a pseudo-rotation are that this is a Hamiltonian system (a Hamiltonian diffeomorphism or a Reeb flow) with finitely many periodic orbits or, alternatively, with the minimal possible number of periodic orbits (perhaps, counted with multiplicity); see [GG18, Sh22, ÇG2M23]. Hypothetically, the two definitions are equivalent and Theorem B establishes this fact for a certain class of Reeb flows, just as [Sh22] did for Hamiltonian diffeomorphisms of ${\mathbb{C}}{\mathbb{P}}^{n}$ and some other manifolds; see also [AL, BX, ÇGG22]. (Interestingly, this is no longer true for symplectomorphisms, [ABF].) Overall, perhaps expectedly so, there is a considerable similarity between the properties of Hamiltonian and Reeb pseudo-rotations; see, e.g., [AK, LRS] vs. [Ka] or [GG18] vs. [ÇG2M23].
Reeb and Hamiltonian pseudo-rotations are of interest because they may have quite complicated dynamics beyond periodic orbits; for instance, they can be ergodic, [AK, Ka, LRS]. However, in all known examples the periodic orbit data of a pseudo-rotation (the linearized return maps and the actions) is the same as a suitably chosen $S^{1}$-action. (This is consistent with the conjecture that all pseudo-rotations are obtained by the Anosov–Katok conjugation method from [AK, Ka].) In Corollary F we show that, under minor non-resonance conditions, the fixed point data of a dynamically convex Reeb pseudo-rotation matches that of an ellipsoid; cf. [GG22]. It is also worth pointing out that all known examples of Reeb pseudo-rotations on $S^{2n-1}$ are dynamically convex and non-degenerate and actually have the same fixed point data as irrational ellipsoids.
Finally, the case of dimension 2n1=32n-1=3 deserves a separate discussion. According to the so called 2-or-infinity conjecture, a Reeb flow on a closed 3-manifold has either exactly two or infinitely many prime closed orbits. This conjecture is almost proved in general, and lens spaces are the only manifolds admitting Reeb pseudo-rotations; see [CGHu, CGHP, CGH2L, CDR]. Furthermore, the conjecture has been completely proved for $S^{3}$. (See also [GH2M] for a different proof of (C-M) for $S^{3}$.) Hence, both (C-M) and (C-HZ) parts of our $n$-or-$\infty$ conjecture hold in this case. However, one cannot expect this clear-cut dichotomy to hold in higher dimensions, even for tight or fillable contact manifolds. Indeed, while it is reasonable to expect that a majority of such contact manifolds do not admit Reeb pseudo-rotations in all dimensions just as in dimension three, it is not clear how to rigorously state this conjecture and a description of manifolds that do seems beyond our reach. Moreover, the lower bound on the number of orbits clearly depends on the manifold when $n>2$.
Remark 1.1 (Smoothness).
Returning to the statement of Ekeland’s conjecture in [Ek, p. 198], it is worth pointing out that there the hypersurface $M$ is assumed to be only $C^{1}$-smooth. To the best of our understanding, our methods require $M$ to be at least $C^{2}$. We would expect Part (C-M) of the $n$-or-$\infty$ conjecture to be still true in dimension 3 in the $C^{1}$-case and accessible by symplectic topological methods, and possibly false in higher-dimensions; cf. [BHS, LRSV].
The key to the proofs of the two main theorems is Theorem C asserting that for a dynamically convex Reeb pseudo-rotation the (non-equivariant) filtered symplectic homology is one-dimensional for every action threshold and any ground field. The proof of this theorem has three main ingredients. The first one is a new version of the index recurrence theorem (Theorem 3.3), refining the original index recurrence theorem from [GG20, Thm. 5.2] which, in turn, is closely related to and can be derived from various variants of the common jump theorem from [DLW, Lo02, LZ]. This is a combinatorial result about, as the name suggests, recurrent behavior of the Conley–Zehnder index type invariants under iterations. This theorem and its consequences are repeatedly used throughout the paper. The other two ingredients are not new: one is an a priori upper bound on the boundary depth of the symplectic homology persistence module (Theorem 2.3, originally from [GS]) and the other is the Smith inequality for symplectic homology (Theorem 2.5, originally from [Se15, ShZh]). Here, due to the nature of the Smith inequality which is used to obtain a lower bound on the dimension, we first need to work over fields of positive characteristic and then extend the result to all fields by the universal coefficient theorem. Once Theorem C is proved, we can set the ground field to be $\mathbb{Q}$} or any field of zero characteristic.
We use Theorem C to show that every homologically visible (over $\mathbb{Q}$}) closed orbit has one-dimensional equivariant local symplectic homology. This allows to assign such a closed orbit a degree and treat the orbit as if it were non-degenerate. We show that the degrees of such orbits form the increasing sequence $n+1,n+3,\ldots$ and the degrees of bars are $n,n+2,\ldots$. All $\mathbb{Q}$}-visible closed orbits have distinct period (action) and the same ratio of the action and the mean index. (See Corollary D and Theorem E.) In other words, the (non-equivariant) symplectic homology persistence module and the orbit degrees behave as they would for an irrational ellipsoid.
Theorem B is a direct consequence of Corollary D and standard results about periodic orbits on centrally symmetric hypersurfaces. The proof of Theorem A uses again the index recurrence theorem, Theorem E and also several facts about the behavior of the local homology under iterations. The last two components are crucially new ingredients of the proof compared to previous multiplicity results.
The paper is organized as follows. In Section 1.2 we precisely state our main results. Sections 2 and 3 cover preliminary material and auxiliary results, some standard and some not, in a reasonably general setting. In Sections 2.1 and 2.2 we spell out our conventions and notation and then review the notion of a persistence module in the context suitable for our purposes. Equivariant and non-equivariant symplectic homology, filtered and local, and their properties needed for the proofs are discussed in Sections 2.3, 2.4 and 2.5. Most of the facts we recall there are well-known but there are also some new results. The main theme of Section 3 is the Conley–Zehnder index. There we discuss several Conley–Zehnder index type invariants and their properties, briefly touch upon dynamical convexity, and then state a new version of the index recurrence theorem and some of its consequences. This is the combinatorial backbone of the proof. In Section 4 we prove the main results of the paper. Finally, the index recurrence theorem is proved in Section 5.
1.2. Main results
Before stating the results, we need to briefly review some terminology, referring the reader to Sections 2 and 3.1 for more details.
Let $M$ be the boundary of a star-shaped domain $W\subset{\mathbb{R}}^{2n}$. Throughout the paper, we assume $M$ to be smooth. We say that a closed Reeb orbit $y$ in $M$ is the $k$th iterate of a closed orbit $x$ and write $y=x^{k}$ if $y(t)=x(kt)$ up to the choice of initial conditions on $x$ and $y$. (In what follows we will always identify the closed orbits $t\mapsto x(t)$ and $t\mapsto x(t+\theta)$ for all $\theta\in{\mathbb{R}}$.) A closed orbit $x$ is prime if it is not an iterate of any other closed orbit.
A closed Reeb orbit $x$ is said to be non-degenerate if $1$ is not an eigenvalue of its Poincaré return map. The Reeb flow on $M$ is non-degenerate if all closed orbits, prime and iterated, are non-degenerate. The parity of a non-degenerate orbit is the parity of its Conley–Zehnder index $\mu$; see Section 3.1.3.
When $x$ is non-degenerate, we call it non-alternating if the number, counted with algebraic multiplicity, of real negative eigenvalues in (1,0)(-1,0) of its Poincaré return map is even. Otherwise, we call $x$ alternating. A non-degenerate even iterate of an alternating closed orbit is called bad; all other non-degenerate closed orbits are called good. When $M$ is centrally symmetric, we call $x$ symmetric if x(t)=x(t+T/2)-x(t)=x(t+T/2), where $T$ is the period of $x$, i.e., $x$ and x-x are the same orbit up to the choice of initial condition.
Fix a ground field $\mathbb{F}$}. (The choice of $\mathbb{F}$} is essential in several instances in this paper.) A closed Reeb orbit $x$ is said to be $\mathbb{F}$}-visible if its local symplectic homology $\operatorname{SH}(x;{\mathbb{F}})$ over $\mathbb{F}$} is non-zero; see Section 2.4. For instance, a non-degenerate closed orbit $x$ is $\mathbb{Q}$}-visible if and only if $x$ is good. We denote the set of $\mathbb{F}$}-visible closed orbits by ${\mathcal{P}}({\mathbb{F}})$ or simply $\mathcal{P}$} when $\mathbb{F}$} is clear from the context.
Finally, recall that the Reeb flow on $M^{2n-1}$ is said to be dynamically convex if $\mu_{-}(x)\geq n+1$ for all closed Reeb orbits $x$, where $\mu_{-}$ is the lower semi-continuous extension of the Conley–Zehnder index $\mu$; see Section 3.1.
The following four theorems and their corollaries are the key results of the paper.
Theorem A (Multiplicity).
Assume that the Reeb flow on the boundary $M^{2n-1}\subset{\mathbb{R}}^{2n}$ of a star-shaped domain is dynamically convex and has finitely many prime closed orbits, i.e., the flow is a Reeb pseudo-rotation.
  • (M1)
    Then the flow has at least $n$ prime closed orbits.
  • (M2)
    If, in addition, the flow is non-degenerate, it has exactly $n$ non-alternating prime closed orbits, and all orbits have parity $n+1$.
As a consequence, the Reeb flow on the boundary of a star-shaped domain has at least $n$ prime closed orbits whenever the flow is dynamically convex.
This theorem is proved in Section 4.1, where in fact we establish a more precise result (Theorem 4.1) identifying $n$ prime orbits meeting a certain visibility criterion. The key new point of Theorem A is, of course, (M1), establishing Ekeland’s conjecture and improving previously known multiplicity lower bounds by roughly a factor of two. As we have already mentioned, the existence of at least $n$ non-alternating orbits for non-degenerate dynamically convex Reeb flows is well-known; see [GK, LZ]. The new part in (M2), essential for the proof of Theorem B, is that in this case the number of non-alternating closed prime orbits is exactly $n$ and the remaining prime orbits, if they exist, must all be alternating.
Our second main result establishes a particular case of the contact HZ-conjecture.
Theorem B (HZ-conjecture).
Assume that the Reeb flow on the boundary $M^{2n-1}\subset{\mathbb{R}}^{2n}$ of a centrally symmetric star-shaped domain has finitely many prime closed orbits and is dynamically convex and non-degenerate. Then the flow has exactly $n$ prime closed orbits. These orbits are symmetric and non-alternating.
This theorem, also proved in Section 4.1, is in fact an easy consequence of (M2) and Corollary D below. Recall in this connection that [ÇG2M23, Thm. A] asserts, in particular, that a dynamically convex Reeb flow on $M$ with a hyperbolic closed orbit has infinitely many prime closed orbits. Interestingly, that theorem does not follow from Theorem A even in the non-degenerate case. Nor is it a consequence of Theorem B when in addition $M$ is symmetric.
Remark 1.2.
In all results of this paper, the requirement that the flow has only finitely many prime closed orbits can be replaced by a hypothetically less restrictive condition that the number of prime eventually homologically visible orbits is finite. Here we call $x$ eventually homologically visible if $\operatorname{SH}\big(x^{k};{\mathbb{F}})\neq 0$ for some $k$ and $\mathbb{F}$}. Furthermore, when ${\mathit{c}har}\,{\mathbb{F}}>0$ in Theorem C below it suffices to assume that the number of prime eventually $\mathbb{F}$}-visible orbits is finite, where $x$ is eventually $\mathbb{F}$}-visible if $\operatorname{SH}\big(x^{k};{\mathbb{F}})\neq 0$ for some $k$ and that particular ground field $\mathbb{F}$}.
Turning to our more technical results, fix a background field $\mathbb{F}$}. When $\mathbb{F}$} has positive characteristic, ${\mathit{c}har}\,{\mathbb{F}}>0$, the reader may take ${\mathbb{F}}={\mathbb{Z}}_{p}$ and when ${\mathit{c}har}\,{\mathbb{F}}=0$ it would be sufficient, without loss of generality, to set $\mathbb{F}$$\mathbb{Q}$}. Denote by $\operatorname{SH}^{t}(W;{\mathbb{F}})$, where the $t\in{\mathbb{R}}$ is the action or equivalently period threshold, the filtered symplectic homology of $W$ with coefficients in $\mathbb{F}$}. For each $t$, this is a $\mathbb{Z}$}-graded vector space over $\mathbb{F}$}. Throughout the paper it is convenient to treat the family of vector spaces $\operatorname{SH}^{t}(W;{\mathbb{F}})$ as a graded persistence module, which we denote by $\operatorname{SH}(W;{\mathbb{F}})$; see Sections 2.2 and 2.3. The next result, proved in Section 4.3, is central to the proofs of Theorems A and B.
Theorem C.
Fix a ground field $\mathbb{F}$}. Assume that the Reeb flow on the boundary $M^{2n-1}$ of a star-shaped domain $W\subset{\mathbb{R}}^{2n}$ is a dynamically convex pseudo-rotation, i.e., it has finitely many prime closed orbits and is dynamically convex. Then, for every $t>0$,
$\dim\operatorname{SH}^{t}(W;{\mathbb{F}})=1.$ (1.1)
In other words, every $t>0$ belongs to exactly one bar in the barcode of the persistence module $\operatorname{SH}(W;{\mathbb{F}})$.
The support of a $\mathbb{Z}$}-graded vector space is the set of degrees in which the vector space is non-zero. Denote by $\operatorname{CH}(x;{\mathbb{F}})$ the equivariant local symplectic homology of $x$ with coefficients in $\mathbb{F}$}; see Section 2.4. An easy consequence of Theorem C is the following corollary also proved in Section 4.3.
Corollary D.
Let $W$ be as in Theorem C. Then
  • all $\mathbb{F}$}-visible closed orbits $x$ have distinct periods and $\dim_{\mathbb{F}}\operatorname{SH}(x;{\mathbb{F}})=2$;
  • the homology $\operatorname{SH}(x;{\mathbb{Q}})$ is supported in two consecutive degrees for $x\in{\mathcal{P}}({\mathbb{Q}})$, i.e.,
    $\operatorname{supp}\operatorname{SH}(x;{\mathbb{Q}})=:\{h,h+1\}$ (1.2)
    for some $h\in{\mathbb{Z}}$, and $\operatorname{CH}_{i}(x;{\mathbb{Q}})={\mathbb{Q}}$ when $i=h$ and zero otherwise.
From now on, we can restrict our attention to ground fields of zero characteristic, e.g., $\mathbb{F}$$\mathbb{Q}$}. In what follows, we call $h$ from (1.2) the degree $\deg(x)$ of $x$. In other words, $h=\deg(x)\in{\mathbb{Z}}$ is the only integer such that $\operatorname{CH}_{h}(x;{\mathbb{Q}})\neq 0$ (and hence is $\mathbb{Q}$}.) For instance, $\deg(x)=\mu(x)$ when $x$ is non-degenerate and good. We emphasize that this notion is specific to the setting of Theorem C and Corollary D, i.e., $\deg(x)$ is defined only when the flow is dynamically convex and has finitely many prime closed Reeb orbits and $x$ is $\mathbb{Q}$}-visible. Recall that ${\mathcal{P}}={\mathcal{P}}({\mathbb{Q}})$ stands for the set of all, not necessarily prime, $\mathbb{Q}$}-visible closed orbits. Denote by ${\mathcal{A}}(x)$ the action (aka the period) of $x$ and by $\operatorname{\hat{\mu}}(x)$ the mean index of $x$; see Sections 2.1 and 3.1.
Our next theorem, proved in Section 4.2, provides yet more detailed information on closed Reeb orbits in $M$ and the persistence module $\operatorname{SH}(W;{\mathbb{Q}})$ and also serves as a bridge between Theorems A and C.
Theorem E.
Assume that $W$ is as in Theorem C. Then $\mathcal{P}$} and the barcode $\mathcal{B}$} of the persistence module $\operatorname{SH}(W;{\mathbb{Q}})$ have the following properties:
  • (O1)
    The map $\deg\colon{\mathcal{P}}\to{\mathbb{Z}}$ is a bijection onto $n-1+2{\mathbb{N}}=\{n+1,n+3,\ldots\}$ and orbits with larger degree have larger period. Likewise, $\deg\colon{\mathcal{B}}\to{\mathbb{Z}}$ is a bijection onto $n-2+2{\mathbb{N}}=\{n,n+2,\ldots\}$ and the bars with larger degree have larger period (at, say, the beginning point).
  • (O2)
    All closed orbits $x\in{\mathcal{P}}$ have the same ratio ${\mathcal{A}}(x)/\operatorname{\hat{\mu}}(x)$.
This theorem asserts, in particular, that the graded barcode of $\operatorname{SH}(W;{\mathbb{Q}})$ looks exactly like that of an irrational ellipsoid: the bars form an ascending infinite staircase with one riser of height one (the degree) for each closed orbit and starting with a tread of degree $n$ beginning at the interior of $W$. We will show in the next section that the similarity with ellipsoids goes farther. Note also that as a consequence of (O1) the flow is automatically lacunary in the terminology from [AM24]. It is also worth noting that the ratios ${\mathcal{A}}(x)/\operatorname{\hat{\mu}}(x)$ playing an important role in these questions were originally, to the best of our knowledge, considered in [Ek, EH].
Remark 1.3.
We conjecture that (1.1) is equivalent to that the Reeb flow is dynamically convex and has only a finite number of prime orbits at least in the non-degenerate case. It not hard to see that then (1.1) implies dynamical convexity. However, proving finiteness appears to be much more difficult. Without non-degeneracy one would also have to account for invisible orbits.
1.3. Application: ellipsoid comparison
The main results of the paper have an interesting application in the context of the question of the extent to which symplectic topology of $W$ or of its interior is tied to the dynamics on $M=\partial W$. Assume that $W\subset{\mathbb{R}}^{2n}$ is star-shaped and the flow on $M=\partial W$ is non-degenerate, dynamically convex and has finitely many prime closed orbits. We further require that the number of such orbits is exactly $n$ and these orbits are non-alternating. For instance, this is the case when $W$ is symmetric as in Theorem B. In this section, in the spirit of [GG22], we show that then, under a very minor non-resonance condition, the linearized flows along the prime closed Reeb orbits on $M=\partial W$ are completely determined up to conjugation by the persistence module $\operatorname{SH}(W;{\mathbb{F}})$ and, moreover, these linearized flows are exactly the same as those for an ellipsoid with the same action spectrum as $M$.
To be more specific, denote by $\vec{\Delta}:=\{\Delta_{1}<\cdots<\Delta_{n}\}\in{\mathbb{R}}^{n}_{+}$ the ordered set of the mean indices of these orbits. (By Theorem E, all closed orbits have distinct mean indices. In what follows, we will always order prime orbits so that their mean indices form an increasing sequence.) Consider the solid ellipsoid $W_{\vec{\Delta}}\subset{\mathbb{C}}^{n}={\mathbb{R}}^{2n}$ given by
$\sum_{j}\frac{|z_{j}|^{2}}{\Delta_{j}}\leq 1$
and its boundary $M_{\vec{\Delta}}=\partial W_{\vec{\Delta}}$. The prime closed orbits on $M_{\vec{\Delta}}$ also have mean indices $\Delta_{1}<\cdots<\Delta_{n}$ and up to scaling $M$ and $M_{\vec{\Delta}}$ have the same action spectra. Indeed, let us scale $M$ so that $\operatorname{\hat{\mu}}(x)={\mathcal{A}}(x)$ for all closed orbits $x$ on $M$. This is possible by (O2). Then for both $M$ and $M_{\vec{\Delta}}$ the action spectrum is simply the union of sequences $\Delta_{j}{\mathbb{N}}$.
Corollary F.
Let $W\subset{\mathbb{R}}^{2n}$ be a star-shaped domain. Assume that the flow on $M=\partial W$ is non-degenerate, dynamically convex, has exactly $n$ prime closed orbits. In addition, let us require the following non-resonance condition: for every $j$ all ratios $\pm\Delta_{j}/\Delta_{i}$, $i\neq j$, are distinct modulo $\mathbb{Z}$}. Then the linearized flows along the corresponding closed orbits in $M$ and $M_{\vec{\Delta}}$ are equal up to conjugation as elements of $\operatorname{\widetilde{Sp}}\big(2(n-1)\big)$.
This corollary, proved in Section 4.4, asserts that within the admittedly very narrow class of hypersurfaces $M$ meeting the conditions of the corollary, the action spectrum completely determines the linearized flows along prime closed orbits and, of course, the actions. In particular, all closed Reeb orbits in $M$ are elliptic. At the same time, dynamics of the Reeb flow on $M$, beyond closed orbits, can be radically different from that on $M_{\vec{\Delta}}$ and quite non-trivial. For instance, the Reeb flow on $M$ can be ergodic even in the setting of Corollary F; see [Ka]. Clearly, then the closed domains $W$ and $W_{\vec{\Delta}}$ are not symplectomorphic and such dynamics features are not detected by their symplectic homology persistence modules, which are isomorphic. We do not know if the interiors of $W$ and $W_{\vec{\Delta}}$ are symplectomorphic or not. (The answer may depend on Liouville/Diophantine properties of $\vec{\Delta}$.)
Acknowledgments
The authors would like to thank Leonardo Macarini, Marco Mazzucchelli and Felix Schlenk for useful remarks and suggestions.
2. Symplectic homology
2.1. Conventions and notation
Let $\alpha$ be the contact form on the boundary $M=\partial W$ of a Liouville domain $W^{2n\geq 4}$. We will also use the same notation $\alpha$ for a primitive of the symplectic form $\omega$ on $W$. The grading of Floer or symplectic homology is essential for our purposes and for the sake of simplicity we require that $c_{1}(TW)=0$. (This condition can be relaxed.) As usual, denote by $\widehat{W}$ the symplectic completion of $W$, i.e.,
$\widehat{W}=W\cup_{M}M\times[1,\,\infty)$
with the symplectic form $\omega=d\alpha$ extended from $W$ to $M\times[1,\infty)$ as
$\omega:=d(r\alpha),$
where $r$ is the coordinate on $1,\,\infty)$. Sometimes it is convenient to extend the function $r$ to a collar of $M=\partial W$ in $W$. Thus we can think of $\widehat{W}$ as the union of $W$ and $M\times[1-\epsilon,\,\infty)$ for small $\epsilon>0$ with $M\times[1-\epsilon,\,1]$ lying in $W$ and the symplectic form given by the same formula.
The Reeb vector field $R$ of $\alpha$ on $M$ is determined by the conditions $i_{R}d\alpha=0$ and $\alpha(R)=1$, and the flow of $R$ is called the Reeb flow. The period or action of a closed orbit $x$ of the flow, denoted by ${\mathcal{A}}(x)$, is the integral of $\alpha$ over $x$. The action spectrum ${\mathcal{S}}(\alpha)$ is the collection of periods of all closed Reeb orbits.
Example 2.1.
The main example we are interested in is where $W$ is a star-shaped domain in ${\mathbb{R}}^{2n}$ with smooth boundary. Then we have a natural identification $\widehat{W}={\mathbb{R}}^{2n}$ as symplectic manifolds.
As we have already mentioned in Section 1.2, a closed Reeb orbit $x$ is non-degenerate if all eigenvalues of the Poincaré return map are different from 1, and $x$ is strongly non-degenerate if all iterates $x^{k}$, $k\in{\mathbb{N}}$, are non-degenerate. The Reeb flow or the form $\alpha$ is said to be non-degenerate if all closed orbits are non-degenerate. For a closed orbit $x$, we denote by $x^{k}$, $k\in{\mathbb{N}}$, its $k$th iterate. This is simply the orbit $x$ traversed $k$ times. Clearly, ${\mathcal{A}}\big(x^{k}\big)=k{\mathcal{A}}(x)$. An orbit is said to be prime if it is not an iterate of any other orbit. Otherwise, we will call an orbit iterated.
When $x$ is non-degenerate, we denote by $\mu(x)$ the Conley–Zehnder index of $x$. Without the non-degeneracy condition, we have the lower and upper semicontinuous extensions $\operatorname{\hat{\mu}}_{\pm}(x)$ of the Conley–Zehnder index and the mean index $\operatorname{\hat{\mu}}(x)$ associated to $x$. Referring the reader to Section 3.1.1 for the definitions and a discussion, we only note here that these indices depend on an additional piece of data (not unique in general) which determines the choice of a symplectic trivialization of the contact structure along $x$; see Section 3.1.3.
Recall also that a non-degenerate closed orbit $x$ is said to be non-alternating if the number (counted with algebraic multiplicity) of eigenvalues in (1,0)(-1,0) of the return map is even. Otherwise, we call $x$ alternating. When $x$ is strongly non-degenerate, these conditions are equivalent to that all indices $\mu\big(x^{k}\big)$, $k\in{\mathbb{N}}$, have the same parity or that the parity alternates. (A word is due on terminology: Non-alternating orbits are sometimes referred to as negative hyperbolic, but we find this term misleading when $\dim M>3$. The frequently used terms “even/odd” are more intuitive; however they clash with the parity of an orbit.) While formally speaking these definitions require no conditions on $x$, we will apply them only when $x$ is prime and strongly non-degenerate. The notion of (non-)alternating closed orbit is not particularly useful when $x^{k}$ is allowed to be degenerate; see Example 2.13. The parity of a closed non-degenerate orbit $x$ is simply the parity of $\mu(x)$. A non-degenerate closed orbit $x$ is said to be bad if it is an even iterate of an alternating prime orbit, and good otherwise.
Throughout the paper we will impose the following additional condition on the Reeb flow of $\alpha$ on $M$:
  • (BA)
    Background assumption: All closed orbits are isolated, i.e., for every $t\in{\mathbb{R}}$ there are only finitely many closed orbits with period less than $t$.
This condition is obviously satisfied in two cases we are interested in here: when $\alpha$ is non-degenerate and when the Reeb flow has only finitely many prime closed orbits. The role of this condition is rather expository than conceptual, and most of the discussion in the next two sections carries over to the general case with only technical changes. For instance, the symplectic homology is still a persistence module (see Section 2.2) when $\mathcal{S}$} is only closed and nowhere dense and $V^{s}$ might be infinite-dimensional, once (PM2) in the definition of a persistence module is suitably modified, [ÇG2M25].
We implicitly adopt the conventions from [ÇG2M23, ÇG2M25, GG20] on the almost complex structures, Hamiltonians and the Floer equation. For the sake of brevity, since these conventions are never explicitly used in the paper, we leave them out. Here we only mention that with these conventions, the Conley–Zehnder index of a non-degenerate minimum with small Hessian in dimension $2n$ is $n$.
The characteristic of the ground field $\mathbb{F}$} is essential in some parts of the paper. When the characteristic is zero, we can and usually will take $\mathbb{F}$$\mathbb{Q}$}. We use the notation ${\mathbb{F}}_{p}$ for the positive characteristic field ${\mathbb{Z}}/p{\mathbb{Z}}$, where $p$ is prime, or any other field of characteristic $p$.
2.2. Graded persistence modules
While one can easily avoid using persistence modules in the proofs of our main results, the language comes handy. In this section we briefly recall relevant definitions in the form and generality suitable for our purposes. We refer the reader to [PRSZ] for a general introduction to persistence modules, their applications in geometry and analysis and further references.
Fix a field $\mathbb{F}$}, which we will suppress in the notation. A (graded) persistence module $(V,\pi)$ is a family of $\mathbb{Z}$}-graded vector spaces $V^{s}$ over $\mathbb{F}$} parametrized by $s\in{\mathbb{R}}$ together with a functorial family $\pi$ of structure maps. These are $\mathbb{Z}$}-graded linear maps $\pi_{st}\colon V^{s}\to V^{t}$, where $s\leq t$, and functoriality is understood as that $\pi_{sr}=\pi_{tr}\pi_{st}$ whenever $s\leq t\leq r$ and $\pi_{ss}={\mathit{i}d}$. In what follows, we often suppress $\pi$ in the notation and simply refer to $(V,\pi)$ as $V$. In such a general form the concept is not particularly useful and one usually imposes additional conditions on $(V,\pi)$. These conditions vary depending on the context. Below we spell out the setting most suitable for our purposes.
Namely, we require that there is a closed, discrete, bounded from below subset $\mathcal{S}$$\subset{\mathbb{R}}$, which we call the spectrum of $V$, and the following four conditions are met:
  • (PM1)
    The persistence module $V$ is locally constant outside $\mathcal{S}$}, i.e., $\pi_{st}$ is an isomorphism whenever $s\leq t$ are in the same connected component of ${\mathbb{R}}\setminus{\mathcal{S}}$.
  • (PM2)
    In every degree $m$, the dimension of $V^{s}$ is finite for every $s$: $\dim V^{s}_{m}<\infty$.
  • (PM3)
    Left-semicontinuity: For all $t\in{\mathbb{R}}$,
[…]
  • (PM4)
    Lower bound: $V^{s}=0$ when $s<s_{0}$ for some $s_{0}\in{\mathbb{R}}$. (Throughout the paper we will assume that $s_{0}=0$.)
In the setting we are interested in, $V^{s}$ is naturally defined only for $s\not\in{\mathcal{S}}$; then the definition is extended to all $s\in{\mathbb{R}}$ by (2.1). By (PM4), we can always assume that $s_{0}\leq\inf{\mathcal{S}}$, which essentially forces $\mathcal{S}$} to be bounded from below. We emphasize, however, that $\mathcal{S}$} is not assumed to be bounded from above and this is actually never the case in our setting. It will be convenient to set
[…]
A basic example motivating requirements (PM1)(PM4) is that of the sublevel homology of a smooth function.
Example 2.2 (Homology of sublevels).
Let $M$ be a smooth manifold and $f\colon M\to{\mathbb{R}}$ be a proper smooth function bounded from below and having a discrete set of critical values. Set $V^{s}:=H_{*}\big(\{f<s\};{\mathbb{F}}\big)$ with the structure maps induced by inclusions. It is not hard to see that conditions (PM1)(PM4) are met with $\mathcal{S}$} being the set of critical values of $f$.
By definition, $\operatorname{supp}V^{s}$, the support of $V^{s}$, is the set of degrees $m$ such that $V_{m}^{s}\neq 0$.
Recall furthermore that a (graded) interval persistence module ${\mathbb{F}}_{(a,\,b]}$, where $\infty<a<b\leq\infty$, is defined by setting
$V^{s}_{m}:=\begin{cases}{\mathbb{F}}&\text{ when }s\in(a,\,b],\\ 0&\text{ when }s\not\in(a,\,b],\end{cases}$
in one degree $m$ and $V^{s}_{l}=0$ in all other degrees $l\neq m$. Furthermore, in degree $m$ we have $\pi_{st}={\mathit{i}d}$ if $a<s\leq t\leq b$ and $\pi_{st}=0$ otherwise. The interval modules are exactly the simple persistence modules, i.e., the persistence modules that are not a (non-trivial) direct sum of other persistence modules.
The normal form or structure theorem asserts that every persistence module meeting the above conditions can be decomposed as a direct sum of a countable collection (i.e., a countable multiset) of interval persistence modules. Moreover, this decomposition is unique up to re-ordering of the sum. We refer the reader to [PRSZ] for a proof of this theorem and further references, and also, e.g., to [CZCG, ZC] for previous or related results. The multiset ${\mathcal{B}}(V)$ of intervals entering this decomposition is referred to as the barcode of $V$ and the intervals occurring in ${\mathcal{B}}(V)$ as bars. In addition, every bar is assigned a degree which is by definition the degree of the component of $V^{s}$ it comes from.
Finally, for a graded persistence module, just as for graded vector spaces, we define the grading shifted persistence module as
$(V[n])_{m}:=V_{m+n}.$
Thus, when $n\geq 0$, the grading of $V$ gets shifted by $n$ to the left.
2.3. Filtered symplectic homology
The first versions of symplectic homology were introduced in [CFH, Vi99] and since then the theory has been further developed. We refer the reader to, e.g., [Abu, Bo, BO09a, BO09b, BO13, Se06] for the definitions and basic constructions, which we mainly omit or just informally sketch for the sake of brevity, focusing on lesser known facts. Here, treating equivariant and non-equivariant symplectic homology as persistence modules, we closely follow [ÇG2M25, GG20].
Let the Liouville domain $W$ be as in Section 2.1. Fixing a field $\mathbb{F}$}, we denote the non-equivariant filtered symplectic homology over $\mathbb{F}$} for the action interval $I\subset{\mathbb{R}}$ by $\operatorname{SH}^{I}(W;{\mathbb{F}})$. This is a vector space over $\mathbb{F}$}, graded by the Conley–Zehnder index. Furthermore, we set
$\operatorname{SH}^{t}(W;{\mathbb{F}}):=\operatorname{SH}^{(-\infty,t)}(W;{\mathbb{F}}).$
This family of vector spaces together with the natural “inclusion” maps form a graded persistence module in the sense of Section 2.2 with ${\mathcal{S}}={\mathcal{S}}(\alpha)\cup\{0\}$. (It is essential at this point that $W$ meets the background requirement (BA). Otherwise, we would need to modify the definition of a persistence module; see [ÇG2M25].) We denote the resulting persistence module by $\operatorname{SH}(W;{\mathbb{F}})$. The global symplectic homology is then $\operatorname{SH}^{\infty}(W;{\mathbb{F}})$.
Somewhat informally, $\operatorname{SH}^{t}(W;{\mathbb{F}})$ can be defined via the following construction. As is pointed out in Section 2.2, it suffices to define $\operatorname{SH}^{t}(W;{\mathbb{F}})$ when $t\not\in{\mathcal{S}}$. Assume first that the Reeb flow is non-degenerate. Let $H$ be an autonomous Hamiltonian on $\widehat{W}$, which is a convex, strictly increasing function $h(r)$ of $r$ on $M\times[1,\infty)$ and 0 on $W$. The function $h$ is required to be linear at infinity, i.e., of the form trctr-c with $c>0$. Furthermore, let $\tilde{H}$ be a small Morse perturbation of $H$ supported near $W$. Then $\tilde{H}$ is a Morse–Bott Hamiltonian in the sense of [Bo, BO09a], and $\operatorname{SH}^{t}(W;{\mathbb{F}})$ is simply the Floer homology of $H$ or equivalently $\tilde{H}$; cf. [ÇG2M25, Thm. 3.5 and Cor. 3.6]. This is the homology of a complex with two generators $\hat{x}$ and $\check{x}$ of degrees, $\mu(x)+1$ and $\mu(x)$ respectively, for every closed Reeb of orbit $x$ with ${\mathcal{A}}(x)<t$ and one generator for every critical point of $\tilde{H}$; see, e.g., [Bo, BO09a, ÇG2M23]. In fact, one can trim down this complex further and make the number of generators of the latter type equal to $\dim\operatorname{H}(W;{\mathbb{F}})$. In any event, when $W$ is star-shaped we can have only one such generator, which we denote by $[W]$. When the Reeb flow is degenerate, we first take a small non-degenerate perturbation of the contact form and then apply this construction. For an interval $I=(t^{\prime},t)$ we define $\operatorname{SH}^{I}(W;\alpha)$ as the quotient of $\operatorname{SH}^{t}(W;{\mathbb{F}})$ by the image of $\operatorname{SH}^{t^{\prime}}(W;{\mathbb{F}})$.
As a consequence of this construction and (BA), for every $t$ the support
$\operatorname{supp}\operatorname{SH}^{t}(W;{\mathbb{F}}):=\{m\in{\mathbb{Z}}\mid\operatorname{SH}^{t}_{m}(W;{\mathbb{F}})\neq 0\}$
is a finite set. Furthermore, when $\delta>0$ is sufficiently small, e.g., $\delta<\inf{\mathcal{S}}(\alpha)$,
$\operatorname{SH}^{\delta}(W;{\mathbb{F}}):=\operatorname{H}(W;{\mathbb{F}})[-n];$ (2.2)
see, e.g., [BO09b, Vi99]. It readily follows from the above description of the symplectic homology as the Floer homology of $\tilde{H}$ that the Euler characteristic of the graded vector space $\operatorname{SH}^{t}(W;{\mathbb{F}})$ is independent of $t>0$. To be more precise, for any field $\mathbb{F}$} and every $t>0$, we have
$\sum_{m}(-1)^{m}\dim\operatorname{SH}^{t}_{m}(W;{\mathbb{F}})=(-1)^{n}\chi(W).$ (2.3)
Less directly, one can also use (2.11), (2.19) and (2.20) to prove (2.3).
The construction of symplectic homology carries over to $\mathbb{Z}$} or any other coefficient ring (see, e.g., [Abu]), and hence we have the universal coefficient theorem. In particular, for a fixed $t$, in every degree $m$ we have
$\dim\operatorname{SH}^{t}_{m}(W;{\mathbb{Q}})=\dim\operatorname{SH}^{t}_{m}(W;{\mathbb{F}}_{p}),$ (2.4)
when the characteristic $p$ is large enough (depending on $t$). Indeed, the torsion part of the group $\operatorname{SH}^{t}_{m}(W;{\mathbb{Z}})$ is finite and (2.4) holds whenever $\operatorname{SH}^{t}_{m}(W;{\mathbb{Z}})$ has no elements of order $p$.
In this paper we will also use two less standard properties of symplectic homology. The first of these is [GS, Prop. 3.5], which is essentially due to K. Irie, combined with [GS, Thm. 3.8].
Theorem 2.3 (Boundary depth upper bound, [GS]).
Assume that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})=0$. Then all bars in the barcode of $\operatorname{SH}(W;{\mathbb{F}})$ are bounded by some constant $C_{\scriptscriptstyle{bar}}$ independent of the location of the bar. When $W$ is displaceable in $\widehat{W}$, the displacement energy can be taken as $C_{\scriptscriptstyle{bar}}$, which is then independent of $\mathbb{F}$}.
Remark 2.4.
Recall in this connection that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})$ vanishes, for instance, whenever $W$ is displaceable in $\widehat{W}$. For example, $\operatorname{SH}(W)^{\infty}=0$ when $W$ is a star-shaped domain in $\widehat{W}={\mathbb{R}}^{2n}$. In this case, vanishing of $\operatorname{SH}^{\infty}(W)$ is established in [Vi99]. The general case is proved in [CFO] via Rabinowitz–Floer homology and direct proofs are given in [Su16] and [GS]. Strictly speaking, Theorem 2.3 is proved in [GS] for $\mathbb{F}$$\mathbb{Q}$}, but the argument carries over word-for-word to any unital ring.
The second fact we will need is a variant of the Smith inequality or Borel’s localization for filtered Floer homology.
Theorem 2.5 (Smith Inequality, [Se15, ShZh]).
For every Liouville domain $W$ and any interval $I$ with end-points outside ${\mathcal{S}}(\alpha)$,
$\dim\operatorname{SH}^{pI}(W;{\mathbb{F}}_{p})\geq\dim\operatorname{SH}^{I}(W;{\mathbb{F}}_{p})$ (2.5)
and, in particular,
$\dim\operatorname{SH}^{pt}(W;{\mathbb{F}}_{p})\geq\dim\operatorname{SH}^{t}(W;{\mathbb{F}}_{p})$ (2.6)
for all $t>0$.
These inequalities were originally proved for the Floer homology of any Hamiltonian $H$ on $\widehat{W}$ of the form H=trcH=tr-c outside a compact set with $t\not\in{\mathcal{S}}(\alpha)$, [Se15, ShZh]. The version stated here readily follows from those results combined with the fact that for any fixed interval $I$ with end-points outside ${\mathcal{S}}(\alpha)$ and a fixed prime $p$, we have $\operatorname{SH}^{I}(W;{\mathbb{F}}_{p})=\operatorname{HF}^{I}(H;{\mathbb{F}}_{p})$ for a suitable choice of $H$; see [ÇG2M23, Sect. 3]. For instance, for $I=(-\infty,t)$ we can take any $H$ such that H=trcH=tr-c at infinity. Then, $\operatorname{SH}^{I}(W;{\mathbb{F}}_{p})=\operatorname{HF}(H;{\mathbb{F}}_{p})$ by [ÇG2M25, Cor. 3.7], which implies (2.6).
Filtered symplectic homology has an $S^{1}$-equivariant counterpart, which we denote by $\operatorname{CH}^{I}(W;{\mathbb{F}})$. (In fact, the construction carries over to any ring and, in particular, $\mathbb{Z}$}.) Again we refer the reader to, e.g., [BO13, Vi99] for definitions; here we closely follow [GG20]. As above, let us set
$\operatorname{CH}^{t}(W;{\mathbb{F}}):=\operatorname{CH}^{(-\infty,t)}(W;{\mathbb{F}}),$
denoting the resulting persistence module by $\operatorname{CH}(W;{\mathbb{F}})$, and also
$\operatorname{CH}^{+}(W;{\mathbb{F}}):=\operatorname{CH}^{(\delta,\infty)}(W;{\mathbb{F}}),$
where $\delta>0$ is sufficiently small. (Namely, $\delta<\inf{\mathcal{S}}(\alpha)$.) Then, when $W$ is a star-shaped domain,
$\dim\operatorname{CH}_{m}^{+}(W;{\mathbb{F}})=\begin{cases}1\textrm{ when }m=n+1+2i,\,i\geq 0,\\ 0\textrm{ otherwise.}\end{cases}$ (2.7)
When the Reeb flow is non-degenerate, for any interval $I\subset(0,\infty)$ with end-points outside ${\mathcal{S}}(\alpha)$, the space $\operatorname{CH}^{I}(W;{\mathbb{F}})$ is the homology of a certain complex generated by closed Reeb orbits with action in $I$ and graded by the Conley–Zehnder index; see Remark 2.14 and [GG20, Sec. 2.5]. (If $\alpha$ is degenerate and the end-points of $I$ are outside ${\mathcal{S}}(\alpha)$, we can first take a small non-degenerate perturbation of $\alpha$ and then apply this construction.)
Remark 2.6.
It is worth keeping in mind that this description does not carry over to a field of positive characteristic $p\geq 2$. The reason is that $x^{p}$ can contribute infinitely many generators to the complex over ${\mathbb{F}}_{p}$. For instance, if $x^{p}$ is non-degenerate and the only orbit with action in $I$ and $x$ is prime, $\operatorname{CH}^{I}(W;{\mathbb{F}}_{p})$ is isomorphic to the homology of the infinite-dimensional lens space $S^{2\infty-1}/{\mathbb{Z}}_{p}$ over ${\mathbb{F}}_{p}$.
Finally, recall from [BO13] that the equivariant and non-equivariant symplectic homology of $W$ fit together into the Gysin exact sequence
$\cdots\to\operatorname{SH}_{m}^{I}(W;{\mathbb{F}})\to\operatorname{CH}^{I}_{m}(W;{\mathbb{F}})\stackrel{{\scriptstyle{\mathcal{D}}}}{{\longrightarrow}}\operatorname{CH}^{I}_{m-2}(W;{\mathbb{F}})\to\operatorname{SH}^{I}_{m-1}(W;{\mathbb{F}})\to\cdots.$
While the construction of symplectic homology in [BO13] uses $\mathbb{Q}$} as the coefficient field and a semi-infinite interval as $I$, it carries over word-for-word to $\mathbb{Z}$}, and hence any ring, and an arbitrary interval $I$; see, e.g., [GG20, Sect. 2] for a detailed discussion of $\mathcal{D}$} and [Abu] for a non-equivariant construction over $\mathbb{Z}$}.
2.4. Local symplectic homology
Throughout this section all closed Reeb orbits, either prime and iterated, are required to be isolated even when this is not explicitly stated. The construction of symplectic homology, equivariant or non-equivariant, has a local counterpart. In this section we review this construction and the properties of local symplectic homology and then discuss its relation with filtered symplectic homology. While some of the definitions and results are standard and well-known to the experts, there are also several new observations.
2.4.1. Basic definitions and facts
Let $x$ be an isolated closed Reeb orbit, not necessarily prime. Fix an isolating neighborhood $U$ of $x({\mathbb{R}})$ in $M$. (By definition, $U$ is a neighborhood of $x$ containing no other closed orbit of period close to ${\mathcal{A}}(x)$.) Consider the product $\widehat{U}=U\times(1-\epsilon,1+\epsilon)\subset\widehat{W}$ and let $H=h(r)$ be a Hamiltonian such that $h^{\prime}(1)={\mathcal{A}}(x)$ and $h^{\prime\prime}(1)>0$. Then $x$ gives rise to an isolated set $S$ of 1-periodic orbits of $H$ with initial conditions on $x({\mathbb{R}})$. By definition, the symplectic homology $\operatorname{SH}(x;{\mathbb{F}})$ is the local Floer homology $\operatorname{HF}(H;{\mathbb{F}})$ of $H$ at $S$, and $\operatorname{CH}(x;{\mathbb{F}})$ is the $S^{1}$-equivariant local Floer homology.
Specifically, let $F$ be a non-degenerate 1-periodic in time perturbation of $H$ and $J$ a 1-periodic in time almost complex structure. When $F$ is sufficiently $C^{2}$-close to $H$, all 1-periodic orbits of the flow of $F$ that $x$ splits into and Floer cylinders between them are contained in $\widehat{U}$. (This is a variant of Gromov’s compactness, which follows, for instance, from the estimates in [Sa, Sec. 1.5].) For a generic choice of $J$ the regularity conditions are satisfied. Hence we obtain a complex and $\operatorname{SH}(x;{\mathbb{F}})$ is the homology of this complex. The equivariant version is defined in a similar fashion with appropriate modifications; see [GG20, Sec. 2.3].
This type of homology is the Floer-theoretic counterpart of critical modules from [GM69a, GM69b] and its definition goes back to [Fl89a, Fl89b]. The (non-equivariant) variant for isolated periodic orbits of Hamiltonian diffeomorphisms is studied in [Gi, GG10]. We refer the reader to [Fe20, Fe21, GG20, McL] for details in the Reeb case.
Similarly to the filtered homology, $\operatorname{SH}(x;{\mathbb{F}})$ and $\operatorname{CH}(x;{\mathbb{Q}})$ can be more explicitly described as follows. Consider a small non-degenerate perturbation of $\alpha$. Then $x$ splits into a finite collection of non-degenerate orbits $x_{i}$. Essentially by definition, $\operatorname{SH}(x;{\mathbb{F}})$ is the homology of a certain complex with generators $\check{x}_{i}$ and $\hat{x}_{i}$ of degrees $\mu(x_{i})$ and, respectively, $\mu(x_{i})+1$ and the differential coming from the Floer–Morse–Bott construction as in [Bo, BO09a] and also, e.g., [ÇG2M23]. Likewise, $\operatorname{CH}(x;{\mathbb{Q}})$ is the homology of a complex over $\mathbb{Q}$} generated by the orbits $x_{i}$ and graded by the Conley–Zehnder index; see [GG20, Sec. 3]. This is no longer literally true when $\mathbb{Q}$} is replaced by a field of positive characteristic; see Remark 2.6.
The absolute grading of the local symplectic homology is determined by fixing a symplectic trivialization of the contact structure along $x$. (Note that such a trivialization gives rise to trivializations along all closed orbits $x_{i}$ that $x$ splits into.) Below we will always assume that this trivialization agrees (up to homotopy) with the trivialization used in the construction of the filtered symplectic homology; see Section 3.1.3.
Remark 2.7.
As is clear from these definitions, $\operatorname{SH}(x;{\mathbb{F}})$ and $\operatorname{CH}(x;{\mathbb{F}})$ are defined even for germs of contact structures near $x({\mathbb{R}})$ as long as $x$ is an isolated closed Reeb orbit for the germ. The absolute grading is again determined by the choice of a trivialization of the contact structure along $x$.
The support of $\operatorname{SH}(x;{\mathbb{F}})$, denoted by $\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})$, is the collection of degrees $m$ such that $\operatorname{SH}_{m}(x;{\mathbb{F}})\neq 0$. The support $\operatorname{supp}\operatorname{CH}(x;{\mathbb{F}})$ is defined in a similar fashion. It readily follows from the definitions that
$\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})\subseteq[\mu_{-}(x),\mu_{+}(x)+1]\subseteq[\operatorname{\hat{\mu}}(x)-n+1,\operatorname{\hat{\mu}}(x)+n]$ (2.8)
and
$\operatorname{supp}\operatorname{CH}(x;{\mathbb{Q}})\subseteq[\mu_{-}(x),\mu_{+}(x)]\subseteq[\operatorname{\hat{\mu}}(x)-n+1,\operatorname{\hat{\mu}}(x)+n-1];$ (2.9)
see [GG20, Prop. 2.20]. Here $\operatorname{\hat{\mu}}$ is the mean index and $\mu_{\pm}$ are the lower and upper semi-continuous extensions of the Conley–Zehnder index; see Section 3.1.1.
For instance, when $x$ is non-degenerate, $\operatorname{SH}_{m}(x;{\mathbb{F}})$ is $\mathbb{F}$} in degrees $\mu(x)$ and $\mu(x)+1$ and zero in all other degrees. When in addition $x$ is good, $\operatorname{CH}_{m}(x;{\mathbb{Q}})={\mathbb{Q}}$ for $m=\mu(x)$ and zero otherwise. If $x$ is bad, $\operatorname{CH}(x;{\mathbb{Q}})=0$.
Furthermore, recall from [GG20] that equivariant and non-equivariant local symplectic homology groups fit into the Gysin exact sequence
$\cdots\to\operatorname{SH}_{m}(x;{\mathbb{F}})\to\operatorname{CH}_{m}(x;{\mathbb{F}})\stackrel{{\scriptstyle{\mathcal{D}}}}{{\longrightarrow}}\operatorname{CH}_{m-2}(x;{\mathbb{F}})\to\operatorname{SH}_{m-1}(x;{\mathbb{F}})\to\cdots.$
This sequence is just a local variant of the Gysin sequence from [BO13]. Moreover, by [GG20, Cor. 3.7], the shift operator $\mathcal{D}$} vanishes in local symplectic homology when $\mathbb{F}$$\mathbb{Q}$}, and the Gysin sequence splits into short exact sequences. (This is no longer true when ${\mathit{c}har}\,{\mathbb{F}}>0$ unless ${\mathit{c}har}\,{\mathbb{F}}$ and the order of iteration of $x$ are relatively prime.) As a consequence,
$\operatorname{SH}(x;{\mathbb{Q}})=\operatorname{CH}(x;{\mathbb{Q}})\oplus\operatorname{CH}(x;{\mathbb{Q}})[-1].$ (2.10)
As was mentioned above, a closed orbit $x$ is said to be $\mathbb{F}$}-visible if $\operatorname{SH}(x;{\mathbb{F}})\neq 0$. Likewise, $x$ is equivariantly $\mathbb{F}$}-visible if $\operatorname{CH}(x;{\mathbb{F}})\neq 0$. An $\mathbb{F}$}-visible closed orbit is automatically equivariantly $\mathbb{F}$}-visible by the Gysin sequence. The converse is also true when $\mathbb{F}$$\mathbb{Q}$} due to (2.10).
It is clear from this construction that the Euler characteristic of $\operatorname{SH}(x;{\mathbb{F}})$ is always zero for any field $\mathbb{F}$} and any closed Reeb orbit $x$, i.e.,
$\chi(x;{\mathbb{F}}):=\sum_{m}(-1)^{m}\dim\operatorname{SH}_{m}(x;{\mathbb{F}})=0.$ (2.11)
Indeed, as stated in the beginning of this section $\operatorname{SH}(x;{\mathbb{F}})$ is the homology of a complex with generators coming in pairs $\check{x}_{i}$ and $\hat{x}_{i}$ where the degree of $\check{x}_{i}$ is $\mu(x_{i})$ and the degree of $\hat{x}_{i}$ is $\mu(x_{i})+1$. Therefore,
$\dim\operatorname{SH}(x;{\mathbb{F}})\textrm{ is always even,}$
and so
$\dim\operatorname{SH}(x;{\mathbb{F}})\geq 2\textrm{ when $x$ is ${\mathbb{F}}$-visible.}$ (2.12)
The equivariant Euler characteristic of an orbit $x$ is defined as
$\chi^{\scriptscriptstyle{eq}}(x)=\sum_{m}(-1)^{m}\dim\operatorname{CH}_{m}(x;{\mathbb{Q}}).$ (2.13)
For instance, when $x$ is non-degenerate, $\chi^{\scriptscriptstyle{eq}}(x)$ is either $(-1)^{\mu(x)}$ or 0 depending on whether $x$ is good or bad. We note that here we restrict our background field $\mathbb{F}$} to $\mathbb{Q}$} because otherwise $\operatorname{supp}\operatorname{CH}(x;{\mathbb{F}})$ can be infinite and then the right-hand side of (2.13) is not defined.
It readily follows from the description of $\operatorname{CH}(x;{\mathbb{Q}})$ in Section 2.4.1 that, whenever $x$ is prime, $\chi^{\scriptscriptstyle{eq}}(x)$ is equal to the Hopf index of the Poincaré return map of $x$ up to the sign $(-1)^{n}$. This is no longer true when $x$ is iterated. For instance, $\chi^{\scriptscriptstyle{eq}}(x)=0$ when $x$ is non-degenerate and bad, while the Hopf index is $\pm 1$. Moreover, let $x=z^{k}$ where $z$ is prime and, for the sake of simplicity, $k$ is odd. Then $\chi^{\scriptscriptstyle{eq}}(x)$ counts the number (with signs) of $k$-periodic orbits of a small non-degenerate perturbation of the return map $\varphi$ of $z$, while the Hopf index of $\varphi^{k}$ counts (with the same signs) the number of $k$-periodic points.
2.4.2. Behavior under iterations
Recall that a positive integer $k$ and the iterate $x^{k}$ are said to be admissible for $x$ if $k$ is not divisible by the degree of any root of unity other than $1$ among the eigenvalues of the linearized Poincaré return map of $x$. For instance, all $k\in{\mathbb{N}}$ are admissible when $x$ is totally degenerate, i.e., all eigenvalues are equal to 1.
We will need the following simple fact which is conceptually closely related to the Shub–Sullivan Theorem and its proof; see [SS].
Lemma 2.8.
Let $x$ be an isolated (not necessarily prime) closed Reeb orbit. Assume that $k$ is admissible for $x$ and odd if the number of eigenvalues of the return map of $x$ in (1,0)(-1,0) is odd. Then
$\chi^{\scriptscriptstyle{eq}}\big(x^{k}\big)=\chi^{\scriptscriptstyle{eq}}(x).$ (2.14)
In particular, $x^{k}$ is $\mathbb{Q}$}-visible when $\chi^{\scriptscriptstyle{eq}}(x)\neq 0$.
Proof.
Fix $k$ and consider a $C^{2}$-small non-degenerate perturbation of the background contact form. Under this perturbation $x$ splits into strongly non-degenerate closed orbits $x_{i}$. The key fact is that $x_{i}^{k}$ are the only closed orbits of the perturbed flow in a neighborhood of $x$ with period close to $k{\mathcal{A}}(x)$, provided that $k$ is admissible and the perturbation is sufficiently small (depending on $k$). This follows for instance from [GG10, Prop. 1.1].
We claim that $x_{i}^{k}$ is good (bad) if and only if so is $x_{i}$, and that $\mu(x_{i})$ and $\mu\big(x^{k}_{i}\big)$ have the same parity. When $k$ is odd this is obvious. Indeed, the eigenvalues in (1,0)(-1,0) remain in that range for all iterates and additional eigenvalues come in pairs, which implies the parity statement. Likewise, an odd iterate of a good (bad) closed orbit is automatically good (bad). If 1-1 is an eigenvalue of the return map of $x$, the iteration $k$ is necessarily odd, since $k$ is admissible, and the claim follows. Assume that 1-1 is not an eigenvalue and $k$ is even. Then for each $x_{i}$ the parity of the number of real eigenvalues in (1,0)(-1,0) is the same as for $x$ and, by the conditions of the lemma, even. Hence, in this case, $\mu(x_{i})$ and $\mu\big(x^{k}_{i}\big)$ also have the same parity. Furthermore, if $x_{i}=z^{r}$ where $z$ is prime, $x_{i}$ is good (bad) if and only $\mu(x_{i})$ and $\mu(z)$ have the same (opposite) parity. This is also true for $x_{i}^{k}=z^{kr}$. As a consequence $x_{i}$ and $x_{i}^{k}$ are good (bad) simultaneously.
As described in Section 2.4.1, $\operatorname{CH}(x;{\mathbb{Q}})$ is the homology of a certain complex generated by good closed orbits $x_{i}$, and $\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)$ is the homology of a complex generated by good $x_{i}^{k}$. In both cases the complex is graded by the Conley–Zehnder index. Therefore,
$\chi^{\scriptscriptstyle{eq}}\big(x^{k}\big)=\sum_{\text{good }x_{i}^{k}}(-1)^{\mu(x^{k}_{i})}=\sum_{\text{good }x_{i}}(-1)^{\mu(x_{i})}=\chi^{\scriptscriptstyle{eq}}(x).$
Next, for an isolating neighborhood $U$ of a closed Reeb orbit $x$, consider the $k$-fold covering $\pi\colon U_{k}\to U$ and the pull-back of the contact form from $U$ to $U_{k}$. This pull-back and hence the Reeb vector field are invariant under the action of ${\mathbb{Z}}_{k}$ on $U_{k}$ by deck transformations. Then $\pi^{-1}(x({\mathbb{R}}))$ is an isolated closed Reeb orbit of the lift, which we denote by $kx$, provided that $x^{k}$ is isolated. The period (action) of $kx$ is $k{\mathcal{A}}(x)$. By Remark 2.7, we have the local symplectic homology $\operatorname{SH}(kx;{\mathbb{F}})$ and $\operatorname{CH}(kx;{\mathbb{F}})$ of $kx$ defined.
Lemma 2.9.
There is a natural ${\mathbb{Z}}_{k}$-action on $\operatorname{SH}(kx;{\mathbb{F}})$ and $\operatorname{CH}(kx;{\mathbb{F}})$, and
$\operatorname{SH}\big(x^{k};{\mathbb{Q}}\big)=\operatorname{SH}(kx;{\mathbb{Q}})^{{\mathbb{Z}}_{k}}\textrm{ and }\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)=\operatorname{CH}(kx;{\mathbb{Q}})^{{\mathbb{Z}}_{k}},$ (2.15)
where the superscript refers to the ${\mathbb{Z}}_{k}$-invariant part of the homology.
Note that here the orbit $x$ need not be prime and $k$ is not required to be admissible. However, $x^{k}$ must be isolated. Furthermore, the condition that $\mathbb{F}$$\mathbb{Q}$} in (2.15) can be relaxed. It would be sufficient to assume that ${\mathit{c}har}\,{\mathbb{F}}$ and $k$ are relatively prime.
Proof.
The argument is quite standard and we only briefly outline it. For the sake of brevity, we focus on the non-equivariant case; for the equivariant version is handled in a similar fashion.
In the notation from Section 2.4.1, let $H=h(r)\colon\widehat{U}\to{\mathbb{R}}$ be a Hamiltonian from the construction of $\operatorname{SH}\big(x^{k};{\mathbb{F}}\big):=\operatorname{HF}(H;{\mathbb{F}})$. Fix a non-degenerate 1-periodic in time perturbation $F$ of $H$ and a regular 1-periodic almost complex structure $J$ on $\widehat{U}$. Thus, in particular, $\operatorname{HF}(H;{\mathbb{F}})=\operatorname{HF}(F;{\mathbb{F}})$. Denote by $H_{k}$, $F_{k}$ and $J_{k}$ the pull-backs of the Hamiltonians and $J$ to the $k$-fold covering $\pi\colon\widehat{U}_{k}\to\widehat{U}$. Then $\operatorname{SH}(kx;{\mathbb{F}}):=\operatorname{HF}(H_{k};{\mathbb{F}})=\operatorname{HF}(F_{k};{\mathbb{F}})$. These pull-backs are invariant under the ${\mathbb{Z}}_{k}$-action on $\widehat{U}_{k}$ by deck transformations. The 1-periodic orbits of $F_{k}$ project to the periodic orbits of $F$, and so do Floer cylinders. Moreover, on the sets of 1-periodic orbits and Floer cylinders these projections are onto and $k$-to-1. To be more precise, since all 1-periodic orbits of $F$ are in the free homotopy class $k[S^{1}]$, the Hamiltonian $F_{k}$ has exactly $k$ distinct 1-periodic orbits lying above every 1-periodic orbit of $F$, and deck transformations act freely on these orbits. It follows that the same is true for Floer cylinders; cf. [Sa, Prop. 1.21].
We claim that the pair $(F_{k},J_{k})$ is regular if and only if the pair $(F,J)$ is regular; cf. [KS, Prop. 5.13]. Indeed, let $u$ be a Floer cylinder in $\widehat{U}_{k}$. Then the pull-backs $u^{*}T\widehat{U}_{k}$ and $(\pi u)^{*}T\widehat{U}$ are naturally isomorphic. (This is true whenever $\pi$ is a local diffeomorphism.) Hence, we obtain an isomorphism between the Sobolev spaces of sections of these bundles; see [Sa, Sec. 2] for the definitions. By construction, this isomorphism intertwines the linearized perturbed Cauchy–Riemann operators $D_{u}$ along $u$ and $D_{\pi u}$ along $\pi u$. Therefore, $D_{u}$ is onto if and only if $D_{\pi u}$ is onto. Finally, since every Floer cylinder in $\widehat{U}$ has a lift to $\widehat{U}_{k}$, we conclude that the pairs $(F,J)$ and $(F_{k},J_{k})$ are regular simultaneously.
As a consequence, ${\mathbb{Z}}_{k}$ acts on the Floer complex $\operatorname{CF}(H_{k},J_{k})$ of $(H_{k},J_{k})$ and hence on $\operatorname{SH}(kx;{\mathbb{F}})=\operatorname{HF}(F_{k};{\mathbb{F}})$. The standard continuation argument shows that this action is independent of the auxiliary choices made in the construction.
When $\mathbb{F}$$\mathbb{Q}$}, the Floer complex $\operatorname{CF}(F,J)$ is naturally isomorphic to the invariant part $\operatorname{CF}(H_{k},J_{k})^{{\mathbb{Z}}_{k}}$ of $\operatorname{CF}(H_{k},J_{k})$. Indeed, in one direction we have (for any ${\mathbb{F}})$ the homomorphism of complexes $\operatorname{CF}(H_{k},J_{k})\to\operatorname{CF}(F,J)$ induced by the projection $\pi\colon\widehat{U}_{k}\to\widehat{U}$. In the opposite direction, we have the averaging map. Namely, for $z\in\operatorname{CF}(F,J)$, pick an arbitrary lift $z^{\prime}\in\operatorname{CF}(H_{k},J_{k})$ and set, in the obvious notation,
$z\mapsto\frac{1}{k}\sum_{g\in{\mathbb{Z}}_{k}}gz^{\prime}\in\operatorname{CF}(H_{k},J_{k})^{{\mathbb{Z}}_{k}}.$
This map gives rise to an isomorphism $\operatorname{CF}(F,J)\cong\operatorname{CF}(H_{k},J_{k})^{{\mathbb{Z}}_{k}}$, which concludes the proof of (2.15). ∎
Local symplectic homology of $x$ over $\mathbb{Q}$} is intimately connected to the local Floer homology $\operatorname{HF}(\varphi;{\mathbb{Q}})$ of the return map $\varphi$, although this connection is not straightforward. When $x$ is prime,
$\operatorname{CH}(x;{\mathbb{Q}})=\operatorname{HF}(\varphi;{\mathbb{Q}})\textrm{ and }\operatorname{SH}(x;{\mathbb{Q}})=\operatorname{HF}(\varphi;{\mathbb{Q}})\oplus\operatorname{HF}(\varphi;{\mathbb{Q}})[-1]$ (2.16)
as is shown in [Fe20, Fe21]. (These two identifications are equivalent due to (2.10).) Applying this to $kx$, which is prime whenever $x$ is prime and using (2.15), we conclude that
$\dim\operatorname{SH}_{m}\big(x^{k};{\mathbb{Q}}\big)\leq\dim\operatorname{HF}_{m}\big(\varphi^{k};{\mathbb{Q}}\big)+\dim\operatorname{HF}_{m-1}\big(\varphi^{k};{\mathbb{Q}}\big)$ (2.17)
for all $k\in{\mathbb{N}}$, provided that $x$ is prime and $x^{k}$ is isolated. A closely related inequality is established in [McL, Lemma 3.4]. The inequalities in (2.17) can be strict. This is the case, for instance, when $x$ is non-degenerate and alternating and $k$ is even; see also Examples 2.12 and 2.13 below.
Conjecturally,
$\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)=\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)^{{\mathbb{Z}}_{k}}$
and
$\operatorname{SH}\big(x^{k};{\mathbb{Q}}\big)=\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)^{{\mathbb{Z}}_{k}}\oplus\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)^{{\mathbb{Z}}_{k}}[-1]$
for all $k\in{\mathbb{N}}$; see [GH2M]. Lemma 2.9 together with (2.16) almost proves these identifications. The missing part is that the isomorphisms in (2.16) commute with the ${\mathbb{Z}}_{k}$-action when applied to $kx$.
Remark 2.10.
When the orbit $x$ is prime, the local cylindrical contact homology of $x$ is defined and isomorphic to $\operatorname{HF}(\varphi;{\mathbb{Q}})$; see [HM]. (Alternatively, one can use here a local version of [EKP, Prop. 4.30].) Furthermore, localizing the isomorphism between the equivariant symplectic and cylindrical contact homology from [BO09b] we see that the local cylindrical contact homology of $x$ is isomorphic to $\operatorname{CH}(x;{\mathbb{Q}})$, assuming again that $x$ is prime. (Hence the notation.) This gives another proof of the isomorphisms in (2.16), albeit quite indirect.
Remark 2.11.
The behavior under iterations of local symplectic or Floer homology (equivariant or not) is complicated and overall poorly understood. For instance, it is not known if a closed orbit can go from invisible to visible under iterations, i.e., if a closed orbit $x^{k}$ can be $\mathbb{F}$}-visible when $x$ is prime and $\mathbb{F}$}-invisible, without additional assumptions on $x$, the ground field $\mathbb{F}$} or $k$. At the same time, $x^{k}$ can in general become $\mathbb{F}$}-invisible when $x$ is prime and $\mathbb{F}$}-visible. A simple situation when this happens is when $x$ is non-degenerate and alternating, $k$ is even and $\mathbb{F}$$\mathbb{Q}$}. In Examples 2.12 and 2.13 we describe more interesting cases of this phenomenon. Lemma 2.8 provides a simple non-vanishing criterion, which is sufficient for our purposes. (The question for a variant of local symplectic homology and admissible iterations has been investigated in [HHM]. The homology there is defined via generating functions and is hypothetically isomorphic to the local homology used here; cf. Remark 4.5.)
Example 2.12 (Vanishing of iterated symplectic homology).
Let $\psi\colon{\mathbb{R}}^{2}={\mathbb{C}}\to{\mathbb{C}}={\mathbb{R}}^{2}$ be the time-one flow of the degree-three monkey saddle Hamiltonian $H=\Re(z^{3})$ and let $R_{2\pi/3}$ be the rotation of ${\mathbb{R}}^{2}$ counterclockwise by $2\pi/3$. (We view $R_{2\pi/3}$ as an element of the universal covering $\operatorname{\widetilde{Sp}}(2)$.) Note that $R_{2\pi/3}$ and the flow of $H$ commute. Set $\varphi=R_{2\pi/3}\psi$ and let $x$ be a closed Reeb orbit with the return map $\varphi$. The orbit $x$ is non-degenerate, $\operatorname{CH}(x;{\mathbb{Q}})={\mathbb{Q}}$ and for a suitable trivialization $\operatorname{CH}(x;{\mathbb{Q}})$ is supported in degree one. The return map of $x^{3}$ is $\varphi^{3}$, which is the time-three flow of $H$. One can show that $\operatorname{CH}(x^{3};{\mathbb{Q}})=0$. (This is not obvious.) However, $\operatorname{HF}(\varphi^{3};{\mathbb{Q}})$ is ${\mathbb{Q}}^{2}$ and supported in degree zero. This gives another example where the inequality in (2.17) is strict; see also Example 2.13 below.
Example 2.13.
The notion of a (non-)alternating closed orbit is not particularly useful in the setting where $x$ is not strongly non-degenerate and the return map $\varphi$ has roots of unity of even degree among its eigenvalues. The reason is that in this case the behavior of the local symplectic homology is not determined by the eigenvalues and depends on higher order terms. For instance, in the spirit of Example 2.12, let $R_{\pi}\colon{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ be the counterclockwise rotation by $\pi$, i.e., central symmetry, viewed as an element of $\operatorname{\widetilde{Sp}}(2)$. Consider the Hamiltonians $H_{0}=p^{4}-q^{4}$ and $H_{1}=(p^{2}+q^{2})^{2}$ on ${\mathbb{R}}^{2}$, where $(p,q)$ are the Darboux coordinates. The flows of $H_{0}$ and $H_{1}$ commute with $R_{\pi}$. Let $\varphi_{i}$, for $i=0,1$, be obtained by composing $R_{\pi}$ with the time-one flow of $H_{i}$ and $x_{i}$ be a closed Reeb orbit with the return map $\varphi_{i}$. Clearly, $\operatorname{CH}(x_{i};{\mathbb{Q}})=\operatorname{HF}(\varphi_{i})={\mathbb{Q}}$ is supported in degree one. However, one can show that $\operatorname{CH}\big(x_{0}^{2};{\mathbb{Q}}\big)=0$ and $\operatorname{HF}\big(\varphi_{0}^{2};{\mathbb{Q}}\big)={\mathbb{Q}}$ (supported in degree two). At the same time, $\operatorname{CH}\big(x_{1}^{2};{\mathbb{Q}}\big)=\operatorname{HF}\big(\varphi_{1}^{2};{\mathbb{Q}}\big)={\mathbb{Q}}$ (supported in degree three).
2.4.3. From local to filtered homology
In what follows, it is convenient to adopt the following convention; cf. Section 2.3. Namely, we will purely formally treat the domain $W$ as a collection of “closed orbits” with action zero. By definition, these orbits are in one-to-one correspondence with the bars in $\operatorname{SH}(W;{\mathbb{F}})$ beginning at zero; each of these orbits has one-dimensional symplectic homology concentrated in the degree of that bar and there are $\dim\operatorname{H}(W;{\mathbb{F}})$ such bars. For instance, when $W$ is star-shaped it contributes one orbit $[W]$ that has action zero and degree $n$.
Local symplectic homology spaces are building blocks for the filtered symplectic homology. This principle can be formalized in several ways. For instance, assume that $a$ is the only point of ${\mathcal{S}}(\alpha)$ in an interval $I=(t,s)$. Then
$\sum_{{\mathcal{A}}(x)=a}\dim\operatorname{SH}_{m}(x;{\mathbb{F}})=\dim\operatorname{SH}^{I}_{m}(W;{\mathbb{F}})$ (2.18)
In fact,
$\bigoplus_{{\mathcal{A}}(x)=a}\operatorname{SH}(x;{\mathbb{F}})=\operatorname{SH}^{I}(W;{\mathbb{F}}).$ (2.19)
by the definition of local symplectic homology. Therefore, we have the long exact sequence
$\ldots\to\operatorname{SH}^{t}_{m}(W;{\mathbb{F}})\to\operatorname{SH}^{s}_{m}(W;{\mathbb{F}})\to\bigoplus_{{\mathcal{A}}(x)=a}\operatorname{SH}_{m}(x;{\mathbb{F}})\to\operatorname{SH}^{t}_{m-1}(W;{\mathbb{F}})\to\ldots.$ (2.20)
Moreover, for any field $\mathbb{F}$} and any interval $I$, there exists a spectral sequence
$E^{r}\Rightarrow\operatorname{SH}^{I}(W;{\mathbb{F}})\textrm{ with }E^{1}=\bigoplus_{{\mathcal{A}}(x)\in I}\operatorname{SH}(x;{\mathbb{F}}),$ (2.21)
where, when $0\in I$, we treat $W$ as a collection of “orbits” with action 0 as described above. This spectral sequence is simply associated with the action filtration; see, e.g., [GG20, Sec. 2].
In a similar vein, we have analogues of (2.19) and (2.20) for equivariant symplectic homology and a spectral sequence
$E^{r}\Rightarrow\operatorname{CH}^{I}(W;{\mathbb{F}})\textrm{ with }E^{1}=\bigoplus_{{\mathcal{A}}(x)\in I}\operatorname{CH}(x;{\mathbb{F}}),$ (2.22)
where for the sake of simplicity we have assumed that $0\not\in I$. As a consequence, when $W$ is a star-shaped domain, for every $m\in n-1+2{\mathbb{N}}$ there must be a closed orbit $x$ with $m\in\operatorname{supp}\operatorname{CH}(x;{\mathbb{F}})$ due to (2.7). Furthermore, if $q\in n+2{\mathbb{N}}$ and $q\in\operatorname{supp}\operatorname{CH}(y;{\mathbb{F}})$ we must have
$\sum_{x\neq y}\dim\operatorname{CH}_{q-1}(x;{\mathbb{F}})\geq 2\quad\textrm{or}\quad\sum_{x\neq y}\dim\operatorname{CH}_{q+1}(x;{\mathbb{F}})\geq 2.$ (2.23)
This can also be easily seen from (2.7) and the equivariant version of (2.20).
Remark 2.14.
When $\mathbb{F}$$\mathbb{Q}$}, the flow is non-degenerate and again $I\subset(0,\infty)$, the spectral sequence (2.22), can be reassembled into a complex. Namely, we have a complex generated over $\mathbb{Q}$} by good closed orbits with action in $I$, graded by the Conley–Zehnder index and filtered by the action, and such that its homology is $\operatorname{CH}^{I}(W;{\mathbb{Q}})$; see [GG20, Sec. 2.5]. The differential in the complex is completely determined by the same auxiliary data as the one from the construction of the equivariant symplectic or Floer homology; it is functorial in a suitable sense, etc. (Without non-degeneracy condition, the complex is the $E^{1}$-term in (2.22) but then we also need to fix more additional data to define the differential.) We omit the details of this construction because it is not explicitly used in the paper. Here we only note that this is no longer literally true when ${\mathit{c}har}\,{\mathbb{F}}\neq 0$; for then $\operatorname{supp}\operatorname{CH}(x;{\mathbb{F}})$ can, for instance, be infinite.
Finally, let $\zeta_{m}^{-}(a)$ be the number of bars in the barcode of $\operatorname{SH}(W;{\mathbb{F}})$ of degree m1m-1 ending at $a$ and $\zeta^{+}_{m}(a)$ be the number of bars of degree $m$ beginning at $a$. Set $\zeta_{m}(a)=\zeta_{m}^{-}(a)+\zeta_{m}^{+}(a)$. The following observation is nearly obvious.
Lemma 2.15 (“Morse Theory”).
For all $a\in{\mathbb{R}}$, we have
$\zeta_{m}(a)=\sum_{{\mathcal{A}}(x)=a}\dim\operatorname{SH}_{m}(x;{\mathbb{F}}).$ (2.24)
Proof.
Throughout the proof we suppress the ground field $\mathbb{F}$} in the notation. It is sufficient to prove the lemma when $a\in{\mathcal{S}}(\alpha)$ or $a=0$; for otherwise both sides of (2.24) are zero. Using the background assumption (BA), fix $\epsilon>0$ so small that $a$ is the only point of the intersection $I\cap({\mathcal{S}}(\alpha)\cup\{0\})$, where $I=(a-\epsilon,a+\epsilon)$. Clearly,
$\zeta_{m}^{-}(a)=\dim\ker\left(\operatorname{SH}^{a-\epsilon}_{m-1}(W)\to\operatorname{SH}^{a+\epsilon}_{m-1}(W)\right)$
and
$\zeta_{m}^{+}(a)=\dim\operatorname{coker}\left(\operatorname{SH}^{a-\epsilon}_{m}(W)\to\operatorname{SH}^{a+\epsilon}_{m}(W)\right).$
From the long exact sequence
$\cdots\to\operatorname{SH}^{a-\epsilon}_{m}(W)\to\operatorname{SH}^{a+\epsilon}_{m}(W)\to\operatorname{SH}^{I}_{m}(W)\to\operatorname{SH}^{a-\epsilon}_{m-1}(W)\to\operatorname{SH}^{a+\epsilon}_{m-1}(W)\to\cdots$
we see that also
$\zeta_{m}^{-}(a)=\dim\operatorname{im}\left(\operatorname{SH}^{I}_{m}(W)\to\operatorname{SH}^{a-\epsilon}_{m-1}(W)\right)$
and
$\zeta_{m}^{+}(a)=\dim\operatorname{im}\left(\operatorname{SH}^{a+\epsilon}_{m}(W)\to\operatorname{SH}^{I}_{m}(W)\right)=\dim\ker\left(\operatorname{SH}^{I}_{m}(W)\to\operatorname{SH}^{a-\epsilon}_{m-1}(W)\right).$
Hence, again since the sequence is exact,
$\zeta_{m}^{-}(a)+\zeta_{m}^{+}(a)=\dim\operatorname{SH}^{I}_{m}(W),$
and (2.24) follows from (2.18). ∎
2.5. The beginning and end maps
For a fixed field $\mathbb{F}$}, denote by $\mathcal{B}$} or ${\mathcal{B}}({\mathbb{F}})$ the collection of bars of the graded persistence module $\operatorname{SH}(W;{\mathbb{F}})$. Assume that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})=0$. Then every bar in $\mathcal{B}$} is finite by Theorem 2.3. Furthermore, let ${\mathcal{P}}={\mathcal{P}}({\mathbb{F}})$ be the set of $\mathbb{F}$}-visible closed Reeb orbits. As in the previous section, let us also formally add to $\mathcal{P}$} the set of the “beginning” of the bars of the form $I=(0,a]$ for some $a>0$, denoting the resulting set by $\mathcal{P}$$^{\prime}$. These additional elements $x$ of $\mathcal{P}$$^{\prime}$ are simply in one-to-one correspondence with such bars. By definition, each such “orbit” $x$ has symplectic homology equal to $\mathbb{F}$} and concentrated in $\deg(I)$, and zero action. (We also set $\mu_{-}(x)=\mu_{+}(x):=\deg(x)$ for such $x$.) When $W$ is star-shaped, we are adding just one extra element corresponding to the domain $W$, which we denote by $[W]$. This element has action zero and is supported in degree $n$.
The sets $\mathcal{P}$} and $\mathcal{P}$$^{\prime}$ depend on $\mathbb{F}$} or, to be more precise, on ${\mathit{c}har}\,{\mathbb{F}}$ even when the Reeb flow is non-degenerate and $W$ is star-shaped. For instance, when ${\mathit{c}har}\,{\mathbb{F}}=2$ we have two generators for each closed orbit of the flow regardless of whether the orbit is good or bad. However, when ${\mathit{c}har}\,{\mathbb{F}}\neq 2$, only good orbits contribute.
Theorem 2.16.
There exist maps $\operatorname{Beg}\colon{\mathcal{B}}\to{\mathcal{P}}^{\prime}$ (the beginning of $I$) and $\operatorname{End}\colon{\mathcal{B}}\to{\mathcal{P}}$ (the end of $I$) such that for every bar $I=(a,b]$ and $x=\operatorname{Beg}(I)$ and $y=\operatorname{End}(I)$ we have
  • ${\mathcal{A}}(x)=a$ and ${\mathcal{A}}(y)=b$,
  • $\deg(I)\in\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})$ and $\deg(I)+1\in\operatorname{supp}\operatorname{SH}(y;{\mathbb{F}})$.
As a consequence of the second of these requirements, we see that
$\mu_{-}(x)\leq\deg(I)\leq\mu_{+}(x)+1$
and
$\mu_{-}(y)-1\leq\deg(I)\leq\mu_{+}(y),$
and therefore
$\mu_{-}(y)-\mu_{+}(x)\leq 2.$ (2.25)
Moreover,
$\deg(I)=\mu_{+}(x)+1=\mu_{-}(y)-1,$ (2.26)
whenever the equality holds in (2.25). Note also that the second condition of the theorem necessitates that the orbits in $\mathcal{P}$} are $\mathbb{F}$}-visible. We will see later that in the setting of Theorem C and as a consequence of that theorem, the maps $\operatorname{Beg}$ and $\operatorname{End}$ are onto. We do not know if this is true in general. Furthermore, in the setting of Theorem C and again as a consequence of that theorem, both maps are one-to-one. This is certainly not true in general. For instance, when $\dim\operatorname{SH}(x;{\mathbb{F}})=3$ and $x$ is the only closed orbit with action ${\mathcal{A}}(x)$, at least one of the pre-images $\operatorname{Beg}^{-1}(x)$ or $\operatorname{End}^{-1}(x)$ must contain more than one point.
In general, even when the Reeb flow is non-degenerate, the maps $\operatorname{Beg}$ and $\operatorname{End}$ are not canonically defined and furthermore, just as the set $\mathcal{P}$}, depend on ${\mathit{c}har}\,{\mathbb{F}}$. However, the construction can be made canonical if all closed orbits have distinct actions.
Proof of Theorem 2.16.
Let $\mathbb{F}$} be the underlying coefficient field and let
$e(m,a)_{1},e(m,a)_{2},\dots,e(m,a)_{\zeta_{m}(a)}$
be an $\mathbb{F}$}-basis for
$E:=\bigoplus_{{\mathcal{A}}(x)=a}\operatorname{SH}_{m}(x;{\mathbb{F}}),$
where $\zeta_{m}(a)$ is the bar counting function from Lemma 2.15. To keep track of the orbits, we require in addition the basis elements $e(m,a)_{i}$ to be adapted to the above direct sum decomposition of $E$. Next, let ${\mathcal{B}}=\{I_{1},I_{2},\ldots\}$ be an arbitrary ordering by $\mathbb{N}$} of the bars in the symplectic homology barcode $\mathcal{B}$}. (Note that, by the background assumption (BA), there are countably many bars in $\mathcal{B}$}.) Let us denote by $\zeta^{i}_{m}(a)$ the restriction of $\zeta_{m}(a)$ to the “truncated” barcodes ${\mathcal{B}}^{i}=\{I_{1},I_{2},\dots,I_{i}\}$. In other words, $\zeta^{i}_{m}(a)$ is the total number of bars of degree m1m-1 ending at $a$ and of degree $m$ beginning at $a$ among the first $i$ bars.
Now we are in a position to define the maps $\operatorname{Beg}$ and $\operatorname{End}$. To simplify the notation, set $(c_{i},b_{i}]:=I_{i}$ and $k_{i}:=\deg(I_{i})$. Let us denote by $\pi$ the natural projection map sending $e(m,a)_{i}$ to the corresponding element in $\mathcal{P}^{\prime}$. (Here we use the requirement that the basis respects the direct sum decomposition). We define $\operatorname{Beg}\colon{\mathcal{B}}\to\mathcal{P}^{\prime}$ by
$I_{i}\mapsto\pi\big(e(k_{i},c_{i})_{\zeta^{i}_{k_{i}}(c_{i})}\big)$
and, similarly, $\operatorname{End}\colon{\mathcal{B}}\to\mathcal{P}^{\prime}$ by
$I_{i}\mapsto\pi\big(e(k_{i}+1,b_{i})_{\zeta^{i}_{k_{i}+1}(b_{i})}\big).$
We leave it to the reader to check that the maps above satisfy the requirements. ∎
Remark 2.17.
Our requirements on the maps $\operatorname{Beg}$ and $\operatorname{End}$ are deliberately minimal. For instance, with more effort, we could construct these maps so that the sum of the number of bars of degree m1m-1 in $\operatorname{End}^{-1}(x)$ and the number of bars in $\operatorname{Beg}^{-1}(x)$ of degree $m$ is equal to $\dim\operatorname{SH}_{m}(x;{\mathbb{F}})$. This would be a refinement of (2.24) in Lemma 2.15. However, in the cases we are interested in, this condition is automatically satisfied as a consequence of (2.24), and we omit this construction.
3. Conley-Zehnder index and index recurrence
The goal of this section section is two-fold. In the first part we review several Conley–Zehnder index type invariants and their properties, and the notion of dynamical convexity, starting with the linear algebra setting and then turning to closed Reeb orbits. This material is quite standard. The second part is devoted to index recurrence and its illustrative applications, and this part is mainly new.
3.1. Background: the Conley–Zehnder index and dynamical convexity
3.1.1. Conley–Zehnder index: linear algebra
In this section for, the reader’s convenience we recall some basic properties of the mean and Conley–Zehnder indices. We refer the reader to, e.g., [Lo02] or [SZ, Sec. 3] or [GG20, Sec. 4], which we closely follow here, for a more thorough treatment.
Let $\operatorname{\widetilde{Sp}}(2m)$ be the universal covering of the group $\operatorname{Sp}(2m)$ of symplectic linear transformations of ${\mathbb{R}}^{2m}$. We view elements of $\operatorname{\widetilde{Sp}}(2m)$ as paths starting at the identity, taken up to homotopy with fixed end-points, and abusing terminology we will often refer to elements of $\operatorname{\widetilde{Sp}}(2m)$ as maps. Unless specifically stated otherwise, we assume that such paths are parametrized by $[0,1]$.
A map $\operatorname{\hat{\mu}}$ from a Lie group to $\mathbb{R}$} is said to be a quasimorphism if it fails to be a homomorphism only up to a constant, i.e.,
$\big|\operatorname{\hat{\mu}}(\Phi\Psi)-\operatorname{\hat{\mu}}(\Phi)-\operatorname{\hat{\mu}}(\Psi)\big|<{\mathit{c}onst},$
where the constant is independent of $\Phi$ and $\Psi$. One can prove that there is a unique quasimorphism $\operatorname{\hat{\mu}}\colon\widetilde{\operatorname{Sp}}(2m)\to{\mathbb{R}}$ which is continuous and homogeneous (i.e., $\operatorname{\hat{\mu}}(\Phi^{k})=k\operatorname{\hat{\mu}}(\Phi)$) and satisfies the normalization condition:
$\operatorname{\hat{\mu}}(\Phi_{0})=2\quad\textrm{for}\quad\Phi_{0}(t)=\exp\big(2\pi\sqrt{-1}t\big)\oplus I_{2m-2}$
with $t\in[0,\,1]$, in the self-explanatory notation; see [BG]. The quasimorphism $\operatorname{\hat{\mu}}$ is called the mean index. The continuity requirement holds automatically and is not necessary for the characterization of $\operatorname{\hat{\mu}}$. (This is not obvious). The mean index is also automatically conjugation invariant, as a consequence of the homogeneity, and $\operatorname{\hat{\mu}}$ restricts to an isomorphism $\pi_{1}(\operatorname{Sp}(2m))\to 2{\mathbb{Z}}$. Furthermore, $\operatorname{\hat{\mu}}(\Phi)\in{\mathbb{Z}}$ whenever the elliptic eigenvalues of $\Phi(1)$ are $\pm 1$ or when $\Phi(1)$ is hyperbolic. (Elliptic eigenvalues are by definition unit eigenvalues and the elliptic part of $A\in\operatorname{Sp}(2m)$ is the restriction of $A$ to the sum of the generalized eigenspaces for all elliptic eigenvalues.)
Recall that $A\in\operatorname{Sp}(2m)$ is said to be non-degenerate if all eigenvalues of $A$ are different from one, and strongly non-degenerate if all powers $A^{k}$, $k\in{\mathbb{N}}$, are non-degenerate. (Similarly, an element $\Phi\in\operatorname{\widetilde{Sp}}(2m)$ is non-degenerate when $\Phi(1)$ is.) We denote the set of non-degenerate symplectic linear maps by $\operatorname{Sp}^{*}(2m)$ and the part of $\operatorname{\widetilde{Sp}}(2m)$ lying over $\operatorname{Sp}^{*}(2m)$ by $\widetilde{\operatorname{Sp}}{}^{*}(2m)$. It is not hard to see that $A=\Phi(1)$ can be connected to a symplectic map with the elliptic part equal to I-I (if non-trivial) by a path $\Psi$ lying entirely in $\operatorname{Sp}^{*}(2m)$. Concatenating this path with $\Phi$, we obtain a new path $\Phi^{\prime}$. By definition, the Conley–Zehnder index $\mu(\Phi)\in{\mathbb{Z}}$ of $\Phi$ is $\operatorname{\hat{\mu}}(\Phi^{\prime})$. One can show that $\mu(\Phi)$ is well-defined, i.e., independent of $\Psi$. The function $\mu\colon\widetilde{\operatorname{Sp}}{}^{*}(2m)\to{\mathbb{Z}}$ is locally constant, i.e., constant on connected components of $\widetilde{\operatorname{Sp}}{}^{*}(2m)$.
The mean index $\operatorname{\hat{\mu}}(\Phi)$ measures the total rotation angle of certain unit eigenvalues of $\Phi(t)$. For instance, for the path $\Phi(t)=\exp\big(2\pi\sqrt{-1}\lambda t\big)$, $t\in[0,\,1]$, in $\operatorname{Sp}(2)$ we have
$\operatorname{\hat{\mu}}(\Phi)=2\lambda\textrm{ and }\mu(\Phi)=\mathit{sign}(\lambda)\big(2\lfloor|\lambda|\rfloor+1\big)\textrm{ when }\lambda\not\in{\mathbb{Z}},$ (3.1)
where $\mathit{sign}(\lambda)$ stands, predictably, for the sign of $\lambda$.
The upper and lower Conley–Zehnder indices are defined as
[…]
The mean index and the upper and lower indices are related by the inequalities
$\operatorname{\hat{\mu}}(\Phi)-m\leq\mu_{-}(\Phi)\leq\mu_{+}(\Phi)\leq\operatorname{\hat{\mu}}(\Phi)+m.$ (3.2)
Moreover, at least one of the inequalities is strict when $\Phi(1)$ is weakly non-degenerate, i.e., at least one of the eigenvalues is different from 1. As a consequence of (3.2), $\mu_{\pm}$ are quasimorphisms $\operatorname{\widetilde{Sp}}(2m)\to{\mathbb{Z}}$, and
$\lim_{k\to\infty}\frac{\mu_{\pm}(\Phi^{k})}{k}=\operatorname{\hat{\mu}}(\Phi);$
hence the name “mean index” for $\operatorname{\hat{\mu}}$.
Whenever $\Phi^{\prime}\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ and $\Phi^{\prime\prime}\in\widetilde{\operatorname{Sp}}(2m^{\prime\prime})$ we have the direct sum $\Phi=\Phi^{\prime}\oplus\Phi^{\prime\prime}\in\widetilde{\operatorname{Sp}}(2m)$, where $m=m^{\prime}+m^{\prime\prime}$, defined in a natural way. This is equivalent to that
$\Phi(1)=\Phi^{\prime}(1)\oplus\Phi^{\prime\prime}(1)\textrm{ and }\operatorname{\hat{\mu}}(\Phi)=\operatorname{\hat{\mu}}(\Phi^{\prime})+\operatorname{\hat{\mu}}(\Phi^{\prime\prime}).$
It is easy to see that any decomposition $\Phi(1)=A\oplus B$ with $A\in\operatorname{Sp}(2m^{\prime})$ and $B\in\operatorname{Sp}(2m^{\prime\prime})$ extends (non-uniquely) to a direct sum decomposition of $\Phi$.
We call $\Psi\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ the non-degenerate part of $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ when $\Psi(1)$ is non-degenerate, $\Phi(1)=\Psi(1)\oplus A$, with $m^{\prime}\leq m$ and $A\in\operatorname{Sp}\big(2(m-m^{\prime})\big)$ totally degenerate (i.e., with all eigenvalues equal to 1), and $\operatorname{\hat{\mu}}(\Psi)=\operatorname{\hat{\mu}}(\Phi)$. (The mean index condition is imposed here to make $\Psi$ unique and ensure that $\Psi=\Phi$ when $\Phi$ is non-degenerate.) When $\Phi$ is itself totally degenerate we say that the non-degenerate part of $\Phi$ is trivial. The non-degenerate part of $\Phi$ is a direct summand.
The indices $\operatorname{\hat{\mu}}$ and $\mu_{\pm}$ are additive under direct sum; see, e.g., [Lo97, SZ] and [GG20, Sec. 4.1]. Namely, for $\Phi^{\prime}\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ and $\Phi\in\widetilde{\operatorname{Sp}}(2m^{\prime\prime})$, we have
$\operatorname{\hat{\mu}}(\Phi^{\prime}\oplus\Phi^{\prime\prime})=\operatorname{\hat{\mu}}(\Phi^{\prime})+\operatorname{\hat{\mu}}(\Phi^{\prime\prime})\quad\textrm{and}\quad\mu_{\pm}(\Phi^{\prime}\oplus\Phi^{\prime\prime})=\mu_{\pm}(\Phi^{\prime})+\mu_{\pm}(\Phi^{\prime\prime}).$ (3.3)
Let us now briefly recall the definitions of several invariants of $\Phi$, which play a central role in the index recurrence theorem (Theorem 3.3). We refer the reader to [GG20, Sec. 4.1.3] for more details and proofs.
Consider first a totally degenerate symplectic linear map $A\in\operatorname{Sp}(2m)$. Then $A=\exp(JQ)$ where $J$ is the matrix of the standard linear symplectic form on ${\mathbb{R}}^{2n}$ and all eigenvalues of $JQ$ are equal zero. The quadratic form $Q$ can be symplectically decomposed into a sum of terms of four types:
  • the identically zero quadratic form on ${\mathbb{R}}^{2\nu_{0}}$,
  • the quadratic form $Q_{0}=p_{1}q_{2}+p_{2}q_{3}+\cdots+p_{d-1}q_{d}$ in Darboux coordinates on ${\mathbb{R}}^{2d}$, where $d\geq 1$ is odd,
  • the quadratic forms $Q_{\pm}=\pm(Q_{0}+p^{2}_{d}/2)$ on ${\mathbb{R}}^{2d}$ for any $d$.
These normal forms, which we will revisit in Section 5.2.1, can be found in [AG, Sec. 2.4] and [CLW, Sect. 2.6] and we find them more convenient to work with than the original Williamson normal forms; see [Wi36] and also, e.g., [Ar, App. 6].
Denoting by $\mathit{sgn}$ the signature, we have $\dim\ker Q_{0}=2$ and $\mathit{sgn}Q_{0}=0$, and $\dim\ker Q_{\pm}=1$ and $\mathit{sgn}Q_{\pm}=\pm 1$. Let $b_{*}(Q)$, where $0,\pm$, be the number of the $Q_{0}$ and $Q_{\pm}$ terms in the decomposition. Let us also set $b_{*}(A):=b_{*}(Q)$ and $\nu_{0}(A):=\nu_{0}(Q)$, and $b_{*}(\Phi):=b_{*}(A)$ when $A=\Phi(1)$ is totally degenerate. These are symplectic invariants of $Q$ and $A$ and $\Phi$, and the geometric multiplicity of the eigenvalue $1$ is $2(b_{0}+\nu_{0})+b_{+}+b_{-}$.
These definitions readily extend to all paths. Namely, every $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ can be written (non-uniquely) as a direct sum $\Phi_{0}\oplus\Psi$ where $\Phi_{0}(1)\in\operatorname{Sp}(2m_{0})$ is totally degenerate and $\Psi(1)\in\operatorname{Sp}(2m_{1})$ is non-degenerate. In particular, $m_{0}$ is the algebraic multiplicity of the eigenvalue 1 of $\Phi(1)$, and $m_{0}+m_{1}=m$. Then we set
$b_{*}(\Phi):=b_{*}\big(\Phi_{0}(1)\big)\textrm{ for $*=0,\pm$ and }\nu_{0}(\Phi):=\nu_{0}\big((\Phi_{0}(1)\big).$
These are symplectic invariants of $\Phi$. Furthermore, set
$\beta_{\pm}:=\nu_{0}+b_{0}+b_{\pm}.$
It readily follows from the definition that
$\beta_{\pm}\big(\Phi^{k}\big)=\beta_{\pm}(\Phi)$ (3.4)
for all $k\in{\mathbb{N}}$ when $\Phi$ is totally degenerate or, more generally, for admissible $k$, where by definition $\beta_{\pm}(\Phi):=\beta_{\pm}(\Phi_{0})$. (This is also true for $b_{*}$ and $\nu_{0}$.)
It is not hard to show that
$\mu_{\pm}(\Phi)=\mu(\Psi)+\operatorname{\hat{\mu}}(\Phi_{0})\pm\beta_{\pm}(\Phi);$ (3.5)
see [GG20, Sec. 4.1]. In particular, when $\Psi$ is the non-degenerate part of $\Phi$, i.e., in addition $\operatorname{\hat{\mu}}(\Phi_{0})=0$, we have
$\mu_{\pm}(\Phi)=\mu(\Psi)\pm\beta_{\pm}(\Phi),$ (3.6)
Furthermore, as readily follows from the definitions, the invariants $b_{*}$ and $\nu_{0}$ and $\beta_{\pm}$ are additive in the sense of (3.3). For instance, for $\Phi^{\prime}\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ and $\Phi\in\widetilde{\operatorname{Sp}}(2m^{\prime\prime})$, we have
$\beta_{\pm}(\Phi^{\prime}\oplus\Phi^{\prime\prime})=\beta_{\pm}(\Phi^{\prime})+\beta_{\pm}(\Phi^{\prime\prime}).$ (3.7)
Likewise,
$\beta_{\pm}\big(\Phi^{-1}\big)=\beta_{\mp}(\Phi).$
Finally, we note that for any $\Phi\in\widetilde{\operatorname{Sp}}(2m)$,
$|\beta_{+}(\Phi)-\beta_{-}(\Phi)|=|b_{+}(\Phi)-b_{-}(\Phi)|\leq m,$ (3.8)
which is essential for what follows.
3.1.2. Dynamical convexity
The notion of dynamical convexity was originally introduced in [HWZ] for Reeb flows on the standard contact $S^{3}$ and, somewhat in passing, for higher-dimensional contact spheres. In [AM17, AM22] the definition was extended to other contact manifolds. In this section we mainly focus on the linear algebra aspect of dynamical convexity and it is convenient for us to adopt the following definition.
Definition 3.1.
A path $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ is said to be dynamically convex (DC) if $\mu_{-}(\Phi)\geq m+2$.
A word of warning is due: the flow generated by a positive definite Hamiltonian need not be dynamically convex. This is the case for, instance, for the time-one flow of a small Hamiltonian; see [GG20, Ex. 4.7] for further discussion. The proof of the following result is quite standard and can be found in, e.g., [Lo02, SZ] and also [GG20, Sec. 4.2].
Lemma 3.2.
For any $\Phi\in\widetilde{\operatorname{Sp}}(2m)$, we have
$\mu_{-}\big(\Phi^{k+1}\big)\geq\mu_{-}\big(\Phi^{k}\big)+\big(\mu_{-}(\Phi)-m\big)$ (3.9)
for all $k\in{\mathbb{N}}$. In particular,
$\mu_{-}\big(\Phi^{k}\big)\geq(\mu_{-}(\Phi)-m)k+m.$
Assume, furthermore, that $\Phi$ is dynamically convex. Then the function $\mu_{-}(\Phi^{k})$ of $k\in{\mathbb{N}}$ is strictly increasing,
$\mu_{-}\big(\Phi^{k}\big)\geq 2k+m,$
and $\operatorname{\hat{\mu}}(\Phi)\geq\mu_{-}(\Phi)-m\geq 2$. Thus $\mu_{-}\big(\Phi^{k}\big)\geq m+2$ and all iterates $\Phi^{k}$ are also dynamically convex.
3.1.3. Conley–Zehnder index: closed orbits
In this section we briefly recall how the definition of the Conley–Zehnder index type invariants from Section 3.1.1 for elements of $\operatorname{\widetilde{Sp}}(2m)$ translates to closed orbits of a Reeb flow. Below all symplectic vector bundles are treated as Hermitian vector bundles.
Now let $M=\partial W$ be the boundary of a Liouville domain $(W^{2n\geq 4},d\alpha)$ such that $c_{1}(TW)=0$ as in Section 2.1 and let $\xi$ be the contact structure $\ker\alpha|_{M}$. Then $c_{1}(\xi)=0$ and hence the top complex wedge power $L_{M}$ of $\xi$ is a trivial complex line bundle over $M$. Likewise, the top complex wedge of $TW$ is a trivial line bundle $L$. Fix a unit section $\mathfrak{s}$} of $L$. Note that the symplectic orthogonal to $\xi$ in $T_{M}W$ is trivial and using, say, the Reeb vector field $R$ as a section of this bundle, we have an isomorphism $T_{M}W=\xi\oplus{\mathbb{C}}$ as bundles over $M$. Let ${\mathfrak{s}}_{M}$ be the unit section of $L_{M}$ such that ${\mathfrak{s}}={\mathfrak{s}}_{M}\otimes R$ over $M$. This section is completely determined by $\mathfrak{s}$}, up to homotopy.
Next, let $x\colon S_{T}^{1}:={\mathbb{R}}/T{\mathbb{Z}}\to M$ be a closed orbit of the Reeb flow, which we do not require to be prime. The pull-back ${\mathfrak{s}}_{M}\circ x$ of ${\mathfrak{s}}_{M}$ by $x$ is a non-vanishing section of the pull-back $x^{*}L_{M}$ over $S_{T}^{1}$ and $x^{*}L_{M}$ is in turn the top complex wedge of $x^{*}\xi$. Pick any Kähler trivialization of $x^{*}\xi$ which, up to homotopy, gives rise to the trivialization ${\mathfrak{s}}_{M}\circ x$ of $x^{*}L$, i.e., such that its top complex wedge is ${\mathfrak{s}}_{M}\circ x$. (Such a trivialization is unique up to homotopy.) Then we use this trivialization to turn the linearized Reeb flow $D\varphi^{t}$, $t\in[0,T]$, along $x$ into a path $\Phi(t)\in\operatorname{Sp}(2m)$ with m=n1m=n-1 starting at the identity, and hence an element $\Phi\in\operatorname{\widetilde{Sp}}(2m)$. By definition, $\mu_{\pm}(x):=\mu_{\pm}(\Phi)$ and $\operatorname{\hat{\mu}}(x):=\operatorname{\hat{\mu}}(\Phi)$. In particular, when $x$ is non-degenerate, $\mu(x):=\mu(\Phi)$. Furthermore, $\operatorname{\hat{\mu}}$ is homogeneous, i.e.,
$\operatorname{\hat{\mu}}\big(x^{k}\big)=k\operatorname{\hat{\mu}}(x),$
and
$\operatorname{\hat{\mu}}(x)-(n-1)\leq\mu_{-}(x)\leq\mu_{+}(x)\leq\operatorname{\hat{\mu}}(x)+(n-1).$
by (3.2). In a similar vein, we set $\beta_{\pm}(x):=\beta_{\pm}(\Phi)$, etc.
A Reeb flow is said to be dynamically convex if $\mu_{-}(x)\geq n+1$ for all closed Reeb orbits $x$. By Lemma 3.2, this is equivalent to that $\mu_{-}(x)\geq n+1$ for all prime orbits. This notion was introduced in [HWZ] where it was shown that the Reeb flow on a convex hypersurface in ${\mathbb{R}}^{2n}$ is dynamically convex. We refer the reader to, e.g., [GG20, Sec. 4.2] for a simple proof of this fact and to [ALM, AM17, AM22, DL2W18, GiM, GGM] for other relevant notions and results.
3.2. Index recurrence and applications
The original version of the index recurrence theorem from [GG20] asserts that arbitrarily long intervals in the sequence of the Conley–Zehnder indices of the iterates of one or a finite collection of symplectic linear maps repeat themselves infinitely many times up to a shift. (That theorem can also be derived in a non-trivial way from the index iteration formula, aka the common jump theorem, in [LZ]; we will return to the relation between the two results later in this section.) A more general variant of the theorem which we need here allows for control of not only the indices but also additional real parameters assigned to the maps, which should be thought of as periods of closed orbits. In addition, it shows that the sequence of “recurrence events” can be made to have a bounded gap.
Thus we consider pairs $(\Phi,a)$ where the “action” $a$ is real and positive, $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ and set $\Delta:=\operatorname{\hat{\mu}}(\Phi)>0$. Assume that we have a finite collection of such pairs $(\Phi_{i},a_{i})$, $i=1,\ldots,r$. Let us relabel and regroup these pairs putting two pairs in the same group, which we will call a cluster, if they have the same ratio $a/\Delta$. To be more specific, we will denote the pairs as $(\Phi_{ij},a_{ij})$, where $i=1,\ldots,i_{0}$ and $j=1,\ldots,j_{0}(i)$ with the index $i$ indicating the cluster, i.e., $i=i^{\prime}$ if and only if $a_{ij}/\Delta_{ij}=a_{i^{\prime}j^{\prime}}/\Delta_{i^{\prime}j^{\prime}}$. We will extend this division into clusters to “iterated” pairs $(\Phi^{k}_{ij},ka_{ij})$, $k\in{\mathbb{N}}$, by the same rule. (In particular, all iterates of a fixed pair are in the same cluster.) We call the union $\mathcal{S}$} of the sequences $a_{ij}{\mathbb{N}}$ the spectrum of the collection $(\Phi_{ij},a_{ij})$. In what follows, we denote by $[t]$ the closest integer to $t\in{\mathbb{R}}$ assuming that $t\not\in 1/2+{\mathbb{Z}}$.
The statement of our next result is admittedly notation heavy. It might be helpful for the reader to keep in mind that in what follows $l$ indexes recurrence events, $k$ with various subscripts and $\ell$ stand for the order of iteration, and $d_{il}$ is the index shift, i.e., an integer approximating the mean indices of the iterates $\Phi^{k_{ijl}}_{ij}$.
Theorem 3.3 (Index recurrence with action matching and bounded gap).
Let $(\Phi_{ij},a_{ij})$ be as above. Then for any $\eta>0$ and any $\ell_{0}\in{\mathbb{N}}$, there exists a sequence $C_{l}\to\infty$ of real numbers and integer sequences $d_{il}\to\infty$ and $k_{ijl}\to\infty$ (as $l\to\infty$) such that for all $i$, $j$ and $l$, and all $\ell\in{\mathbb{Z}}$ in the range $1\leq|\ell|\leq\ell_{0}$, we have
  • (IR1)
    $\big|\operatorname{\hat{\mu}}\big(\Phi^{k_{ijl}}_{ij}\big)-d_{il}\big|<\eta$ and, in particular, $d_{il}=\big[\operatorname{\hat{\mu}}\big(\Phi^{k_{ijl}}_{ij}\big)\big]$, and
    $d_{il}-m\leq\mu_{-}\big(\Phi^{k_{ijl}}_{ij}\big)\leq\mu_{+}\big(\Phi^{k_{ijl}}_{ij}\big)\leq d_{il}+m;$
  • (IR2)
    $\mu_{\pm}\big(\Phi^{k_{ijl}+\ell}_{ij}\big)=d_{il}+\mu_{\pm}\big(\Phi^{\ell}_{ij}\big)$ when $0<\ell\leq\ell_{0}$;
  • (IR3)
    $\mu_{+}\big(\Phi^{k_{ijl}-\ell}_{ij}\big)=d_{il}-\mu_{-}\big(\Phi^{\ell}_{ij}\big)+\big(\beta_{+}\big(\Phi^{\ell}_{ij}\big)-\beta_{-}\big(\Phi^{\ell}_{ij}\big)\big)$, where $0<\ell\leq\ell_{0}<k_{ijl}$ and $\beta_{+}-\beta_{-}=b_{+}-b_{-}\leq m$ with $\beta_{\pm}=0$ when $\Phi_{ij}^{\ell}$ is non-degenerate;
  • (IR4)
    Assertions (IR1)(IR3) continue to hold with the same values of $d_{il}$ and $k_{ijl}$ when $\Phi_{ij}$ is replaced by its non-degenerate part $\Psi$ with all invariants of $\Phi_{ij}$ and its iterates replaced by their counterparts for $\Psi$, i.e., $m$ replaced by $m^{\prime}$, $\mu_{\pm}\big(\Phi^{k_{ijl}}_{ij}\big)$ replaced by $\mu\big(\Psi^{k_{ijl}}\big)$, $\beta_{\pm}$ set to be 0, etc.;
  • (IR5)
    $C_{l}-\eta<k_{ijl}a_{ij}<C_{l}$, and $ka_{ij}<C_{l}-\eta$ when $k<k_{ijl}$ and $ka_{ij}>C_{l}$ when $k>k_{ijl}$.
Moreover, we can ensure that the sequences $C_{l}$, $d_{il}$ and $k_{ijl}$ are strictly increasing and have bounded gap. In addition, we can make all $k_{ijl}$ and $d_{il}$ divisible by any pre-assigned integer and all $C_{l}$ lie outside the spectrum $\mathcal{S}$}.
In what follows we will refer to the collection $\big\{k_{ijl},d_{il},C_{l}\big\}$ for a fixed $l$ as a recurrence event.
Remark 3.4.
Condition (IR4) in Theorem 3.3 and its corollaries can be generalized a bit further. Namely, (IR1)(IR3) continue to hold with the same values of $d_{il}$ and $k_{ijl}$ when $\Phi_{ij}$ is replaced by any $\Psi\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ such that $\Phi(1)=\Psi(1)\oplus A$, where $m^{\prime}\leq m$ and $A\in\operatorname{Sp}\big(2(m-m^{\prime})\big)$ and $\operatorname{\hat{\mu}}(\Psi)=\operatorname{\hat{\mu}}(\Phi)$. As in (IR4), we also need to replace $m$ by $m^{\prime}$, $\mu_{\pm}\big(\Phi^{k_{ijl}}_{ij}\big)$ by $\mu_{\pm}\big(\Psi^{k_{ijl}}\big)$, $\beta_{\pm}\big(\Phi^{k_{ijl}}_{ij}\big)$ by $\beta_{\pm}\big(\Psi^{k_{ijl}}\big)$, etc. This fact readily follows from the proof of the theorem.
Let us now state some of the corollaries of Theorem 3.3 and show how this theorem is related to several previous results of the same type.
Corollary 3.5.
In the setting of Theorem 3.3, assume that all maps $\Phi_{ij}$ are dynamically convex and $\ell_{0}$ is taken large enough. (For instance, $2\ell_{0}\geq 3(m+1)$.) Then as a formal consequence of (IR2), (IR3) and (IR5), we have
  • (IR2’)
    $\mu_{-}\big(\Phi^{k}_{ij}\big)\geq d_{il}+m+2$ and $ka_{ij}>C_{l}$ when $k>k_{ijl}$;
  • (IR3’)
    $\mu_{+}\big(\Phi^{k}_{ij}\big)\leq d_{il}-2$ and $ka_{ij}<C_{l}-\eta$ when $k<k_{ijl}$, and $\mu\big(\Phi^{k}_{ij}\big)\leq d_{il}-m-2$ when in addition $\Phi_{ij}$ is strongly non-degenerate.
In particular, for any $i$, none of the intervals $\big[\mu_{-}\big(\Phi^{k}_{ij}\big),\mu_{+}\big(\Phi^{k}_{ij}\big)\big]$, $k\in{\mathbb{N}}$, contains the point $d_{il}+m+1$. When $\Phi_{ij}$ is in addition strongly non-degenerate this is also true for $d_{il}-m-1$. Furthermore, $ka_{ij}\in(C_{l}-\eta,C_{l})$ only when $k=k_{ij}$.
Proof.
Only the index part in (IR2’) and (IR3’) requires a proof. The “action” part follows directly from (IR5). By (IR1), the “In particular” part of the corollary is essentially a consequence of (IR2’) and (IR3’).
Recall from Lemma 3.2 that $\operatorname{\hat{\mu}}(\Phi_{ij})\geq 2>0$ due to the dynamical convexity condition. Fix $\ell_{0}$ so that
$\ell_{0}\min_{ij}\operatorname{\hat{\mu}}(\Phi_{ij})\geq 2\ell_{0}\geq 3(m+1).$
Then, when $k>k_{ijl}+\ell_{0}$, by the triangle inequality and (3.2),
$\begin{split}\mu_{-}(\Phi^{k}_{ij})-d_{il}&\geq\big(\operatorname{\hat{\mu}}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})\big)-\big|\mu_{-}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k}_{ij})\big|-\big|\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})-d_{il}\big|\\ &\geq(k-k_{ijl})\operatorname{\hat{\mu}}(\Phi_{ij})-m-\eta\\ &\geq 2(k-k_{ijl})-m-1\\ &\geq 2m+2.\end{split}$
When $k=k_{ijl}+\ell$ with $1\leq\ell\leq\ell_{0}$, we have
$\mu_{-}\big(\Phi^{k}_{ij}\big)=d_{il}+\mu_{-}\big(\Phi^{\ell}_{ij}\big)\geq d_{il}+m+2$
by (IR2) and the dynamical convexity assumption. This proves (IR2’).
Condition (IR3’) is established by a similar calculation. Namely, for $k<k_{ijl}-\ell_{0}$,
$\begin{split}\mu_{+}(\Phi^{k}_{ij})-d_{il}&\leq\big(\operatorname{\hat{\mu}}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})\big)+\big|\mu_{+}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k}_{ij})\big|+\big|\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})-d_{il}\big|\\ &\leq(k-k_{ijl})\operatorname{\hat{\mu}}(\Phi_{ij})+m+\eta\\ &\leq 2(k-k_{ijl})+m+1\\ &\leq-2m-2.\end{split}$
When $k=k_{ijl}$-$\ell$ with $1\leq\ell\leq\ell_{0}$, we have
$\begin{split}\mu_{+}\big(\Phi^{k}_{ij}\big)&=d_{il}-\mu_{-}\big(\Phi^{\ell}_{ij}\big)+\big(\beta_{+}\big(\Phi^{\ell}_{ij}\big)-\beta_{-}\big(\Phi^{\ell}_{ij}\big)\big)\\ &\leq d_{il}-m-2+m\\ &=d_{il}-2\end{split}$
by (IR3) and again the dynamical convexity assumption, proving (IR3’). When $\Phi_{ij}$ is strongly non-degenerate, the term $\beta_{+}-\beta_{-}$ disappears and we obtain the upper bound $d_{il}-m-2$. ∎
Setting $a_{ij}:=\operatorname{\hat{\mu}}(\Phi_{ij})$, we force all maps $\Phi_{ij}$ to be in one cluster, and then we can suppress $i$ in the notation, labeling the maps as $\Phi_{j}$. In particular, $d_{il}$ becomes independent of $i$. Thus, the above results immediately yield the following.
Corollary 3.6.
In the setting of Theorem 3.3 or Corollary 3.5, for a collection of maps $\Phi_{j}\in\widetilde{\operatorname{Sp}}(2m)$, we can find bounded gap sequences $C_{l}$ and $d_{l}$ independent of $j$, and bounded gap sequences $k_{jl}$ such that all other conditions are satisfied. In particular, the sequences $k_{jl}$ and $d_{l}$ can be made divisible by any fixed integer. Furthermore, with $\Delta_{j}:=\operatorname{\hat{\mu}}(\Phi_{j})$ condition (IR5) takes the following form: $C_{l}-\eta<k_{jl}\Delta_{j}<C_{l}$, and $k\Delta_{j}<C_{l}-\eta$ when $k<k_{jl}$ and $k\Delta_{j}>C_{l}$ when $k>k_{jl}$.
Thus, in this case, a recurrence event takes a simpler form $\big\{k_{jl},d_{l},C_{l}\big\}$. In fact, the sequence $C_{l}$ is somewhat redundant now and carries essentially the same information as the sequence $d_{l}:=[k_{jl}\Delta_{j}]$.
This corollary is close to the original version of the index recurrence theorem from [GG20, Thm. 5.2]. This theorem is in turn closely related to and can be derived from the common jump theorem from [Lo02, LZ] or from its further refined variant proved in [DLW], which is done in detail in [DLWZ]. The new point of Corollary 3.6 as compared to these results is the part that all sequences involved have bounded gap. A simple proof from [GG20] of the following multiplicity result, quite standard now, is an illustrative application of the index recurrence theorem in the form of Corollaries 3.5 and 3.6. Interestingly, the argument relies on the existence of just one recurrence event.
Corollary 3.7 (Multiplicity; [LZ] and [GK]).
Assume that the Reeb flow on the boundary $M$ of a star-shaped domain $W\subset{\mathbb{R}}^{2n}$ is non-degenerate and dynamically convex. Then the flow has at least $n$ non-alternating prime closed orbits with index of the same parity as $n+1$.
Proof.
Without loss of generality, we may assume that there are only finitely many prime closed orbits $x_{i}$ in $M$. Denote the linearized flows along these orbits by $\Phi_{1},\ldots,\Phi_{r}$. Consider one recurrence event from Corollary 3.6 for these elements of $\operatorname{Sp}(2(n-1))$. Suppressing $l$ in the notation and ignoring $C_{l}$, we denote this event by $\{d,k_{1},\ldots,k_{r}\}$ and require $d$ and all $k_{j}$ to be divisible by two. Since $\operatorname{CH}_{n-1+2i}(W;{\mathbb{Q}})\neq 0$, $i\in{\mathbb{N}}$, there must be at least one closed orbit with index n1+2in-1+2i. There are exactly $n$ such integer points in the range ${\mathcal{I}}:=[d-n+1,d+n-1]$. By Corollary 3.5, the only closed Reeb orbits with index in $\mathcal{I}$} are $x_{j}^{k_{j}}$. Moreover, since $k_{j}$ is even, to contribute to the symplectic homology, $x_{j}$ must be non-alternating and hence, the parity of $\mu(x_{j})$ must be the same as that of $\mu\big(x_{j}^{k_{j}}\big)$, i.e., $n+1$. Therefore, there are at least $n$ non-alternating prime closed orbits with parity $n+1$. ∎
A more subtle application allows one to remove the non-degeneracy condition when $W$ is centrally symmetric and convex.
Corollary 3.8 (Multiplicity in the convex, symmetric case; [LLZ]).
Assume that $W$ is centrally symmetric and strictly convex. Then the Reeb flow on $M=\partial W$ has at least $n$ prime closed orbits.
Outline of the proof following [GiM].
As in the proof of Corollary 3.7, assume that the flow has only finitely many prime closed orbits. Keeping the notation from that proof, denote these orbits by $x_{j}$. The key observation is that with our conventions $\mu_{-}\big(x_{i}^{k}\big)-\beta_{+}\big(x_{i}^{k}\big)+\beta_{-}\big(x_{i}^{k}\big)\geq n+1$ for all $k\in{\mathbb{N}}$ as a consequence of symmetry and convexity; see [GiM, Sec. 4], [LLZ, Lemma 4.1] and [Lo02, Lemma 15.6.3]. (This fact is formalized in [GiM] as a symplectically invariant notion of strong dynamical convexity.) Considering an index recurrence event, we now see that $\operatorname{supp}\operatorname{CH}\big(x_{j}^{k};{\mathbb{Q}}\big)$ does not overlap with the interval $\mathcal{I}$} unless $k=k_{j}$ which we can assume to be even. Using the $S^{1}$-equivariant spectral invariants over $\mathbb{Q}$} as in [GG20], we conclude that among the prime orbits $x_{i}$ there must be at least $n$ orbits to fill in all $n$ spots in $\mathcal{I}$}. Alternatively, we can use the fact that, by Corollary D, $\operatorname{supp}\operatorname{CH}\big(x_{j}^{k};{\mathbb{Q}}\big)$ contains at most one point for all $k\in{\mathbb{N}}$. ∎
Remark 3.9.
As we have already mentioned in the introduction, Corollary 3.7 had been generalized and strengthened in several directions. When the non-degeneracy condition is removed, a proof similar to the above one but relying on equivariant spectral invariants gives a lower bound $\lfloor n/2\rfloor+1$. The argument can be further pushed (in a non-trivial way) to improve this lower bound to $\lceil n/2\rceil+1$; see [Wa16b] and also [DL, GG20]. In dimensions 6 and 8, the lower bound is 3 and, respectively 4; see [DL, Wa16a, WHL].) Furthermore, for non-degenerate flows, the dynamical convexity condition has been replaced by a much less restrictive index requirement, [DL2W18, DL2W24, GGM]. Corollary 3.8 has also been refined and generalized in a variety of ways; see, e.g., [AM22, ALM, GiM] and references therein.
4. Proofs of the main results
Here we first derive Theorems A and B from Corollary D and Theorem E in Section 4.1. Then, in Section 4.2, we establish the latter as a consequence of Theorem C. In Section 4.3, we turn to the proofs of Theorem C, which is the central part of the argument, and Corollary D. Finally, Corollary F is proved in Section 4.4
4.1. Proofs of Theorems A and B
The proof of Theorem A comprises two unequal parts with most of the work done in the proof of the first part. Theorem B is proved in Section 4.1.2.
4.1.1. Proof of Theorem A
Our proof of the theorem actually establishes a more precise result asserting the existence of exactly $n$ prime orbits satisfying certain visibility conditions. Although this result is more technical, we state and prove it here, for it sheds more light on the proof and is of independent interest.
Let $y$ be a prime orbit and let $p(y)$ be the least common multiple of the degrees of the roots of unity among the elliptic eigenvalues of the return map of $y$. Thus $y^{p(y)}$ is a “maximally degenerate” iterate of $y$. Note that $y^{p(y)}$ can fail to be $\mathbb{Q}$}-invisible even when $y$ is elliptic and non-degenerate; see Example 2.12. Set $N(y):=2p(y)$. We will call $y$ virtually $\mathbb{Q}$}-visible when $y^{N(y)}$ is $\mathbb{Q}$}-visible. For instance, a strongly non-degenerate prime orbit is virtually $\mathbb{Q}$}-visible if and only if it is non-alternating. A totally degenerate $\mathbb{Q}$}-visible prime orbit is virtually $\mathbb{Q}$}-visible. The orbit from Example 2.12 is not. As of writing, it is not known if a virtually $\mathbb{Q}$}-visible prime orbit must be $\mathbb{Q}$}-visible even in the setting of Theorem A.
Theorem 4.1.
Assume that the Reeb flow on the boundary $M^{2n-1}\subset{\mathbb{R}}^{2n}$ of a star-shaped domain is dynamically convex and has finitely many prime closed orbits. Then the flow has exactly $n$ virtually $\mathbb{Q}$}-visible prime orbits. Moreover, these orbits $y$ have the same ratio ${\mathcal{A}}(y)/\operatorname{\hat{\mu}}(y)$.
Remark 4.2.
The $n$ virtually $\mathbb{Q}$}-visible closed orbits from Theorem 4.1 have an additional property that $\operatorname{\hat{\mu}}(y)/\operatorname{\hat{\mu}}(y^{\prime})\not\in{\mathbb{Q}}$ for any two such distinct orbits $y$ and $y^{\prime}$. Indeed, assume that this ratio is rational: $\operatorname{\hat{\mu}}(y)/\operatorname{\hat{\mu}}(y^{\prime})=p/q$ for some positive integers $p$ and $q$. Let $m$ be a common multiple of $N(y)$ and $N(y^{\prime})$. Then, by (4.1) below, $y^{qm}$ and $(y^{\prime})^{pm}$ are both $\mathbb{Q}$}-visible. The iterates $y^{qm}$ and $(y^{\prime})^{pm}$ have the same action (aka period), which contradicts Corollary D. This observation, added after we have learned about [Wa25], sharpens the main result from that paper, which in turn is a refinement of a theorem from [HO].
Part (M1) of Theorem A immediately follows from Theorem 4.1. To prove (M2), recall that in the non-degenerate case a prime orbit is virtually $\mathbb{Q}$}-visible if and only if it is non-alternating. Hence there are exactly $n$ such orbits. The parity statement is a consequence Part (O1) from Theorem E since now the degree is equal to the Conley–Zehnder index. As we have already mentioned, the lower bound in (M2) is not new; cf. Corollary 3.7 and [GG20, GK, LZ].
Proof of Theorem 4.1.
The “Moreover” part follows immediately from (O2) in Theorem E. We begin with some general remarks on virtually $\mathbb{Q}$}-visible orbits. Then we turn to the multiplicity statement which comprises the main part and the bulk of the proof.
Let $y$ be a closed orbit. Then, in the setting of the theorem,
[…]
In other words, $y$ is virtually $\mathbb{Q}$}-visible if and only if there exists $k\in{\mathbb{N}}$ such that $y^{kN(y)}$ is $\mathbb{Q}$}-visible, and if and only if $y^{kN(y)}$ is $\mathbb{Q}$}-visible for all $k\in{\mathbb{N}}$. (In particular, again within the theorem’s framework, for any orbit $z$ which is not virtually $\mathbb{Q}$}-visible, no iterate $z^{kN(y)}$, $k\in{\mathbb{N}}$, is $\mathbb{Q}$}-visible.)
Indeed, $y^{kN(y)}$ is an admissible iterate of $y^{N(y)}$ and the return map of $y^{N(y)}$ has no real negative eigenvalues by the definition of $N(y)$. Hence,
$\chi^{\scriptscriptstyle{eq}}\big(y^{kN(y)}\big)=\chi^{\scriptscriptstyle{eq}}\big(y^{N(y)}\big)$
for all $k\in{\mathbb{N}}$ by Lemma 2.8. Under the conditions of the theorem, $\chi^{\scriptscriptstyle{eq}}(x)$ is either $(-1)^{n+1}$ or 0 depending on whether $x$ is $\mathbb{Q}$}-visible or not, as follows from Corollary D. Hence, $y^{N(y)}$ and $y^{kN(y)}$ are both either $\mathbb{Q}$}-visible or $\mathbb{Q}$}-invisible.
The strategy of the proof of the multiplicity part is to use the index recurrence theorem to identify an interval ${\mathcal{I}}:=[d-n+1,d+n-1]\subset{\mathbb{N}}$ such that for each eventually $\mathbb{Q}$}-visible prime orbit $x_{j}$ only one iterate $x_{j}^{k_{j}}$ can have degree in $\mathcal{I}$}. Moreover, we ensure that $x_{j}^{k_{j}}$ is $\mathbb{Q}$}-visible if and only if $x_{j}$ is virtually $\mathbb{Q}$}-visible and thus there are exactly $n$ such orbits. The main difficulty lies in showing that $\deg(x_{j}^{k<k_{j}})\not\in{\mathcal{I}}$ and this is where we crucially rely on Corollary D and Theorem E. We prove this first for the least action orbit $x$ by analyzing the behavior of the sequence $\deg(x^{k})$. Dealing with the iterates of the remaining prime orbits $x_{j}$, we show that $x_{j}^{k}$ must be $\mathbb{Q}$}-invisible whenever $\deg(x_{j}^{k})\in{\mathcal{I}}$ and $k<k_{j}$, by comparing the degree and the action of $x_{j}^{k}$ with the degree and the action of the iterates of $x$. It is convenient to break down the proof into three steps.
Step 1: Setup. As in Theorem E, denote by $\mathcal{P}$} the set of all $\mathbb{Q}$}-visible closed orbits. Let $x\in{\mathcal{P}}$ be the orbit with the least action. Then $\deg(x)=n+1$ since the ordering of $\mathcal{P}$} by the action agrees with the ordering by the degree due to Part (O1) of that theorem. As a consequence, $\mu_{-}(x)\leq n+1$. We claim that $x$ is prime. Indeed, if $x=z^{k}$ with $k\geq 2$, by Lemma 3.2 and dynamical convexity we have
$\mu_{-}(x)=\mu_{-}\big(z^{k}\big)>n+1,$
which is impossible.
Remark 4.3.
When $M$ is convex, $x$ is, in fact, a least period closed orbit on $M$, among all (not necessarily $\mathbb{Q}$}-visible) closed orbits. In other words, at least one of the least period closed orbits is necessarily $\mathbb{Q}$}-visible. This is a consequence of the fact that for convex hypersurfaces the (non-equivariant) homological capacity is equal to the least period, [AbK, Ir].
Next, let $x_{0}:=x,\,x_{1},\ldots,x_{r}$ be all prime closed orbits which are eventually $\mathbb{Q}$}-visible, i.e., $x_{j}^{k}\in{\mathcal{P}}$ for some $k\in{\mathbb{N}}$. Note that the orbits $x_{j\geq 1}$ might be $\mathbb{Q}$}-invisible and/or have action below ${\mathcal{A}}(x)$. In fact, $x_{j}^{s}$ with $j\geq 1$ is necessarily $\mathbb{Q}$}-invisible when ${\mathcal{A}}\big(x_{j}^{s}\big)\leq{\mathcal{A}}(x)$. For each $j\geq 1$, we denote by $s_{j}$ the largest such $s$, setting $s_{j}=0$ if ${\mathcal{A}}(x_{j})>{\mathcal{A}}(x)$. Our goal is to show that there are exactly $n$ virtually $\mathbb{Q}$}-visible orbits among the orbits $x_{j}$.
By (O2) and homogeneity of $\mathcal{A}$} and $\operatorname{\hat{\mu}}$, the orbits $x_{j}$, $j\geq 0$, form one cluster, i.e., have the same ratio ${\mathcal{A}}(x_{j})/\operatorname{\hat{\mu}}(x_{j})$. Without loss of generality, we may assume that this ratio is 1, i.e.,
${\mathcal{A}}(y)=\operatorname{\hat{\mu}}(y)$ (4.2)
for all $y\in{\mathcal{P}}$.
Consider an index recurrence event as in Corollary 3.6, where we require $\eta>0$ to be sufficiently small, all $d_{l}$ to be even and $k_{jl}$ divisible by $N\in{\mathbb{N}}$ specified as follows. Let $p$ be the least common multiple of the degrees of the roots of unity among the elliptic eigenvalues of all closed orbits $x_{j}$. Then
$N=2p\prod_{j=1}^{r}s_{j}!$ (4.3)
where, as usual, we set $0!=1$. (Since there is only one cluster we can suppress $i$ in the notation.) Furthermore, we will also drop $l$ from the notation, concentrating on one event. Thus $d=d_{l}$ and $k_{j}=k_{jl}$, etc.
We have three groups (not as algebra objects) of $\mathbb{Q}$}-visible closed orbits $x_{j}^{k}$:
  • Group $\Gamma_{+}$ formed by the $\mathbb{Q}$}-visible orbits $x_{j}^{k>k_{j}}$. Then $\mu_{-}\big(x_{j}^{k}\big)\geq d+n+1$, and hence $\deg\big(x_{j}^{k}\big)\geq d+n+1$.
  • Group $\Gamma_{0}$ comprising the $\mathbb{Q}$}-visible orbits $x_{j}^{k_{j}}$. Then
    $d-n+1\leq\mu_{-}\big(x_{j}^{k_{j}}\big)\leq\mu_{+}\big(x_{j}^{k_{j}}\big)\leq d+n-1,$
    and hence
    $\deg\big(x_{j}^{k_{j}}\big)\in{\mathcal{I}}:=[d-n+1,d+n-1]\subset{\mathbb{N}}.$
  • Group $\Gamma_{-}$ formed by the $\mathbb{Q}$}-visible orbits $x_{j}^{k<k_{j}}$. Then $\mu_{+}\big(x_{j}^{k}\big)\leq d-2$, and hence $\deg\big(x_{j}^{k}\big)\leq d-2$.
It is enough to show that the orbits from $\Gamma_{-}$ cannot have degree in $\mathcal{I}$}, i.e.,
$\deg\big(x_{j}^{k}\big)\leq d-n\textrm{ when }k<k_{j}\textrm{ and $x_{j}^{k}\in{\mathcal{P}}$}.$ (4.4)
Indeed, then the only closed orbits with degree in $\mathcal{I}$} must be from $\Gamma_{0}$. By (4.1), since $N$ is divisible by all $N(x_{j})$, an orbit $x_{j}^{k_{j}}$ is $\mathbb{Q}$}-visible if and only if $x_{j}$ is virtually $\mathbb{Q}$}-visible. There are exactly $n$ integer spots in $\mathcal{I}$} of parity $n+1$, which must all be filled. Different $\mathbb{Q}$}-visible orbits have different degrees by Theorem E. Therefore, there are exactly $n$ virtually $\mathbb{Q}$}-visible prime orbits.
Up to this point our reasoning has followed closely the proof of multiplicity results in [GG20, Sec. 6]. The rest of the argument is new and different and critically relies on Theorem E, Corollary D and Lemma 2.8 precisely describing the behavior of the equivariant Euler characteristic $\chi^{\scriptscriptstyle{eq}}$ under admissible iterations.
Step 2: The orbit $x_{0}=x$. To establish (4.4), we start with $j=0$, i.e., $x_{0}=x$, and $k=k_{0}-1$. Then $k$ is admissible and odd due to our choice of $N$. Thus
$\chi^{\scriptscriptstyle{eq}}(x^{k})=\chi(x)=(-1)^{n+1}\neq 0.$
Here in the first equality we used Lemma 2.8. The second equality follows from the fact that $\operatorname{CH}(x;{\mathbb{Q}})$ is concentrated in degree $n+1$ by Corollary D. As a consequence, $x^{k}\in{\mathcal{P}}$. We will need the following observation.
Lemma 4.4.
Let $x$ and $k$ be as above.
  • (D1)
    Then
    $\deg(x^{k})=\deg(x)+(k-1)\operatorname{\hat{\mu}}(x)$
    if $x$ is totally degenerate. (In fact, this is true for all $k\in{\mathbb{N}}$.)
  • (D2)
    When $x$ is not totally degenerate, let $\Psi$ be the non-degenerate part of $x$. Then
    $\deg(x^{k})=\deg(x)+\big[\mu\big(\Psi^{k}\big)-\mu(\Psi)\big].$
Proof.
We already know that $\operatorname{CH}(x;{\mathbb{Q}})$ and $\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)$ are supported in one degree and one-dimensional. Hence we only need to determine the degree shift $L$. Let $\varphi$ be the Poincaré return map of $x$. Then the return map of $x^{l}$ is $\varphi^{l}$ for all $l\in{\mathbb{N}}$. By (2.10) and (2.17),
$\dim\operatorname{CH}\big(x^{l};{\mathbb{Q}}\big)\leq\dim\operatorname{HF}\big(\varphi^{l};{\mathbb{Q}}\big),$
again for all $l\in{\mathbb{N}}$, and
$\dim\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)=1$
when $k$ is admissible. Moreover, $\operatorname{CH}(x;{\mathbb{Q}})=\operatorname{HF}(\varphi;{\mathbb{Q}})$ due to [Fe20, Thm. 1.2]; see also [Fe21]. Furthermore, as readily follows from [GG10, Thm. 1.1] (and its proof),
$\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)=\operatorname{HF}(\varphi;{\mathbb{Q}})[-L]=\operatorname{CH}(x;{\mathbb{Q}})[-L]$
since $k$ is admissible. Here the shift $L$ is exactly as in (D1) and (D2) depending on whether $x$ is totally degenerate or not. (Indeed, in the former case $\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)=\operatorname{HF}(\psi\varphi;{\mathbb{Q}})$ for a loop $\psi$ of local Hamiltonian diffeomorphisms with $\operatorname{\hat{\mu}}(\psi)=(k-1)\operatorname{\hat{\mu}}(x)$ by [GG10, Claim 4.1]. Hence, $\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)=\operatorname{HF}(\varphi;{\mathbb{Q}})[-(k-1)\operatorname{\hat{\mu}}(x)]$ by [GG10, (LF6)]. The latter case follows from [GG10, Sect. 4.5].) Combining these facts we conclude that $\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)$ is one-dimensional and supported in $\deg(x^{k})$, and thus $\deg(x^{k})=\deg(x)+L$. ∎
Remark 4.5.
Lemma 4.4 is a particular case of a much more general hypothetical statement. Namely, we conjecture that for any isolated closed Reeb orbit $x$,
$\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)=\operatorname{CH}(x;{\mathbb{Q}})[(1-k)\operatorname{\hat{\mu}}(x)]$
if $x$ is totally degenerate. When $x$ is not totally degenerate, let as above $\Psi$ be the non-degenerate part of $x$. Then, at least if $x$ is prime,
$\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)=\operatorname{CH}(x;{\mathbb{Q}})\big[\mu(\Psi)-\mu\big(\Psi^{k}\big)\big]$
where $k$ is admissible and odd. For non-equivariant symplectic homology we expect similar statements to be true for any field. A variant of this conjecture is proved in [HHM, Thm.s 1.10 and 1.11] with a different definition of the local homology, which is hypothetically equivalent to the one used here.
We are now in a position to return to the proof of (4.4) for $x_{0}=x$ and $k=k_{0}-1$. We will prove the strict inequality
$\deg\big(x^{k_{0}-1}\big)<d-n.$ (4.5)
There are two cases to consider depending on whether $x$ is totally degenerate or not.
Assume first that $x$ is totally degenerate. In this case $x^{k}\in{\mathcal{P}}$ for all $k\in{\mathbb{N}}$. Let us define $q$ by the condition
$n+1=\deg(x)=\operatorname{\hat{\mu}}(x)+q.$
Note that by dynamical convexity and (3.5) with $\Psi$ absent we have
$n+1\leq\mu_{-}(x)=\operatorname{\hat{\mu}}(x)-\beta_{-}(x),\textrm{ where }\beta_{-}(x)\geq 0.$
Thus,
$q\leq 0\quad\textrm{and}\quad\operatorname{\hat{\mu}}(x)\geq n+1.$ (4.6)
It follows from (D1) of Lemma 4.4 that
$\deg\big(x^{k}\big)=k\operatorname{\hat{\mu}}(x)+q.$
(In fact this is true for all $k\in{\mathbb{N}}$ since all $k$ are admissible in this case.) Applying this to $k=k_{0}-1$ and using (4.6) in the first inequality and then (IR1) from Theorem 3.3 in the penultimate one, we have
$\begin{split}\deg\big(x^{k_{0}-1}\big)&=(k_{0}-1)\operatorname{\hat{\mu}}(x)+q\\ &=\operatorname{\hat{\mu}}\big(x^{k_{0}}\big)-\operatorname{\hat{\mu}}(x)+q\\ &\leq\operatorname{\hat{\mu}}\big(x^{k_{0}}\big)-(n+1)\\ &\leq d+\eta-(n+1)\\ &<d-n.\end{split}$
In the last inequality, we use the fact that $\eta$ is small and, in particular, $\eta<1$. (In fact, in this case $d=\operatorname{\hat{\mu}}\big(x^{k_{0}}\big)$ because $x$ is totally degenerate, and hence $\operatorname{\hat{\mu}}\big(x^{k_{0}}\big)\in{\mathbb{Z}}$.) Thus $\deg\big(x^{k}\big)<d-n$ proving (4.5) when $x$ is totally degenerate.
Next, assume that $x$ has a non-trivial non-degenerate part which we denote by $\Psi$ as above. The argument is quite similar to the totally degenerate case with $\operatorname{\hat{\mu}}(x)$ replaced by $\mu(\Psi)$. Namely, note that $k:=k_{0}-1$ is admissible and odd, due to our divisibility assumptions on $k_{0}$; see (4.3). Define $q$ by the condition
$n+1=\deg(x)=\mu(\Psi)+q.$
By dynamical convexity and (3.5) or more specifically (3.6),
$n+1\leq\mu_{-}(x)=\mu(\Psi)-\beta_{-}(x)\textrm{ where }\beta_{-}(x)\geq 0.$
It follows that
$n+1=\mu(\Psi)+q\leq\mu(\Psi)-\beta_{-}(x)\textrm{ with }\beta_{-}(x)\geq 0,$
and hence
$q\leq 0\quad\text{and}\quad\mu(\Psi)\geq n+1.$ (4.7)
Now by (D2) with $k=k_{0}-1$, we have
$\deg\big(x^{k}\big)=\mu(\Psi)+q+\mu\big(\Psi^{k}\big)-\mu(\Psi)=\mu\big(\Psi^{k}\big)+q.$
Applying Parts (IR3) and (IR4) of Theorem 3.3 to $\Psi^{k_{0}-1}=\Psi^{k}$, we have
$\mu\big(\Psi^{k}\big)=d-\mu(\Psi),$
Putting these two facts together and using (4.7) we obtain that
$\deg\big(x^{k}\big)=d-\mu(\Psi)+q\leq d-n-1.$
Therefore, $\deg\big(x^{k}\big)<d-n$, i.e., (4.5) holds when the non-degenerate part of $x$ is non-trivial.
When $j=0$, i.e., $x_{0}=x$, and $k=k_{0}$-$\ell$ with $\ell>1$, condition (4.4) is an easy consequence of (4.5). Indeed, then
${\mathcal{A}}\big(x^{k}\big)<{\mathcal{A}}\big(x^{k_{0}-1}\big).$
When in addition $x^{k}\in{\mathcal{P}}$, i.e., $x^{k}$ is $\mathbb{Q}$}-visible, we must have
$\deg\big(x^{k}\big)<\deg\big(x^{k_{0}-1}\big)<d-n$
by Part (O1) of Theorem E and (4.5). This proves (4.4) for $j=0$.
Step 3: The orbits $x_{j\geq 1}$. Next, let us focus on (4.4) for $j\geq 1$. Fixing $j$, assume that (4.4) fails:
$x_{j}^{k_{j}-s}\in{\mathcal{P}}\textrm{ and }\deg\big(x_{j}^{k_{j}-s}\big)>d-n\textrm{ for some }s>0.$
Then we also have
$\deg\big(x_{j}^{k_{j}-s}\big)>\deg\big(x^{k_{0}-1}\big),$
and hence
${\mathcal{A}}\big(x_{j}^{k_{j}-s}\big)>{\mathcal{A}}\big(x^{k_{0}-1}\big)$ (4.8)
as a consequence again of (O1).
By (4.2), the action is equal to the mean index. By Part (IR1) of Theorem 3.3,
$\big|{\mathcal{A}}\big(x_{j}^{k_{j}}\big)-{\mathcal{A}}\big(x^{k_{0}}\big)\big|<2\eta.$
Therefore, provided that $\eta>0$ is sufficiently small, (4.8) is equivalent to the condition that
${\mathcal{A}}\big(x_{j}^{s}\big)<{\mathcal{A}}(x).$ (4.9)
To summarize, to complete the proof we need to show that $x_{j}^{k_{j}-s}\not\in{\mathcal{P}}$, i.e., $x_{j}^{k_{j}-s}$ is $\mathbb{Q}$}-invisible, whenever (4.9) is satisfied.
Note that then $s\leq s_{j}$ and hence $k_{j}$ is divisible by $s$ due to the divisibility requirement, (4.3). Furthermore, $x_{j}^{s}$ is $\mathbb{Q}$}-invisible and $\chi^{\scriptscriptstyle{eq}}\big(x_{j}^{s}\big)=0$. Set $k^{\prime}_{j}=k_{j}/s$. We have
$x_{j}^{k_{j}-s}=\big(x_{j}^{s}\big)^{k^{\prime}_{j}-1}.$
This guarantees, again by (4.3), that $x_{j}^{k_{j}-s}$ is an admissible odd iterate of $x_{j}^{s}$. Finally, by Lemma 2.8,
$\chi^{\scriptscriptstyle{eq}}\big(x_{j}^{k_{j}-s}\big)=\chi^{\scriptscriptstyle{eq}}\big(x_{j}^{s}\big)=0,$
and $x_{j}^{k_{j}-s}$ is $\mathbb{Q}$}-invisible due to Corollary D.
This completes the proof of (4.4) and Theorem 4.1. ∎
4.1.2. Proof of Theorem B
The idea of the proof is that symmetry and Corollary D rule out the existence of asymmetric orbits of the flow. Furthermore, prime symmetric orbits are necessarily non-alternating and the result follows from Part (M2).
Here is a detailed argument. Recall that when $M$ is centrally symmetric, there are two types of closed Reeb orbits in $M$: symmetric and asymmetric. The former are the closed orbits $x$ such that x(t)=x(t+T/2)-x(t)=x(t+T/2) where $T={\mathcal{A}}(x)$ is the period of $x$, and the latter are the orbits $x$ such that x-x and $x$ are geometrically distinct. Clearly, asymmetric orbits come in pairs, (x,x)(x,-x), with both $x$ and x-x having the same action and the same index. Thus, by Corollary D, the flow has no asymmetric orbits.
To finish the proof it suffices to show that all symmetric closed orbits are non-alternating. This is a standard observation going back to [LLZ]; below we closely follow [GiM].
Consider the homogeneous of degree two Hamiltonian $H$ on ${\mathbb{R}}^{2n}$ with $\{H=1\}=M$. The flow of $H$ on $M$ agrees with the Reeb flow on $M$. Let $\Psi^{t}$ be the linearized flow of $H$ in ${\mathbb{R}}^{2n}$ along a closed symmetric orbit $x$ in $M$ with period $T$. Using the constant trivialization $T{\mathbb{R}}^{2n}\cong{\mathbb{R}}^{2n}\times{\mathbb{R}}^{2n}$ we can view $\Psi^{t}$ as a path in $\operatorname{Sp}(2n)$. Then $\Psi^{T}=(\Psi^{T/2})^{2}$, since the trivialization and $H$ are both centrally symmetric, and therefore, $\Psi^{T}$ is non-alternating. To be more precise, $\Psi^{T}$ has an even number of eigenvalues in the range (1,0)(-1,0) and 1-1 is not an eigenvalue since the Reeb flow is non-degenerate. Fix a point $z$ on $x$ and consider the decomposition $T_{z}{\mathbb{R}}^{2n}=\xi_{z}\oplus{\mathbb{R}}^{2}$, where $\xi_{z}\subset T_{z}M$ is the contact plane and the second factor is the span of the Reeb vector field vector at $z$ and the radial direction. The decomposition is invariant under $\Psi^{T}\colon T_{z}{\mathbb{R}}^{2n}\to T_{z}{\mathbb{R}}^{2n}$, which then decomposes as $\Phi\oplus{\mathit{i}d}$, where $\Phi\colon\xi_{z}\to\xi_{z}$ is the linearized Reeb flow in time $T$. It follows that $\Phi$ and $\Psi^{T}$ have the same eigenvalues other than 1. Therefore, $\Phi$ and hence, by definition, $x$ are also non-alternating.
This completes the proof of the theorem. ∎
Remark 4.6 (Role of symmetry and non-degeneracy).
It is illuminating to examine the role of symmetry, non-degeneracy and our other main results in the proof of Theorem B. The non-degeneracy condition is used to have non-alternating closed orbits defined and then, in some sense even more crucially, to guarantee that all iterates except the bad ones are $\mathbb{Q}$}-visible; cf. Examples 2.12 and 2.13. The lower bound $n$ on the number of non-alternating prime orbits follows in essence from [GK, LZ]; cf. Corollary 3.7. The genuinely new part of Theorem A is not needed here. However, we would have the same lower bound on the total number of prime closed orbits even without non-degeneracy by (M1). Likewise, if we replaced dynamical convexity by convexity or strong dynamical convexity in Theorem B, which are admittedly more restrictive conditions, by [LLZ] or [GiM]; cf. Corollary 3.8.
The symmetry condition together with the fact that distinct orbits have distinct actions (Corollary D) is central to the proof of Theorem B to rule out the existence of asymmetric closed orbits. This step is independent of the non-degeneracy requirement.
4.2. Proof of Theorem E
The proof relies on Theorem C and Corollary D, which are the central parts of the argument. Recall that $\mathcal{P}$} is the set of $\mathbb{Q}$}-visible closed orbits and $\mathcal{P}$$^{\prime}$ stands for $\mathcal{P}$} with the extra element $[W]$ of degree $n$ and action 0. We start with the proof of Part (O1), noting first that the existence of $x\in{\mathcal{P}}$ with $\deg(x)=m$ for any $m\in n-1+2{\mathbb{N}}$ is an immediate consequence of (2.7) and (2.22).
For every $x\in{\mathcal{P}}$, there exists exactly one bar $I_{-}$ ending at $x$ and one bar $I_{+}$ beginning at $x$ by Lemma 2.15 and Theorem C. Furthermore, again due to Lemma 2.15, there are two possibilities:
[…]
or
[…]
Next, denote by $x_{1},x_{2},\ldots$ the elements of $\mathcal{P}$} arranged in the order of (strictly) increasing action. We also set $x_{0}$ to be the extra element $[W]$ of $\mathcal{P}$$^{\prime}$. By definition, this element has action 0 and degree $n$. Let $I_{i}$ be the bar beginning at $x_{i}$.
It immediately follows from (4.10) and (4.11) that the sequence $\deg(I_{i})$ has constant parity. Since $\deg(I_{0})=n$, all $\deg(I_{i})$ have parity $n$.
Again by (4.10) and (4.11), the sequences $\deg(x_{i})$ and $\deg(I_{i})$ are (not necessarily strictly) increasing. Moreover, (4.11) cannot happen when $\deg(x_{i})$ has parity $n+1$; for then $\deg(I_{i})$ would have parity $n+1$. Hence, $\deg(x_{i})$ is strictly increasing at the points of $n-1+2{\mathbb{N}}$. It follows that for every $m$ in this set there is exactly one $x\in{\mathcal{P}}$ with $\deg(x)=m$.
Furthermore, if we had $y\in{\mathcal{P}}$ with $q:=\deg(y)\in n+2{\mathbb{N}}$ we would have at least two orbits in $\mathcal{P}$} of degree q1q-1 or $q+1$ due to (2.22) or, more specifically, (2.23) and Theorem C. This is impossible since, as we have shown, for every $m\in n-1+2{\mathbb{N}}$ an orbit $x\in{\mathcal{P}}$ with $\deg(x)=m$ is unique. This completes the proof of the statement that the degree map $\deg$ on $\mathcal{P}$} takes values in $n-1+2{\mathbb{N}}$ and is a bijection.
Together with the fact that all $\deg(I_{i})$ have parity $n$, this also completely rules out (4.11). Therefore, the degree map $\deg$ on $\mathcal{B}$} takes values in $n-2+2{\mathbb{N}}$ and is also a bijection.
The statement that the ordering of the orbits in $\mathcal{P}$} (or the bars in $\mathcal{B}$}) by the action agrees with the ordering by the index is simply a reformulation of the fact that the sequence $\deg(x_{i})$ is strictly increasing. (Recall that the sequence $x_{i}$ is originally ordered by the action.) Alternatively, this statement is also a consequence of [GG20, Thm. 1.3]. This concludes the proof of Part (O1).
Turning to Part (O2), assume that $x\in{\mathcal{P}}$. Then $\chi^{\scriptscriptstyle{eq}}(x)=(-1)^{\deg(x)}\neq 0$ by (2.13). Then by Lemma 2.8 and, in particular, (2.14), $x^{k}\in{\mathcal{P}}$ for all odd admissible $k\in{\mathbb{N}}$. Thus, using the terminology from [GG20], we see that infinitely many iterates $x^{k}$ occur as carriers of equivariant action selectors (aka equivariant capacities). Therefore, all $\mathbb{Q}$}-visible closed orbits have the same ratio ${\mathcal{A}}(x)/\operatorname{\hat{\mu}}(x)$ by [GG20, Thm. 6.4]. ∎
Remark 4.7.
When the flow is non-degenerate, Part (O1) has an analogue for any field $\mathbb{F}$} with $\deg(x)$ replaced by $\mu(x)$, although the statement is then less precise. Namely, the sequences $\mu(x_{i})$ and $\deg(I_{i})$ are still increasing although now not necessarily strictly. However, for every $m\in n-1+2{\mathbb{N}}$ we still have exactly one $\mathbb{F}$}-visible closed orbit $x$ with $\mu(x)=m$ and all $\deg(I_{i})$ have parity $n$. The key difference is that in this case we cannot rule out the existence of closed $\mathbb{F}$}-visible orbits $x$ with $\mu(x)$ of parity $n$ and intervals in the sequences $\deg(x_{i})$ and $\deg(I_{i})$ where these sequences are constant. The proof is a subset of the proof of Part (O1) combined with the fact that $\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})=\{\mu(x),\mu(x)+1\}$ for any field $\mathbb{F}$} when $x$ is non-degenerate, i.e., the second part of Corollary D holds automatically. We do not know to what degree Part (O2) extends to this case.
4.3. Proofs of Theorem C and Corollary D
The strategy of the proof of the theorem is as follows. Due to the universal coefficient theorem, it suffices to prove the theorem for ${\mathbb{F}}={\mathbb{F}}_{p}$. By the Smith inequality (see Theorem 2.5 or [Se15, ShZh]), $D_{t}:=\dim\operatorname{SH}^{t}(W;{\mathbb{F}}_{p})\geq 1$ for all $t>0$. Moreover, for every $t\in{\mathbb{R}}\setminus{\mathcal{S}}(\alpha)$, there are arbitrarily long intervals $I\subset{\mathbb{R}}$ such that $D_{\tau}\geq D_{t}$ for all $\tau\in I$. Next, we infer from the index recurrence theorem (Theorem 3.3) and an upper bound on the boundary depth (see Theorem 2.3 or [GS]) that there exists $C\in I$ such that $\operatorname{SH}^{C}(W;{\mathbb{F}}_{p})$ is supported only in degrees of parity $n$. This is the central point of the proof, and this is where dynamical convexity and index recurrence in the form of Theorem 3.3 become crucial. Then
$D_{C}=(-1)^{n}\chi(W)=1$
by (2.3). Hence, $D_{t}=1$ for all $t>0$.
4.3.1. Generalities
We start by stating several facts used in the proof of the theorem. In this section, we will also reduce the theorem to an auxiliary result (Proposition 4.8) which is then proved in Section 4.3.2.
Recall that by Theorem 2.5 proved in [Se15, ShZh], for every Liouville domain $W$, a characteristic $p$ field ${\mathbb{F}}_{p}$ and an interval $I$,
$\dim\operatorname{SH}^{pI}(W;{\mathbb{F}}_{p})\geq\dim\operatorname{SH}^{I}(W;{\mathbb{F}}_{p}).$ (4.12)
(The reader may simply set here ${\mathbb{F}}_{p}$ to be the standard field ${\mathbb{Z}}/p{\mathbb{Z}}$ of characteristic $p$.) In particular, $\dim\operatorname{SH}^{pt}(W;{\mathbb{F}}_{p})\geq\dim\operatorname{SH}^{t}(W;{\mathbb{F}}_{p})$ for all $t>0$. Therefore, for all $t>0$ and any field $\mathbb{F}$}, we have
$\dim\operatorname{SH}^{t}(W;{\mathbb{F}})\geq\dim\operatorname{H}(W;{\mathbb{F}}).$ (4.13)
As a consequence, $\dim\operatorname{SH}^{t}(W;{\mathbb{F}})\geq 1$. Indeed, to prove (4.13), we first note that for all $k\in{\mathbb{N}}$ and $t>0$,
$\dim\operatorname{SH}^{t}(W;{\mathbb{F}}_{p})\geq\dim\operatorname{SH}^{t/p^{k}}(W;{\mathbb{F}}_{p})$
due to (2.6). On the other hand, when $k$ is large, by (2.2) we have
$\operatorname{SH}^{t/p^{k}}(W;{\mathbb{F}}_{p})=\operatorname{H}(W;{\mathbb{F}}_{p})[-n].$
Thus, (4.13) holds for every positive characteristic field. Then the case of zero characteristic follows from the universal coefficient theorem and, in particular, (2.4).
With the Smith inequality, (4.13), in mind, the proof of the theorem is based on the following key result giving the opposite bound for a certain sequence of action values.
Proposition 4.8.
In the setting of Theorem C, for every field $\mathbb{F}$} there exists a positive real sequence $C_{l}\to\infty$ with bounded gap such that
$\dim\operatorname{SH}^{C_{l}}(W;{\mathbb{F}})=1$
for all $C_{l}$.
This proposition is proved in Section 4.3.2. Here we only point out that this is the only part of the proof of Theorem C where the finiteness and dynamical convexity conditions are used. The rest of the argument is independent of these requirements. Note also that the sequence $C_{l}$ can be taken independent of the field $\mathbb{F}$}. However, we do not need this fact.
Proof of Theorem C.
As above, by the universal coefficient theorem and, in particular (2.4), it is enough to prove (1.1) for all positive characteristic fields ${\mathbb{F}}_{p}$. Setting
$D_{t}:=\dim\operatorname{SH}^{t}(W;{\mathbb{F}}_{p}),$
our goal is to show that $D_{t}=1$ for all $t>0$. By (4.13), we have $D_{t}\geq 1$, and hence we only need to prove the opposite inequality. Clearly, it suffices to show this for $t>0$ outside the action (period) spectrum ${\mathcal{S}}(\alpha)$ of the contact form $\alpha$.
Thus, fix $t\in{\mathbb{R}}_{+}\setminus{\mathcal{S}}(\alpha)$ and let $t^{\prime}>t$ be such that $t,t^{\prime}]\cap{\mathcal{S}}(\alpha)=\emptyset$. Since $t,t^{\prime}]\cap{\mathcal{S}}(\alpha)=\emptyset$, we have $D_{t}=D_{\tau}$ for all $\tau\in[t,t^{\prime}]$. Then, by the Smith inequality, (4.12), again, applied $k$ times to $I=[0,\tau]$, we have
$D_{\tau}\leq D_{p^{k}\tau}$
for all $k\in{\mathbb{N}}$. Here $p^{k}\tau$ ranges through the interval $I_{k}=[p^{k}t,p^{k}t^{\prime}]$ as $\tau$ ranges through $t,t^{\prime}]$. Now let $C_{l}\to\infty$ be as in Proposition 4.8. Since the sequence $C_{l}$ has bounded gap, for all sufficiently large $k$, the intersection of the sequence $\{C_{l}\}$ and $I_{k}$ is non-empty: $\{C_{l}\}\cap[p^{k}t,p^{k}t^{\prime}]\neq\emptyset$. Therefore, there exists $\tau\in[t,t^{\prime}]$ and $k\in{\mathbb{N}}$ and $l\in{\mathbb{N}}$ such that $p^{k}\tau=C_{l}$. By Proposition 4.8, $D_{C_{l}}\leq 1$. Combining these facts, we arrive at
$D_{t}=D_{\tau}\leq D_{p^{k}\tau=C_{l}}=1.$
Hence, $D_{t}\leq 1$, which concludes the proof of the theorem. ∎
4.3.2. Proof of Proposition 4.8
Fix a field $\mathbb{F}$}, which we suppress in the notation. As in Section 3.2, let us break down the collection ${\mathcal{P}}={\mathcal{P}}({\mathbb{F}})$ of $\mathbb{F}$}-visible closed orbits into several groups, called clusters. We say that two such orbits $x$ and $y$ belong to the same cluster if ${\mathcal{A}}(x)/\operatorname{\hat{\mu}}(x)={\mathcal{A}}(y)/\operatorname{\hat{\mu}}(y)$. (Thus, for instance, (O2) asserts that there is only one cluster.)
Consider now the maps $\operatorname{Beg}\colon{\mathcal{B}}\to{\mathcal{P}}^{\prime}$ and $\operatorname{End}\colon{\mathcal{B}}\to{\mathcal{P}}$ from Theorem 2.16, where $\mathcal{B}$} is the collection of bars of the graded persistence module $\operatorname{SH}(W)$. Let $I\in{\mathcal{B}}$, and $x=\operatorname{Beg}(I)$ and $y=\operatorname{End}(I)$. Then
  • (B1)
    $a={\mathcal{A}}(x)$ and $b={\mathcal{A}}(y)$;
  • (B2)
    $\mu_{-}(y)-\mu_{+}(x)\leq 2$;
  • (B3)
    $\deg(I)=\mu_{+}(x)+1=\mu_{-}(y)-1$ if $\mu_{-}(y)-\mu_{+}(x)=2$.
Here (B1) is the first item in Theorem 2.16, (B2) is (2.25), and (B3) is (2.26).
Our first observation is that beyond some action threshold there are no inter-cluster bars.
Lemma 4.9.
There exists a constant $K_{\scriptscriptstyle{bar}}>0$ such that for every bar $(a,b]$ with $a>K_{\scriptscriptstyle{bar}}$ beginning at an orbit $x$ and ending at $y$, the orbits $x$ and $y$ are in the same cluster.
Remark 4.10.
The constant $K_{\scriptscriptstyle{bar}}$ can be taken independent of the background field $\mathbb{F}$}, but we do not need this fact.
Proof.
Recall that by Theorem 2.3, for any field $\mathbb{F}$} all bars in $\operatorname{SH}(W)$ are finite and bounded from above by a constant $C_{\scriptscriptstyle{bar}}$. (This constant can also be taken independent of $\mathbb{F}$}, but this is not essential for our purposes.) In other words, there exists $C_{\scriptscriptstyle{bar}}>0$ such that any bar $(a,b]$ has length $b-a<C_{\scriptscriptstyle{bar}}$.
Observe that by (B2) the difference $\big|\operatorname{\hat{\mu}}(y)-\operatorname{\hat{\mu}}(x)\big|$ is a priori bounded from above by $2n$. Here we use the fact that $|\mu_{\pm}-\operatorname{\hat{\mu}}|\leq n-1$ by (3.2). Set $\rho={\mathcal{A}}(x)/\operatorname{\hat{\mu}}(x)$ and $\rho^{\prime}={\mathcal{A}}(y)/\operatorname{\hat{\mu}}(y)$ and recall that $b={\mathcal{A}}(y)$ and $a={\mathcal{A}}(x)$. Then a direct calculation shows that
$b-a\geq|\rho^{\prime}/\rho-1|\cdot b-2n\rho.$
We also have $b-a<C_{\scriptscriptstyle{bar}}$. Therefore, since there are only finitely many clusters, we must have $\rho=\rho^{\prime}$ when $a>K_{\scriptscriptstyle{bar}}$ for some constant $K_{\scriptscriptstyle{bar}}$. ∎
Next, let us label the prime orbits as $x_{ij}$, where $i=1,\ldots,i_{0}$ and $j=1,\ldots,j_{0}(i)$, with the index $i$ indicating the cluster. In other words, $x_{ij}^{k}$ and $x_{i^{\prime}j^{\prime}}^{k^{\prime}}$ are in the same cluster if and only if $i=i^{\prime}$. (Note that $x_{ij}^{k}$ may fail to be $\mathbb{F}$}-visible, i.e., $x_{ij}^{k}$ is not necessarily in $\mathcal{P}$}, for some $k\in{\mathbb{N}}$.)
By Theorem 3.3 and Corollary 3.5, there exists a positive real sequence $C_{l}\to\infty$ and for each $i=1,\ldots,i_{0}$ there is a sequence of positive even integers $d_{il}\to\infty$ as $l\to\infty$ such that for every iterated orbit $x_{ij}^{k}$ we have
$\mu_{+}\big(x_{ij}^{k}\big)\leq d_{il}+n-1\textrm{ when }{\mathcal{A}}\big(x_{ij}^{k}\big)<C_{l}$ (4.14)
and
$\mu_{-}\big(x_{ij}^{k}\big)\geq d_{il}+n+1\textrm{ when }{\mathcal{A}}\big(x_{ij}^{k}\big)>C_{l}.$ (4.15)
Indeed, in the setting of Theorem 3.3 and Corollary 3.5 for a recurrence event $k_{ijl}$ we have $k\leq k_{ijl}$ in (4.14) and $k>k_{ijl}$ in (4.15). Here both action inequalities are strict, since the sequence $C_{l}$ can be chosen outside the action spectrum ${\mathcal{S}}(\alpha)$.
The next lemma asserts that for sufficiently large $C_{l}$, the graded vector space $\operatorname{SH}^{C_{l}}(W)$ is supported in degrees $\{d_{il}+n\mid i=1,\ldots,i_{0}\}$.
Lemma 4.11.
Let $C_{l}\to\infty$ and $d_{il}\to\infty$ be as above, i.e., these sequences satisfy (4.14) and (4.15) and all $d_{il}$ are even. Then there exists a constant $K>0$ such that once $C_{l}>K$ we have
$\dim\operatorname{SH}^{C_{l}}(W)=\sum_{i}\dim\operatorname{SH}_{d_{il}+n}^{C_{l}}(W).$ (4.16)
In particular, $\operatorname{SH}^{C_{l}}(W)$ is supported in degrees of the same parity as $n$.
Proof.
Note that $\dim\operatorname{SH}^{C_{l}}(W)$ is equal to the number of bars containing $C_{l}$. Since $b-a<C_{\scriptscriptstyle{bar}}$ by Theorem 2.3 for any bar $(a,b]$, the bars $(a,b]$ with $a<C_{l}-C_{\scriptscriptstyle{bar}}$ do not contribute to $\dim\operatorname{SH}^{C_{l}}(W)$. Next, fix a bar $(a,b]$ containing $C_{l}$, beginning at $x_{ij}^{k}$ and ending at $x_{i^{\prime}j^{\prime}}^{k^{\prime}}$. Let us also assume from now on that $C_{l}>K:=K_{\scriptscriptstyle{bar}}+C_{\scriptscriptstyle{bar}}$. Then $a>K_{\scriptscriptstyle{bar}}$, and hence, by Lemma 4.9, $i=i^{\prime}$, i.e., these orbits are in the same cluster.
We claim that when $C_{l}>K$ the degree of $I=(a,b]$ is in the set $\{d_{il}+n\}$. This will prove (4.16) since the equality simply asserts that $\dim\operatorname{SH}_{m}^{C_{l}}(W)=0$ when $m\neq d_{il}+n$.
To see that $\operatorname{SH}^{C_{l}}(W)$ is supported in the degrees $d_{il}+n$, first note that
$a={\mathcal{A}}\big(x_{ij}^{k}\big)<C_{l}<{\mathcal{A}}\big(x_{ij^{\prime}}^{k^{\prime}}\big)=b.$
Hence, by (4.14) and (4.15),
$\mu_{+}\big(x_{ij}^{k}\big)\leq d_{il}+n-1\quad\textrm{and}\quad\mu_{-}\big(x_{ij^{\prime}}^{k^{\prime}}\big)\geq d_{il}+n+1.$
(Here we are using the fact that $i=i^{\prime}$, i.e., the orbits are in the same cluster.) Due to (B2), we have equality in both of these inequalities and the index difference is exactly 22. Therefore, by (B3), $\deg(I)=\mu_{+}\big(x_{ij}^{k}\big)+1=d_{il}+n$. ∎
It remains to show that $D:=\dim\operatorname{SH}^{C_{l}}(W)=1$. By (2.3),
$\sum_{m}(-1)^{m}\dim\operatorname{SH}^{C_{l}}_{m}(W)=(-1)^{n}\chi(W)=(-1)^{n}.$
At the same time, the left-hand side is equal to $(-1)^{n}D$ by Lemma 4.11. Indeed, each term on the left is either 0 when $m\neq d_{il}+n$ or $(-1)^{n}\dim\operatorname{SH}^{C_{l}}_{m}(W)$ when $m=d_{il}+n$ since $d_{il}$ is even. Thus
$\sum_{m}(-1)^{m}\dim\operatorname{SH}^{C_{l}}_{m}(W)=(-1)^{n}\dim\operatorname{SH}^{C_{l}}(W)=(-1)^{n}D.$
As a consequence, $(-1)^{n}D=(-1)^{n}$, i.e., $D=1$. This completes the proof of the proposition and hence of the theorem. ∎
4.3.3. Proof of Corollary D
The first assertion is a consequence of (2.12), (2.24) and Theorem C. Indeed, if we had two $\mathbb{F}$}-visible orbits $x$ and $x^{\prime}$ with action $a$, by (2.12) and (2.24), we would have at least four bars with end-point $a$ and hence $\dim\operatorname{SH}^{t}(W)\geq 2$ for some $t$ close to $a$. Likewise, $\dim\operatorname{SH}(x;{\mathbb{F}})=2$ for all $\mathbb{F}$}-visible orbits by (2.24) and Theorem C. Now the second assertion follows immediately from (2.10) and the first one. ∎
4.4. Proof of Corollary F
As was already mentioned in Section 1.3, by Part (O2) of Theorem E, we can scale $M$ so that $\operatorname{\hat{\mu}}(x)={\mathcal{A}}(x)$ for all closed orbits in $M$. From now on we assume that this condition is satisfied. Denote by $x_{1},\ldots,x_{n}$ the prime closed orbits in $M$. Then the action spectrum of $M$ is the same as the mean index spectrum and is simply the union of the sequences $\Delta_{j}{\mathbb{N}}$. Note also that we have $\Delta_{j}/\Delta_{i}\not\in{\mathbb{Q}}$, when $i\neq j$, again by (O2), since all closed orbits have distinct actions due to Corollary D.
The prime closed orbits in $M_{\vec{\Delta}}$ have the form
$y_{j}(t)=(0,\ldots,0,z_{j}(t),0,\ldots,0),$
where $z_{j}(t)=\sqrt{\Delta_{j}}\exp\big(2\sqrt{-1}t/\Delta_{j}\big)$ with $t\in[0,\pi\Delta_{j})$. Therefore, in suitable coordinates, the Poincaré return map $\Phi_{j}$ along $y_{j}$ is the $(n-1)\times(n-1)$ diagonal matrix with entries $\exp(2\pi\sqrt{-1}\lambda_{ij})$, where $\lambda_{ij}=\Delta_{j}/\Delta_{i}$ and $i\neq j$. Since $\Delta_{j}/\Delta_{i}\not\in{\mathbb{Q}}$, all orbits $y_{j}$ are strongly non-degenerate and all eigenvalues of $\Phi_{j}$ are elliptic. Note, however, that the “logarithmic” eigenvalues $\lambda_{ij}$ are not normalized in any sense. For instance, the eigenvalues $\lambda_{ij}$ can be arbitrarily large or small and need not be of first Krein type.
Recall that the non-resonance condition is that for every $j$ the ratios $\pm\Delta_{j}/\Delta_{i}=\pm\lambda_{ij}$ are distinct modulo $\mathbb{Z}$} for every fixed $j$ and all $i\neq j$. This guarantees that the first Krein type eigenvalues are distinct for every $j$ and are not complex conjugate to each other.
A straightforward calculation shows that
$\operatorname{\hat{\mu}}(y_{j})=2\Delta_{j}\sum_{i=1}^{n}\frac{1}{\Delta_{i}}.$
As a consequence of Theorem A, the prime orbits $x_{1},\ldots,x_{n}$ are non-alternating. By [Vi89], the sum in this formula is $1/2$; see also [GiK]. Therefore,
$\operatorname{\hat{\mu}}(y_{j})=\Delta_{j}=\operatorname{\hat{\mu}}(x_{j}).$ (4.17)
Clearly, ${\mathcal{A}}(x_{j}^{k})=k\Delta_{j}={\mathcal{A}}(y_{j}^{k})/\pi$. By Theorem E, for each $m\in{\mathbb{N}}$, the $m$th entry in the action spectrum, for $M$ and $M_{\vec{\Delta}}$, is the value of the $m$th $S^{1}$-equivariant spectral invariant over $\mathbb{Q}$}, which is attained on an orbit of index $n+m$; see [GG20]. It follows that $\mu\big(x_{j}^{k}\big)=\mu\big(y_{j}^{k}\big)$. By [GG22, Cor. 4.3], for every $j$ the linearized return maps along $y_{j}$ and $x_{j}$ have the same first Krein type eigenvalues, which are all distinct. Furthermore, the linearized flows also have the same mean index by (4.17). As a consequence, the linearized flows along $x_{j}$ and $y_{j}$ agree as elements of $\widetilde{\operatorname{Sp}}(2(n-1))$ up to conjugation. ∎
Remark 4.12.
Without the non-resonance condition, the argument still provides some comparison information about the linearized flows along $x_{j}$ and $y_{j}$. Namely, it shows that the return maps have the same first Krein type eigenvalues, counted with multiplicity which is defined as follows. Assume that a linear symplectic map has both $w$ and $\bar{w}$ as first Krein type eigenvalues and their algebraic multiplicities are $m$ and, respectively, $l$. Then we set, by definition, the multiplicity of $w$ to be mlm-l and the multiplicity of $\bar{w}$ to be lml-m; see [GG22, Sec. 4]. (Note that with this way of counting multiplicity an elliptic orbit such that its first Krein type eigenvalues come in complex conjugate pairs is indistinguishable from a hyperbolic orbit.) Furthermore, as in the non-resonance case, $x_{j}$ and $y_{j}$ have the same mean index (see (4.17)) and also, up to a factor, the same action.
5. Proof of the index recurrence theorem
Our goal in this section is to prove the index recurrence theorem, Theorem 3.3. The section is broken down into two parts: we first review some standard facts from geometry of numbers and then turn to the actual proof.
5.1. Input from geometry of numbers
The existence of the sequences $C_{l}$, $d_{il}$ and $k_{ijl}$ with the required properties hinges on two standard facts along the lines of Kronecker’s and Minkowski’s theorems; see, e.g., [Ca]. These facts are completely standard and obvious to the experts.
Let $U\subset{\mathbb{T}}^{n}={\mathbb{R}}^{n}/{\mathbb{Z}}^{n}$ be a neighborhood of 0 and $\theta\in{\mathbb{T}}^{n}$. Consider the sequence $k_{1}\leq k_{2}\leq k_{3}\leq\ldots$ of all positive integers such that $k_{l}\theta\in U$. Clearly, $k_{l}\to\infty$. Moreover, it is well known that the orbit $\{k\theta\mid k\in{\mathbb{N}}\}$ is equidistributed in the closure $G$ of $\theta{\mathbb{Z}}$ in ${\mathbb{T}}^{n}$. In particular, the density of this sequence in $\mathbb{N}$} is equal to the $\dim G$-dimensional volume of $U\cap G$.
Lemma 5.1.
The sequence $k_{l}$ has bounded gap.
For the sake of completeness we include a short proof of the lemma.
Proof.
Since $G$ is compact, there exists a finite collection of integers $S=\{s_{1},\ldots,s_{N}\}$ such that $G$ is covered by the shifted neighborhoods $s_{i}\theta+U$. Thus, for every $\ell\in{\mathbb{N}}$, we have $\ell\theta\in s_{i}\theta+U$ for some $i$, which might not be unique. We pick one such $i$ and write $s_{i}=:s(\ell)$.
Set $q_{\ell}=\ell-s(\ell)$. Clearly, $q_{\ell}\theta\in U$. Thus the sequence $q_{\ell}$ takes values in the set $M:=\{k_{l}\mid l\in{\mathbb{N}}\}$ and $q_{\ell}\to\infty$ since the set $S$ is finite. The maximal gap in the sequence $k_{l}$ is equal to the maximal gap in the set $M$ which is bounded from above by the maximal gap in the sequence $q_{\ell}$. The latter sequence has bounded gap, for
$|q_{\ell+1}-q_{\ell}|\leq 1+|s(\ell+1)-s(\ell)|\leq 1+\max_{i,j\in S}|s_{i}-s_{j}|.$
We will also need the following consequence of Lemma 5.1 along the lines of Minkowski’s theorem on linear forms; see, e.g., [Ca]. Consider a collection of $p<n$ linear maps
$f_{s}\colon{\mathbb{R}}^{n}\to{\mathbb{R}},\quad s=1,\ldots,p$
and positive constants $\delta_{1}>0,\ldots,\delta_{p}>0$. We are interested in integer solutions $K\in{\mathbb{Z}}^{n}$ of the system of inequalities
$|f_{s}(K)|<\delta_{s},\quad s=1,\ldots,p.$ (5.1)
Lemma 5.2.
The system of inequalities (5.1) has a sequence of non-zero solutions $K_{l}\to\infty$ in ${\mathbb{Z}}^{n}$ with bounded gap, i.e., such that the norm of $K_{l}$ goes to $\infty$, but the norm of the difference $K_{l+1}-K_{l}$ is bounded by a constant independent of $l$. Moreover, we can make all components of $K_{l}$ divisible by any fixed integer.
Since $p<n$, this system has infinitely many distinct solutions by the Minkowski theorem; see [Ca]. However, the standard proof of this theorem does not immediately imply the bounded gap part. Hence we include a proof of the lemma for the sake of completeness.
Proof of Lemma 5.2.
Let $\tilde{U}\subset{\mathbb{R}}^{n}$ be a small ball centered at the origin such that $|f_{s}|_{\tilde{U}}|<\delta_{s}$ for all $s$ and the balls $K+\tilde{U}$ for all $K\in{\mathbb{Z}}^{n}$ do not overlap. Denote by $U$ the projection of $\tilde{U}$ to ${\mathbb{T}}^{n}={\mathbb{R}}^{n}/{\mathbb{Z}}^{n}$.
Since $p<n$, the hyperplanes $\ker f_{s}$ have a non-zero intersection. Pick a non-zero vector
$\tilde{\theta}\in\bigcap_{s=1}^{p}\ker f_{s}.$
Without loss of generality, we may assume that $\tilde{\theta}\not\in{\mathbb{Z}}^{n}$ for otherwise we can simply set $K_{l}=l\tilde{\theta}$. Thus the projection $\theta$ of $\tilde{\theta}$ to ${\mathbb{T}}^{n}$ is also non-zero. By Lemma 5.1, there exists a bounded gap sequence $k_{l}\to\infty$ such that $k_{l}\theta\in U$ or equivalently
$k_{l}\tilde{\theta}\in K_{l}+\tilde{U}$
for some sequence $K_{l}\in{\mathbb{Z}}^{n}$. Clearly, $K_{l}$ has bounded gap and $K_{l}\to\infty$.
We claim that $K_{l}$ satisfies the system of inequalities, (5.1). Indeed, let us write
$k_{l}\tilde{\theta}=K_{l}+v_{l}$
where $v_{l}\in\tilde{U}$. Then $f_{s}(k_{l}\tilde{\theta})=0$, and hence
$|f_{s}(K_{l})|=|f_{s}(v_{l})|<\delta_{s}.$
Replacing $\tilde{\theta}$ by $N\tilde{\theta}$, we can make all components of $K_{l}$ divisible by $N\in{\mathbb{N}}$ if needed. ∎
5.2. Proof of Theorem 3.3
The strategy of the proof is as follows. After settling some preliminaries in Sections 5.2.1 and 5.2.2, we show in Section 5.2.3 that for any element $\Phi:=\Phi_{ij}$ assertions (IR1)(IR5) are formal consequences of the requirement that $k=k_{ijl}$ satisfies a certain system of inequalities completely determined by $\Phi$. Here $d:=d_{il}$ is determined by $k$ and, of course, $\Phi$ by the first part of (IR1). Furthermore, (IR5) can be taken as the defining condition for $C_{l}$. (The existence of an infinite bounded gap sequence of solutions $k$ of this system of inequalities is a consequence of Lemma 5.1.)
As the second step of the proof, we prove in Section 5.2.4 that for a finite collection of maps $\Phi_{ij}$ as in the theorem the resulting systems of inequalities have bounded gap sequences of solutions $k_{ijl}$ such that $d_{il}$ is independent of $j$ within every cluster and $C_{l}$ is independent of $i$ and $j$. This is done by using Lemma 5.2. This independence is the key new component of the second step. For simply applying the first step to each $\Phi_{ij}$ individually would result in $d_{il}$ depending on $j$ and $C_{l}$ depending in addition on $i$ and $j$.
5.2.1. Isospectral families
In the proof we will use a general fact from symplectic linear algebra, which allows one to simplify certain index calculations. Recall that a linear map is said to be semisimple if for every eigenvalue its algebraic multiplicity is equal to the geometric multiplicity. A family of linear maps $P_{t}$, $t\in[0,1]$, is said to be isospectral if the eigenvalues of $P_{t}$ are constant. (Then, as a consequence, they are constant as a multiset.)
For such a family, the mean index $\operatorname{\hat{\mu}}(P_{t})$ and, when $P$ is non-degenerate, the Conley–Zehnder index $\mu(P_{t})$ are constant as follows immediately from the definitions; see, e.g., [Lo02, SZ]. (The non-degeneracy condition is essential: more sensitive invariants such as $\mu_{\pm}$ or $\beta_{\pm}$ need not be constant.)
Lemma 5.3.
Every $P\in\operatorname{Sp}(2m)$ can be connected in $\operatorname{Sp}(2m)$ to a semisimple linear map by an isospectral family.
As a consequence, in many index calculations we can assume that the map is semisimple by replacing $P_{0}$ by $P_{1}$. Lemma 5.3 has certainly been known in some contexts, but we are not aware of any reference and we include a proof for the sake of completeness. In fact, we need the lemma only when $P$ is elliptic and non-degenerate, but the general case might be of independent interest and the proof under these conditions is not much shorter.
Proof.
By decomposing $P$ as a direct sum depending on the type of eigenvalues, we only need to prove the lemma when $P$ is one of the following three types:
  • All eigenvalues of $P$ are elliptic (i.e., unit) and different from 1-1.
  • The only eigenvalue is 1-1.
  • All eigenvalues are hyperbolic, i.e., not unit.
We will start with the first case. Since $P$ has no real negative eigenvalues, it lies in the image of the exponential map $\exp\colon{\mathfrak{sp}}(2m)\to\operatorname{Sp}(2m)$ by [CLW, Sect. 2.6]. Thus $P=e^{A}$, where $A$ is a linear Hamiltonian operator, aka a linear Hamiltonian vector field, and it is sufficient to connect $A$ to a semisimple operator in ${\mathfrak{sp}}(2m)$. When we order the linear Darboux coordinates as $(p_{1},\ldots,p_{m},q_{1},\ldots,q_{m})$, the operator $A$ takes the form
$A=\left[\begin{array}[]{cc}A_{11}&A_{12}\\ A_{21}&A_{22}\end{array}\right],$
where $A_{11}\in{\mathfrak{gl}}(n)$ is an arbitrary matrix, $A_{22}=-A_{11}^{\top}$, and $A_{12}$ and $A_{21}$ are symmetric: $A_{12}=A_{12}^{\top}$ and $A_{21}=A_{21}^{\top}$. Furthermore, since $P$ is elliptic, $A$ has pure imaginary eigenvalues. Let us denote the spectrum of $A$ by $\sigma(A)$ with eigenvalues taken without multiplicity. Without loss of generality, we may assume that $A$ does not have semisimple eigenvalues. Then, depending on $\sigma(A)$, the operator $A$ breaks down into a direct sum of the following four normal forms:
  • (NF1)
    In this case $\sigma(A)=\{0\}$, $m$ is odd, and $A_{22}$ is the negative Jordan block of size $m$ with eigenvalue 0, i.e.,
    $A_{22}=\left[\begin{array}[]{cccc}0&-1&&\\ &\ddots&\ddots&\\ &&\ddots&-1\\ &&&0\\ \end{array}\right]$
    with the remaining empty spots standing for zeros, and $A_{12}=0=A_{21}$.
  • (NF2)
    As above, $\sigma(A)=\{0\}$ and $A_{22}$ is the negative Jordan block of size $m$ with eigenvalue 0 and $A_{12}=0$, but $m$ is arbitrary and all entries of $A_{21}$ are zero except the bottom right corner which is $\pm 1$.
  • (NF3)
    Here $\sigma(A)=\{\sqrt{-1}b\}$ with $b>0$, $m$ is even, and
    $A_{22}=\left[\begin{array}[]{cccc}B&-\tilde{I}&&\\ &\ddots&\ddots&\\ &&\ddots&-\tilde{I}\\ &&&B\\ \end{array}\right],$
    where
    $B=\left[\begin{array}[]{cc}0&-b\\ b&0\end{array}\right]\textrm{ and }\tilde{I}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]$
    and the empty spots are again zeros. Furthermore, $A_{12}=0$ and all entries of $A_{21}$ are zero except the $2\times 2$ block $\pm\tilde{I}$ at the bottom right corner.
  • (NF4)
    In this case $\sigma(A)=\{\sqrt{-1}b\}$ with $b\neq 0$, $m$ is odd, and $A_{22}$ is the negative Jordan block as in (NF1). However, $A_{12}$ and $A_{21}$ are anti-diagonal matrices:
    $A_{12}=\left[\begin{array}[]{ccccc}&&&&-b\\ &&&b&\\ &&\reflectbox{$\ddots$}&&\\ &b&&&\\ -b&&&&\\ \end{array}\right]=-A_{21}.$
These normal forms are taken from [CLW, Sect. 2.6] and are simpler to use for our purposes than their original variant from [Wi36].
Next, let us replace $\pm 1$ (but not $\pm b$) by $\pm t$ in all occurrences in these matrices. For instance, in (NF3) $\tilde{I}$ becomes $t\tilde{I}$ in $A_{22}$, etc. In each case, we obtain a family of Hamiltonian operators $A(t)$, $t\in[0,1]$, with $A(1)=A$.
It is easy to see that $A(0)$ is semisimple in all four cases. Indeed, $A(0)=0$ in (NF1) and (NF2). In (NF3) and (NF4), $A(0)$ is equivalent via a linear symplectic change of variables to a direct sum of infinitesimal rotations with eigenvalues $\pm\sqrt{-1}b$ in $m$ symplectic planes.
We claim that the four families $A(t)$ are isospectral. Let $\rho_{t}(z)=\det\big(zI-A(t)\big)$ be the characteristic polynomial. It is sufficient to show that $\rho_{t}(z)$ is independent of $t$ and this can be done by a direct calculation. In (NF1) or (NF2), we have the determinant of a block-diagonal or block-triangular matrix with triangular matrices on the diagonal, and $\rho_{t}(z)=z^{2m}$ for all $t$. Likewise, in (NF3) we have the determinant of a block-triangular matrix with block-triangular matrices on the diagonal, and $\rho_{t}(z)=(z^{2}+b^{2})^{m}$. To calculate $\rho_{t}(z)$ in (NF4), we expand the determinant in the first row and then expand each of the resulting two non-zero terms in the last row. Then $\rho_{t}(z)$ factors as the product of $z^{2}+b^{2}$ and the characteristic polynomial of a $(2m-2)\times(2m-2)$ matrix of exactly the same form, but with $b$ replaced by b-b. (The resulting matrix is not a normal form because m1m-1 is even. However, this does not matter since we are just calculating the characteristic polynomial.) Proceeding inductively, we again see that $\rho_{t}(z)=(z^{2}+b^{2})^{m}$. This concludes the proof for an elliptic map $P$ without eigenvalue 1-1.
The second case is where 1-1 is the only eigenvalue of $P$. Then P-P is covered by the previous case. Thus there exists an isospectral family connecting P-P to a semisimple symplectic map. (In fact, this map is $I$ because $1$ is the only eigenvalue of P-P.) Multiplying this family by I-I, we obtain the required isospectral family for $P$.
The remaining case is where $P$ is hyperbolic. This case can be dealt with in a fashion similar to the first two, but a much simpler proof is available. Namely, ${\mathbb{R}}^{2m}=W_{-}\oplus W_{+}$ be the decomposition of ${\mathbb{R}}^{2m}$ into subspaces corresponding to the eigenvalues $\lambda$ of $P$ with $|\lambda|<1$ and, respectively, $|\lambda|>1$. These are $P$-invariant Lagrangian subspaces. As a consequence, the symplectic form gives rise to an isomorphism $W_{+}=W_{-}^{*}$ and $P=(P_{-},\,P_{+})$ with $P_{+}=(P_{-}^{-1})^{\top}$ in this decomposition. (In other words, identifying ${\mathbb{R}}^{2m}=W_{-}\oplus W_{-}^{*}$ with $T^{*}W_{-}$ we identify $P\colon T^{*}W_{-}\to T^{*}W_{-}$ with the map induced by $P_{-}^{-1}\colon W_{-}\to W_{-}$.) Let us connect $P_{-}$ to an semisimple linear map by a family $P_{-}(t)\colon W_{-}\to W_{-}$ of isospectral linear maps. The existence of such a family immediately follows from the Jordan normal form theorem. Then $P_{t}:=\big(P_{-}(t),\,(P_{-}^{-1}(t))^{\top}\big)$ is a family of isospectral maps in $\operatorname{Sp}(2m)$ connecting $P$ to a semisimple map. This completes the proof of Lemma 5.3. ∎
Remark 5.4.
Throughout the proof of the lemma, one could have used the normal forms for symplectic matrices from [Wi36], see also [Lo02], but these normal forms have a complicated format and the proof would be more involved. The last step of the proof can also be used to directly obtain the normal forms of hyperbolic symplectic or Hamiltonian operators from the Jordan normal form theorem.
5.2.2. Preliminary observations
Before turning to the actual proof of Theorem 3.3 it is useful to make a couple of simple observations.
The claim that the sequences $C_{l}$, $d_{il}$ and $k_{ijl}$ can be taken strictly increasing is a formal consequence of the rest of the theorem. Indeed, observe that every bounded gap sequence $d_{il}\to\infty$ has a bounded gap strictly increasing refinement $s\mapsto d_{il_{s}}$. Moreover, for any $c>0$ we may ensure by further thinning the sequence that $d_{il_{s+1}}-d_{l_{s}}>c$. Hence, by passing to subsequences of $d_{il}$ and $k_{ijl}$, we can guarantee that $d_{il}$ is strictly increasing, $d_{il+1}-d_{il}>c$ and these sequences still have bounded gap. Then, since $\operatorname{\hat{\mu}}\big(\Phi^{k_{ijl}}_{ij}\big)=k_{ijl}\operatorname{\hat{\mu}}(\Phi_{ij})$ and $\operatorname{\hat{\mu}}(\Phi_{ij})>0$, we conclude from (IR1) that the sequences $k_{ijl}$ are automatically strictly increasing when $c$ is sufficiently large. Likewise, the sequence $C_{l}$ is forced to be strictly increasing by (IR5) once $\eta>0$ is sufficiently small.
In a similar vein, again by (IR1), the sequences $d_{il}$ and $k_{ijl}$ have bounded gap if and only if one of them does. Thus it suffices to prove this for the sequences $k_{ijl}$.
To summarize, we only need to show that there exist sequences $C_{l}$, $d_{il}\to\infty$ and $k_{ijl}\to\infty$ satisfying the requirements of the theorem and such that one of the sequences $k_{ijl}$ has bounded gap, and that we can make $d_{il}$ and $k_{ijl}$ divisible by any pre-assigned integer.
Finally, note that throughout the proof we can assume without loss of generality that $\eta>0$ is sufficiently small, e.g., $\eta<\min a_{ij}$.
5.2.3. The case of one map: $r=1$.
Let $\Phi=\Phi_{1j}\in\widetilde{\operatorname{Sp}}(2m)$. Throughout the argument we suppress $i$ and $j$ in the notation, i.e., we will write $k=k_{l}$ for $k_{1jl}$, etc.
Let us decompose $\Phi$ as the product of a loop $\varphi$ and a direct sum $\Phi_{\scriptscriptstyle{deg}}\oplus\Phi_{h}\oplus\Phi_{e}$ with the following properties. The end-point $\Phi_{\scriptscriptstyle{deg}}(1)$ is totally degenerate and $\operatorname{\hat{\mu}}(\Phi_{\scriptscriptstyle{deg}})=0$. The path $\Phi_{h}$ has a hyperbolic end-point $\Phi_{h}(1)$, while $\Phi_{e}$ is the elliptic part of $\Phi$, i.e., all eigenvalues of $\Phi_{e}(1)$ are on the unit circle and different from 1 due to the presence of $\Phi_{\scriptscriptstyle{deg}}$. We write the eigenvalues of $\Phi_{e}(1)$ as $\exp(\pm 2\pi\sqrt{-1}\lambda_{q})$, where $0<|\lambda_{q}|\leq 1/2$ and $q=1,\ldots,m^{\prime}$. Here for each pair of complex conjugate eigenvalues it is convenient to pick the sign of $\lambda_{q}$ so that the resulting eigenvalue is of the first Krein type; see [SZ]. (We will elaborate on this choice later.) Furthermore, we then require that
$\operatorname{\hat{\mu}}(\Phi_{e})=2\sum_{q}\lambda_{q}.$ (5.2)
Note that one or more terms in this decomposition might be absent. For instance, $\Phi_{\scriptscriptstyle{deg}}$ is not present when $\Phi(1)$ is non-degenerate.
It is easy to see that such a decomposition exists; see, e.g., [SZ, Sec. 3]. (In general, the decomposition is not unique in $\widetilde{\operatorname{Sp}}(2m)$ without extra assumptions on the index of $\Phi_{h}$. For instance, the loop $\varphi$ could have been absorbed into $\Phi_{h}$; it is needed only when $\Phi$ has no hyperbolic eigenvalues.) Throughout the rest of the proof, we will assume that $\Phi_{e}$ is present in the decomposition, i.e., $m^{\prime}\geq 1$. It is not hard to see from the proof that the argument goes through with minor modifications when $\Phi_{e}$ is absent.
Set
$\epsilon_{0}=\min\{\|\lambda_{q}\ell\|\mid\lambda_{q}\ell\not\in{\mathbb{Z}},1\leq|\ell|\leq\ell_{0},1\leq q\leq m^{\prime}\}>0,$ (5.3)
where $\|\cdot\|$ stands for the distance to the nearest integer. We will require $\epsilon>0$ (and $\eta>0$) to be so small that
$\epsilon\leq\epsilon_{0}\quad\textrm{and}\quad 2m^{\prime}\epsilon<\eta<1/2.$ (5.4)
Furthermore, denote by $U\subset{\mathbb{T}}^{m^{\prime}}$ the neighborhood of the unit given by
$U=\big\{(e^{2\pi\sqrt{-1}\tau_{1}},\ldots,e^{2\pi\sqrt{-1}\tau_{m^{\prime}}})\,\big|\,|\tau_{1}|<\epsilon,\ldots,|\tau_{m^{\prime}}|<\epsilon\big\},$
and set
$\theta=\big(e^{2\pi\sqrt{-1}\lambda_{1}},\ldots,e^{2\pi\sqrt{-1}\lambda_{m^{\prime}}}\big)\in{\mathbb{T}}^{m^{\prime}}.$
In what follows, to be consistent with Section 5.1, we will use additive notation for the group operations in ${\mathbb{T}}^{m^{\prime}}$.
Lemma 5.5.
Assume that $k$ is such that $k\theta\in U$, or equivalently
$\|k\lambda_{q}\|<\epsilon\textrm{ for $q=1,\ldots,m^{\prime}$},$ (5.5)
and that $k$ is divisible by the least common multiple $D$ of the degrees of all roots of unity among the elliptic eigenvalues of $\Phi$. Let $d$ be the nearest integer to $\operatorname{\hat{\mu}}\big(\Phi^{k}\big)$ or, more explicitly,
$d=\big[k\operatorname{\hat{\mu}}(\Phi)\big]=k\big(\operatorname{\hat{\mu}}(\varphi)+\operatorname{\hat{\mu}}(\Phi_{h})\big)+[k\operatorname{\hat{\mu}}(\Phi_{e})].$ (5.6)
Then conditions (IR1)(IR4) are satisfied with $\Phi:=\Phi_{ij}$, $k=k_{ijl}$, etc.
Before proving the lemma, we note that (5.5) has a bounded gap sequence of solutions $k_{l}\to\infty$ by Lemma 5.1 and then, as we have already pointed out, the corresponding sequence $d=d_{l}\to\infty$ has bounded gap automatically. For any $a:=a_{ij}>0$ to find a bounded gap sequence $C_{l}\to\infty$ meeting condition (IR5), we can just pick any $C_{l}\not\in a{\mathbb{N}}$ in the range $(k_{l}a,k_{l}a+\eta)$ assuming that $\eta<a$. Furthermore, to make sure that $k$ is divisible by $D$ or any fixed integer $N>0$, it suffices to replace $\epsilon$ by $\epsilon/N$ in the definition of $U$. Then, after applying Lemma 5.1 for this new neighborhood $U$, we replace sequence of solutions $k$ by $Nk$ and have a bounded gap sequence of solutions satisfying (5.5) and divisible by $N$. As the next step, to meet the requirement of (5.6) we have $d$ replaced by $Nd$, and hence also divisible by $N$, provided that $\epsilon>0$ is sufficiently small. Thus, to complete this step of the proof, it remains to prove Lemma 5.5. The proof has a substantial overlap with [GG20, Sec. 5.2.1] and is implicitly contained in that section.
Proof of Lemma 5.5.
We start with conditions (IR1)(IR3), which we restate now for the reader’s convenience:
  • (IR1”)
    $\big|\operatorname{\hat{\mu}}\big(\Phi^{k}\big)-d\big|<\eta$ and
    $d-m\leq\mu_{-}\big(\Phi^{k}\big)\leq\mu_{+}\big(\Phi^{k}\big)\leq d+m;$
  • (IR2”)
    $\mu_{\pm}\big(\Phi^{k+\ell}\big)=d+\mu_{\pm}\big(\Phi^{\ell}\big)$ when $0<\ell\leq\ell_{0}$;
  • (IR3”)
    $\mu_{+}\big(\Phi^{k-\ell}\big)=d-\mu_{-}\big(\Phi^{\ell}\big)+\big(\beta_{+}\big(\Phi^{\ell}\big)-\beta_{-}\big(\Phi^{\ell}\big)\big)$, where $0<\ell\leq\ell_{0}<k$ and $\beta_{+}-\beta_{-}=b_{+}-b_{-}\leq m$ with $\beta_{\pm}=0$ when $\Phi^{\ell}$ is non-degenerate.
In spite of a minor change of notations, these are exactly conditions (IR1)(IR3) stated for one map $\Phi=\Phi_{ij}$ with $k=k_{ijl}$, etc.
Note that concatenation with the loop $\varphi$ simply results in the shift of $d$ and the indices in each of these formulas by $k\operatorname{\hat{\mu}}(\varphi)\in{\mathbb{Z}}$. Hence, without loss of generality, we can assume that $\varphi={\mathit{i}d}$. Likewise, the effect of $\Phi_{h}$ is the shift by $k\operatorname{\hat{\mu}}(\Phi_{h})$ and again we can assume that $\Phi_{h}$ is trivial, i.e., $\Phi=\Phi_{\scriptscriptstyle{deg}}\oplus\Phi_{e}$.
To prove the first part of (IR1”), observe that by (5.2) and homogeneity of the mean index we have
$\operatorname{\hat{\mu}}(\Phi^{k})=2k\sum_{q}\lambda_{q},$ (5.7)
since $\operatorname{\hat{\mu}}(\Phi_{\scriptscriptstyle{deg}})=0$. By (5.3) for $\ell=1$, (5.4) and (5.5), we have
$d=\sum_{q}\mathit{sign}(\lambda_{q})\big[2k|\lambda_{q}|\big],$ (5.8)
where $\cdot\,$] stands for the nearest integer. Then
$\big|d-\operatorname{\hat{\mu}}\big(\Phi^{k}\big)\big|<2m^{\prime}\epsilon<\eta,$ (5.9)
again by (5.5). This also shows that $d$ is unambiguously defined.
To prove the second part of (IR1”), we combine (3.2) and (5.9) and use the triangle inequality. Thus we have
$\big|d-\mu_{\pm}\big(\Phi^{k}\big)\big|\leq\big|d-\operatorname{\hat{\mu}}\big(\Phi^{k}\big)\big|+\big|\mu_{\pm}\big(\Phi^{k}\big)-\operatorname{\hat{\mu}}\big(\Phi^{k}\big)\big|\leq\eta+m^{\prime}.$
Then, since the left-hand side is an integer,
$\big|d-\mu_{\pm}(\Phi^{k})\big|\leq m^{\prime}\leq m,$
establishing the second inequality in (IR1”).
Let us now turn to the proof of (IR2”) and (IR3”). The fact that
$\beta_{+}(\Phi)-\beta_{-}(\Phi)=b_{+}(\Phi)-b_{-}(\Phi)\leq m$
in (IR3”) follows directly from the definitions of these invariants; cf. (3.8).
Next, observe that these formulas are additive in $\Phi$ in the following sense. Assume that $\Phi=\Phi^{\prime}\oplus\Phi^{\prime\prime}$, i.e.,
$\Phi(1)=\Phi^{\prime}(1)\oplus\Phi^{\prime\prime}(1)\textrm{ and }\operatorname{\hat{\mu}}(\Phi)=\operatorname{\hat{\mu}}(\Phi^{\prime})+\operatorname{\hat{\mu}}(\Phi^{\prime\prime}).$
Then
$\mu_{\pm}\big(\Phi^{k}\big)=\mu_{\pm}\big((\Phi^{\prime})^{k}\big)+\mu_{\pm}\big((\Phi^{\prime\prime})^{k}\big)$
and
$\beta_{\pm}\big(\Phi^{k}\big)=\beta_{\pm}\big((\Phi^{\prime})^{k}\big)+\beta_{\pm}\big((\Phi^{\prime\prime})^{k}\big)$
for all $k\in{\mathbb{N}}$ by (3.3) and (3.7). Moreover, when $k$ satisfies (5.5) for $\Phi$, it also satisfies this condition for $\Phi^{\prime}$ and $\Phi^{\prime\prime}$. Then setting, as in Lemma 5.5 and (5.6), $d\big((\Phi^{\prime})^{k}\big)$ to be the nearest integer to $\operatorname{\hat{\mu}}\big((\Phi^{\prime})^{k}\big)$ and $d\big((\Phi^{\prime\prime})^{k}\big)$ to be the nearest integer to $\operatorname{\hat{\mu}}\big((\Phi^{\prime\prime})^{k}\big)$ and using self-explanatory notation, we have
$d\big(\Phi^{k}\big)=d\big((\Phi^{\prime})^{k}\big)+d\big((\Phi^{\prime\prime})^{k}\big).$
It follows that (IR2”) and (IR3”) hold for $\Phi$ whenever they hold for $\Phi^{\prime}$ and $\Phi^{\prime\prime}$ assuming that $k$ satisfies (5.5) for $\Phi$.
We will use this observation to reduce the proof to several particular cases. Throughout the argument we will assume without specifically stating that $k\in{\mathbb{N}}$ and $\ell\in{\mathbb{N}}$ and $k>\ell$ in (IR3”).
When $\Phi$ is totally degenerate and $\operatorname{\hat{\mu}}(\Phi)=0$, i.e., $\Phi=\Phi_{\scriptscriptstyle{deg}}$, we have $d=0$ and $\mu_{\pm}\big(\Phi^{k}\big)=\mu_{\pm}(\Phi)=\pm\beta_{\pm}(\Phi)$ for all $k\in{\mathbb{N}}$ by (3.5). Then (IR2”) reduces to the identity
$\beta_{\pm}\big(\Phi^{k+\ell}\big)=\beta_{\pm}(\Phi)=\beta_{\pm}\big(\Phi^{\ell}\big)$
and, in a similar fashion, (IR3”) turns into
$\beta_{+}\big(\Phi^{k-\ell}\big)=-\big(-\beta_{-}\big(\Phi^{\ell}\big)\big)+\big(\beta_{+}\big(\Phi^{\ell}\big)-\beta_{-}\big(\Phi^{\ell}\big)\big).$
Therefore, (IR2”) and (IR3”) hold again for all $k$ and $\ell$.
The remaining case is when $\Phi=\Phi_{e}$ is elliptic. We break it down into two subcases.
Subcase 1: $\Phi$ is elliptic and strongly non-degenerate. Then, by definition, no eigenvalue of $\Phi(1)$ is a root of unity and $\Phi^{k}$ is non-degenerate for all $k\in{\mathbb{N}}$. As a consequence,
$\mu_{-}\big(\Phi^{k}\big)=\mu\big(\Phi^{k}\big)=\mu_{+}\big(\Phi^{k}\big).$
By Lemma 5.3, without loss of generality, we may assume that map $\Phi(1)$ is semisimple. Indeed, consider an isospectral (i.e., with constant eigenvalues) family $\Psi_{s}$, $s\in[0,1]$, connecting $\Phi(1)$ to a semisimple linear symplectic map with the same eigenvalues as $\Phi(1)$. Let $\tilde{\Phi}_{s}\in\widetilde{\operatorname{Sp}}(2m)$ be obtained by concatenating $\Phi$ and $\Psi$. Then $\mu\big(\tilde{\Phi}_{s}^{k}\big)=\mu\big(\Phi^{k}\big)$ and $\operatorname{\hat{\mu}}\big(\tilde{\Phi}_{s}^{k}\big)=\operatorname{\hat{\mu}}\big(\Phi_{s}^{k}\big)$ for all $k$ and $s$, and hence in the proof of the theorem we can replace $\Phi$ by $\tilde{\Phi}$. We keep the notation $\Phi$.
Then, since $\Phi$ is semisimple, it decomposes as a direct sum of $m$ “short” rotations
$R_{q}(t):=\exp(2\pi\sqrt{-1}\lambda_{q}t)\colon{\mathbb{C}}\to{\mathbb{C}}$, (5.10)
where $0<|\lambda_{q}|\leq 1/2$, $q=1,\ldots,m^{\prime}$, and $t$ ranges from 0 to $1$. Since no eigenvalue is a root of unity all $\lambda_{q}$ are irrational. Furthermore, $\operatorname{\hat{\mu}}(R_{q})=2\lambda_{q}$ and $\mu(R_{q})=\mathit{sign}(\lambda_{q})$ and $\operatorname{\hat{\mu}}(\Phi)=\sum\operatorname{\hat{\mu}}(R_{q})$. Therefore, by additivity, we only need to establish (IR2”) and (IR3”) for one rotation $R_{q}$ to complete the proof of this subcase. For the sake of brevity, set $R:=R_{q}$ and $\lambda:=\lambda_{q}$. Then these identities amount to
$\mu\big(R^{k\pm\ell}\big)=\big[\operatorname{\hat{\mu}}\big(R^{k}\big)\big]\pm\mu\big(R^{\ell}\big),$ (5.11)
where $1\leq\ell\leq\ell_{0}$ and
$\|k\lambda\|<\min\{\|\lambda\ell\|\mid 1\leq\ell\leq\ell_{0}\}<1/2,$ (5.12)
as a combination of (5.3), (5.4) and (5.5). By (3.1),
$\operatorname{\hat{\mu}}\big(R^{k}\big)=2k\lambda\quad\textrm{and}\quad\mu\big(R^{k}\big)=\mathit{sign}(\lambda)\big(2\lfloor k|\lambda|\rfloor+1\big).$
Thus, (5.11) turns into the identity
$\mathit{sign}(\lambda)\big(2\lfloor(k\pm\ell)|\lambda|\rfloor+1\big)=[2k\lambda]\pm\mathit{sign}(\lambda)\big(2\lfloor\ell|\lambda|\rfloor+1\big).$
It is routine to see that this condition is indeed satisfied whenever (5.12) holds; see, e.g., [GG20, Sec. 5.2.1] for a detailed argument. This finishes the proof of the first subcase.
Subcase 2: Roots of unity. In this subcase all eigenvalues of $\Phi$ are roots of unity. This is the only case where it matters that $k$ is divisible by $D$, the common multiple of their degrees, which we will assume from now on. Hence, $\Phi^{k}$ is totally degenerate and $d=\operatorname{\hat{\mu}}\big(\Phi^{k}\big)$. Fixing $\ell$, let us decompose $\Phi$ as $\Phi_{0}\oplus\Phi_{1}$, where the degrees of the eigenvalues of $\Phi_{0}(1)$ divide $\ell$ and the degrees of the eigenvalues of $\Phi_{1}(1)$ do not. In other words, $\Phi_{1}^{\ell}$ and $\Phi_{1}^{k\pm\ell}$ are non-degenerate, and $\Phi_{0}^{\ell}$ and $\Phi_{0}^{k\pm\ell}$ are totally degenerate.
We claim that
$\mu\big(\Phi_{1}^{k\pm\ell}\big)=\operatorname{\hat{\mu}}\big(\Phi_{1}^{k}\big)\pm\mu\big(\Phi_{1}^{\ell}\big).$ (5.13)
In spite of its appearance, this statement is not entirely obvious and requires a proof. An easy way to prove (5.13) is as follows. By Lemma 5.3, there exists an isospectral family $\Psi_{s}$, $s\in[0,1]$, of transformations with $\Psi_{1}=\Phi_{1}$ and $\Psi_{0}(1)$ semisimple. Then $\Psi_{0}^{k}(1)={\mathit{i}d}$, i.e., $\Psi_{0}^{k}$ is a loop, and $\operatorname{\hat{\mu}}\big(\Phi_{1}^{k}\big)=\operatorname{\hat{\mu}}\big(\Psi_{0}^{k}\big)$ because all eigenvalues of $\Psi_{s}^{k}(1)$ are equal to 1 for all $s$. Furthermore, $\Psi_{s}^{k\pm\ell}$ and $\Psi_{s}^{\ell}$ remain non-degenerate for all $s$. As a consequence, $\mu\big(\Phi_{1}^{\ell}\big)=\mu\big(\Psi_{0}^{\ell}\big)$ and $\mu\big(\Phi_{1}^{k\pm\ell}\big)=\mu\big(\Psi_{0}^{k\pm\ell}\big)$. It follows that (5.13) holds if and only if it holds for $\Psi_{0}$. The latter, however, is obvious:
$\mu\big(\Psi_{0}^{k\pm\ell}\big)=\mu\big(\Psi_{0}^{k}\Psi_{0}^{\pm\ell}\big)=\operatorname{\hat{\mu}}\big(\Psi_{0}^{k}\big)\pm\mu\big(\Psi_{0}^{\ell}\big)$
since $\Psi_{0}^{k}$ is a loop.
Let $q$ be the least common multiple of the degrees of eigenvalues of $\Phi_{0}$. Then $\Phi_{0}^{q}$ is totally degenerate. Furthermore, $q$ divides both $k$ and $\ell$, and $\Phi_{0}^{\ell}=\big(\Phi_{0}^{q}\big)^{\ell/q}$ and $\Phi_{0}^{k+\ell}=\big(\Phi_{0}^{q}\big)^{(k+\ell)/q}$ are admissible iterates of $\Phi_{0}^{q}$. Hence
$\beta_{\pm}\big(\Phi_{0}^{k+\ell}\big)=\beta_{\pm}(\Phi_{0}^{q})=\beta_{\pm}\big(\Phi_{0}^{\ell}\big)$
by (3.4) applied to $\Phi_{0}^{q}$. Therefore, using (5.13) in the third line below, we have
$\begin{split}\mu_{\pm}\big(\Phi^{k+\ell}\big)&=\mu_{\pm}\big(\Phi_{1}^{k+\ell}\big)+\mu_{\pm}\big(\Phi_{0}^{k+\ell}\big)\\ &=\mu\big(\Phi_{1}^{k+\ell}\big)+\operatorname{\hat{\mu}}\big(\Phi_{0}^{k+\ell}\big)\pm\beta_{\pm}\big(\Phi_{0}^{k+\ell}\big)\\ &=\operatorname{\hat{\mu}}\big(\Phi_{1}^{k}\big)+\mu\big(\Phi_{1}^{\ell}\big)+\operatorname{\hat{\mu}}\big(\Phi_{0}^{k+\ell}\big)\pm\beta_{\pm}(\Phi_{0})\\ &=\big(\operatorname{\hat{\mu}}\big(\Phi_{1}^{k}\big)+\operatorname{\hat{\mu}}\big(\Phi_{0}^{k}\big)\big)+\mu\big(\Phi_{1}^{\ell}\big)+\big(\operatorname{\hat{\mu}}\big(\Phi_{0}^{\ell}\big)\pm\beta_{\pm}\big(\Phi_{0}^{\ell}\big)\big)\\ &=d+\mu_{\pm}\big(\Phi^{\ell}\big),\end{split}$
proving (IR2”). Condition (IR3”) is established by a similar calculation:
$\begin{split}\mu_{+}\big(\Phi^{k-\ell}\big)&=\mu_{+}\big(\Phi_{1}^{k-\ell}\big)+\mu_{+}\big(\Phi_{0}^{k-\ell}\big)\\ &=\operatorname{\hat{\mu}}\big(\Phi_{1}^{k}\big)-\mu\big(\Phi_{1}^{\ell}\big)+\operatorname{\hat{\mu}}\big(\Phi_{0}^{k-\ell}\big)+\beta_{+}\big(\Phi_{0}^{k-\ell}\big)\\ &=\big(\operatorname{\hat{\mu}}\big(\Phi_{1}^{k}\big)+\operatorname{\hat{\mu}}\big(\Phi_{0}^{k}\big)\big)-\mu\big(\Phi_{1}^{\ell}\big)-\big(\operatorname{\hat{\mu}}\big(\Phi_{0}^{\ell}\big)-\beta_{-}\big(\Phi_{0}^{\ell}\big)\big)\\ &\quad+\big(\beta_{+}\big(\Phi_{0}^{\ell}\big)-\beta_{-}\big(\Phi_{0}^{\ell}\big)\big)\\ &=d-\mu_{-}\big(\Phi^{\ell}\big)+\big(\beta_{+}\big(\Phi^{\ell}\big)-\beta_{-}\big(\Phi^{\ell}\big)\big).\end{split}$
In the last two lines we have used the fact that $\Phi_{1}^{\ell}$ is non-degenerate, and hence we can replace $\Phi_{0}^{\ell}$ by $\Phi^{\ell}$ when calculating $\beta_{\pm}$. This completes the proof of Subcase 2.
To finish the proof of the lemma, it remains to verify condition (IR4) asserting in this case that (IR1”), (IR2”) and (IR3”) with the same $k$ and $d$ continue to hold with $\Phi$ replaced by its non-degenerate part $\Psi$. To this end, we note that the elliptic eigenvalues of $\Psi(1)$ still satisfy (5.5). Thus, since $\operatorname{\hat{\mu}}(\Psi)=\operatorname{\hat{\mu}}(\Phi)$, using the self-explanatory notation as above, we have
$d\big(\Phi^{k}\big)=d\big(\Psi^{k}\big).$
Then a simple examination of the rest of the proof shows that (IR1”)(IR3”) still hold for $\Psi$. This completes the proof of Lemma 5.5. ∎
5.2.4. The general case: $r\geq 1$.
Let $(\Phi_{ij},a_{ij})$ be as in the statement of the theorem. We denote the total number of maps $\Phi_{ij}$ by $r$. As we have already pointed out, if we apply the argument from Section 5.2.3 to each of these maps individually, we obtain integer sequences $l\mapsto k_{ijl}$ and integer sequences $d_{ijl}$ such that conditions (IR1)(IR5) are satisfied with $d_{ijl}$ depending on $j$ and $C_{ijl}$ depending on $i$ and $j$. Thus our goal in this section is to modify the process to remove this dependence. This is where Lemma 5.2 enters the proof.
As in Section 5.2.3, denote the elliptic eigenvalues of $\Phi_{ij}$ by $\exp(\pm 2\pi\sqrt{-1}\lambda_{ijq})$, where $0<|\lambda_{ijq}|\leq 1/2$ and $q=1,\ldots,m_{ij}^{\prime}$, and for each pair of complex conjugate eigenvalues we pick the sign of $\lambda_{ijq}$ so that the resulting eigenvalue is of the first Krein type; see [SZ]. Set $\Delta_{ij}:=\operatorname{\hat{\mu}}(\Phi_{ij})$ and let $D$ be the least common multiple of the degrees of the roots of unity among the eigenvalues of all $\Phi_{ij}$.
Fix $\epsilon>0$ and $\sigma>0$, which we will eventually need to make sufficiently small, and consider the system of inequalities
$\|k_{ij}\lambda_{ijq}\|$ $<\epsilon\quad\textrm{ for all $i$, $j$ and $q$,}$ (5.14)
$|k_{11}a_{11}-k_{ij}a_{ij}|$ $<\sigma\quad\textrm{ for $i+j\geq 2$},$
where we treat the integer vector $\vec{k}\in{\mathbb{Z}}^{r}$ with components $k_{ij}$ as a variable. Introducing additional integer variables $c_{ijq}$, we can rewrite the first group of inequalities in the form
$|k_{ij}\lambda_{ijq}-c_{ijq}|<\epsilon.$ (5.15)
With this in mind, system (5.14) has one fewer equation than the number of variables.
This reformulation puts the question within the framework of Lemma 5.2. Namely, thinking for a moment of $k_{ij}$ and $c_{ijq}$ as real coordinates on ${\mathbb{R}}^{n}$ with $n=r+\sum m^{\prime}_{ij}$, we can view (5.14) as a system in the form of (5.1), where $f_{s}=k_{11}a_{11}-k_{ij}a_{ij}$ for $s=1,\ldots,r-1$ and the remaining linear maps are $k_{ij}\lambda_{ijq}-c_{ijq}$. By Lemma 5.2, there exists a bounded gap sequence $K_{l}\to\infty$ of non-zero solutions of (5.14) in ${\mathbb{Z}}^{n}$. Projecting to the first $r$-coordinates, we obtain a bounded gap sequence $\vec{k}_{l}\in{\mathbb{Z}}^{r}$ solving (5.14). Moreover, we can make all $k_{ijl}$ divisible by any fixed integer $N$ as in the lemma. In particular, we can have all $k_{ijl}$ divisible by $D$.
Our next goal is to ensure that $k_{ijl}\to\infty$ as $l\to\infty$ for all $i$ and $j$. To this end, we need to require that $\epsilon<1$ and $\sigma<\min a_{ij}$. (Recall that by the assumptions of the theorem $a_{ij}>0$ for all $i$ and $j$.) First, note that the sequence $\vec{k}_{l}$ is necessarily unbounded; for otherwise the entire sequence $K_{l}$ would be bounded due to (5.15). Furthermore, $\vec{k}_{l}\neq 0$ since otherwise we would have $K_{l}=0$ by again (5.15). (Here we use the fact that $0<\epsilon<1$.) In particular, for every $l$ at least one of the integers $k_{ijl}\neq 0$. Then by the second inequality in (5.14), for a fixed $l$, all $k_{ijl}$ are non-zero and have the same sign when $\sigma<\min a_{ij}$. Therefore, replacing $\vec{k}_{l}$ by $\vec{k}_{l}$ when necessary, we can guarantee that $k_{ijl}>0$ for all $i$, $j$ and $l$. Clearly, after this sign change the sequence still has bounded gap. Finally, since the norm of $\vec{k}_{l}$ goes to infinity, we must have $k_{ijl}\to\infty$ for all $i$ and $j$ again by the second inequality in (5.14).
It follows that when $\epsilon>0$ is sufficiently small the requirements of Lemma 5.5 are met for each $\Phi_{ij}$, and hence conditions (IR1)(IR4) hold with
$d_{ijl}:=[k_{ijl}\Delta_{ij}].$
More specifically, fix $\ell_{0}$ and $\eta>0$ which we assume to be sufficiently small (e.g., $\eta<1/4$) and similarly to (5.3), set
$\epsilon_{0}=\min_{i,j,q}\{\|\lambda_{ijq}\ell\|\mid\lambda_{ijq}\ell\not\in{\mathbb{Z}},1\leq|\ell|\leq\ell_{0}\}>0.$
Furthermore, we let $\epsilon>0$ be so small that again
$\epsilon\leq\epsilon_{0}\quad\text{and}\quad 2\epsilon\max_{i,j}m^{\prime}_{ij}<\eta$
as in (5.4). Then (5.5) is satisfied for each $\Phi_{ij}$.
Next, we claim that $d_{ijl}$ is in fact independent of $j$ when $\sigma>0$ is sufficiently small. To see this, let us focus on one particular cluster, i.e., fixing $i$ we have the ratio $\rho_{i}=\Delta_{ij}/a_{ij}>0$ independent of $j$. First, note that as a consequence of the second condition in (5.14),
$|k_{ij^{\prime}l}a_{ij^{\prime}}-k_{ijl}a_{ij}|<2\sigma.$
Therefore, multiplying this inequality by $\rho_{i}>0$, we have
$|k_{ij^{\prime}l}\Delta_{ij^{\prime}}-k_{ijl}\Delta_{ij}|<2\sigma\rho_{i}.$
It follows that $d_{il}:=d_{ijl}$ is independent of $j$ for all $i$ when $\sigma\max_{i}\rho_{i}<1/8$.
Turning to (IR5), we can pick any
$C_{l}\in(\max_{ij}k_{ijl}a_{ij},\,\min_{ij}k_{ijl}a_{ij}+\eta)$
lying outside the spectrum (i.e., the union of the sequences $a_{ij}{\mathbb{N}}$) assuming again that $\sigma$ is so small that $2\sigma<\eta$ and $\eta<\min a_{ij}$.
We have already shown that the sequences $k_{ijl}$ can be made divisible by any given integer $N$. This, however, does not automatically guarantee that $d_{il}$ is divisible by $N$. To achieve this, we argue exactly as in the case of one map. Namely, we first replace the upper bounds $\epsilon$ and $\sigma$ in (5.14) by $\epsilon/N$ and $\sigma/N$ and then multiply the resulting sequences $k_{ijl}$ by $N$. (This also makes $k_{ijl}$ divisible by $N$, although we have already guaranteed this by other means.) Then the sequence $d_{il}$ gets multiplied by $N$.